### v2

# Characterization of the directed landscape from the KPZ fixed point

Abstract
We show that the directed landscape is the unique coupling of the KPZ fixed point from all initial conditions at all times satisfying three natural properties: independent increments, monotonicity, and shift commutativity. Equivalently, we show that the directed landscape is the unique directed metric on $\mathbb{R}^{2}$ with independent increments and KPZ fixed point marginals. This gives a framework for proving convergence to the directed landscape given convergence to the KPZ fixed point. We apply this framework to prove convergence to the directed landscape for a range of models, some without exact solvability: asymmetric exclusion processes with potentially non-nearest neighbour interactions, exotic couplings of ASEP, the random walk and Brownian web distance, and directed polymer models. All of our convergence theorems are new except for colored ASEP and the KPZ equation, where we provide alternative proofs.
1 Introduction
The KPZ (Kardar-Parisi-Zhang) universality class is a loose term for a collection of $1$-dimensional models of random interface growth and $2$-dimensional random metric or polymer models that exhibit the same behaviour under rescaling. The past twenty-five years have seen a period of intense and fruitful research on this class, with progress propelled by the discovery of a handful of exactly solvable models, including exclusion processes, directed polymers, last passage percolation models, and the KPZ equation itself. See [Qua12, Cor12, Rom15, BG16, Zyg22, Gan22] and references therein for background on the KPZ universality class and related areas.
The KPZ fixed point is the universal scaling limit for growth models in the KPZ class, and the directed landscape is the universal scaling limit for metrics and polymer models. The KPZ fixed point can be constructed as a marginal of the directed landscape, and one may optimistically hope the opposite is also true: that we can construct the directed landscape from the KPZ fixed point. This is the goal of the present paper.
Random growth and the KPZ fixed point.
To describe the KPZ fixed point, we first consider one particular family of growth models: exclusion processes on $\mathbb{Z}$. Let $p$ be a measure on $\mathbb{Z}$ with finite support. An exclusion process on $\mathbb{Z}$ with jump distribution $p$ is an interacting particle system where particles occupy sites in $\mathbb{Z}$, with at most one particle per site. A particle at site $x$ will attempt to jump to site $x+v$ at rate $p(v)$, succeeding if the site $x+v$ is unoccupied (the exclusion rule). All particle jumps are independent, and to avoid degeneracies we assume that the symmetrized support $\{v\in\mathbb{Z}:p(v)+p(-v)>0\}$ additively generates $\mathbb{Z}$. We can associate to an exclusion process at time $t$ a height function $h_{t}:\mathbb{Z}\to\mathbb{Z}$, by letting
$h_{t}(x+1)=\begin{cases}h_{t}(x)+1,\quad\text{there is a particle at site $x$},\\ h_{t}(x)-1,\quad\text{there is a hole at site $x$}.\end{cases}$
This rule determines the height function up to an anchor point at $h_{t}(0)$. While we allow ourselves flexibility in choosing $h_{0}(0)$, in order to define the height functions consistently as $t$ changes, we stipulate that $h_{t}(0)=h_{t-}(0)-2$ if a particle jumps from the interval $(-\infty,-1]$ to the interval $[0,\infty)$ at time $t$, and that $h_{t}(0)=h_{t-}(0)+2$ if a particle jumps from $[0,\infty)$ to $(-\infty,-1]$ at time $t$. Otherwise we set $h_{t}(0)=h_{t-}(0)$.
When the expectation $\sum_{v\in\mathbb{Z}}vp(v)\neq 0$, we call such a process an asymmetric exclusion process (AEP). Without loss of generality, we may assume this expectation is positive, and by rescaling time, that it equals $1$. In an AEP, the height function decreases at a linear rate, and so the process $(h_{t})_{t\geq 0}$ produces an growing interface where the growth is governed by independent forcing, but is constrained by the local structure of the interface through the exclusion rule. These features are typical of growth models in the KPZ universality class. The special case when the jump distribution $p$ is supported on $\{-1,1\}$ is known as the asymmetric simple exclusion process (ASEP). If $p(1)=1$ and $p(v)=0$ for all $v\neq 1$ then this is the totally asymmetric simple exclusion process (TASEP). TASEP height function evolutions are particularly easy to understand: the height function evolves by a flipping local maxima to a local minima, $h_{t}(x)\mapsto h_{t}(x)-2$, at rate $1$.
Matetski, Quastel, and Remenik [MQR21] showed that for TASEP, the process
$h^{\epsilon}_{t}(x):=\epsilon^{1/2}h_{2\epsilon^{-3/2}t}(2\epsilon^{-1}x)+\epsilon^{-1}t$
converges as $\epsilon\to 0$ to a limiting Markov process whose state space is the set of upper semicontinuous functions from ${}\mathbb{R}\to\mathbb{R}\cup\{-\infty\}$ satisfying certain growth conditions. This limiting process is the KPZ fixed point. Convergence to the KPZ fixed point has since been verified for a handful of models, including solvable last passage percolation models [NQR20, DV21b], the KPZ equation and polymer models [QS23, Vir20], the AEPs described above [QS23], and more recently the stochastic six-vertex model [ACH24a].
Random metrics and coupled random growth.
The KPZ universality class also contains $2$-dimensional models of random metrics naturally built from a family of independent random variables or a noise on the plane. These models can be connected back to models of random growth in various ways, for example by considering the process of distances from a point $o$ to a collection of lines indexed by time $t$. Varying $o$ gives rise to a coupling of many different random growth processes. Conversely, if we are given a particularly natural coupling of random growth processes simultaneously from all initial conditions, then we can often realize this coupling through a random metric.
To make this discussion more concrete, consider Liggett’s basic coupling for exclusion processes started from different initial configurations [Lig76]. In the basic coupling for AEPs with jump distribution $p$, we can view each pair of sites $n,n+v\in\mathbb{Z}$ as being attached to an independent rate-$p(v)$ Poisson clock $\Pi_{n,v}$. When the Poisson clock $\Pi_{n,v}$ rings, in all of the AEPs, if there is a particle at site $n$ then this particle will attempt to jump to site $n+v$, succeeding whenever the exclusion rule allows it. We can also couple AEPs started at different times with this coupling. This coupling is equivalent to colored or multi-type AEP, see [AAV11, ACH24b] for background.
In the special case of TASEP, this coupling can be fully realized with a directed metric construction on $\mathbb{Z}\times\mathbb{R}$. For each $x\in\mathbb{Z}$, let $\Delta_{x}:y\mapsto-|x-y|$ be the narrow wedge centred at $x\in\mathbb{Z}$. Define the function
$d_{\Pi}(x,s;y,t)=h_{t}(y;\Delta_{x},s),$ (1)
for any $s\leq t$ and $x,y\in\mathbb{Z}$. Here the right hand side is the height function at time $t$ and location $y$ for TASEP started from $\Delta_{x}$ at time $s$ in the basic coupling through the Poisson clocks $\Pi=\{\Pi_{n,v}\}$. Set $d_{\Pi}(x,s;y,t)=-\infty$ for $t>s$. We have $d_{\Pi}(p;p)=0$ for all $p\in\mathbb{Z}\times\mathbb{R}$ and the function $d_{\Pi}$ satisfies a reverse triangle inequality
$d_{\Pi}(p;q)+d_{\Pi}(q;r)\leq d_{\Pi}(p,r).$ (2)
This makes $d_{\Pi}$ a directed metric, see Section 2 for background. Moreover, for any initial condition $h_{0}$, the TASEP evolution of $h_{0}$ started at time $s$ in the coupling governed by the clocks $\Pi$ is given by the variational formula
$h_{t}(\cdot;h_{0},s):=\max_{x\in\mathbb{Z}}[h_{0}(x)+d_{\Pi}(x,s;\cdot,t)].$ (3)
Note that if we replace TASEP by a general ASEP in (1) then formula (3) no longer holds, and if we further replace ASEP with a general AEP, then we lose the triangle inequality (2) as well.
Just as TASEP and other growth models converge to the KPZ fixed point, random metrics also have a KPZ scaling limit. Indeed, if we define
$d_{\Pi}^{\epsilon}(x,s;y,t)=\epsilon^{1/2}d_{\Pi}(2\epsilon^{-1}x,2\epsilon^{-3/2}s;2\epsilon^{-1}y,2\epsilon^{-3/2}t)+\epsilon^{-1}(t-s)$
then $d_{\Pi}^{\epsilon}$ converges as $\epsilon\to 0$ to the random function $\mathcal{L}:\mathbb{R}^{4}_{\uparrow}\to\mathbb{R},\mathbb{R}^{4}_{\uparrow}:=\{(x,s;y,t)\in\mathbb{R}^{4}:s<t\}$ known as the directed landscape, constructed in [DOV22] two years after the construction of the KPZ fixed point and with a completely different method. For this particular coupling of TASEP, this convergence was recently proven by Aggarwal, Corwin, and Hegde [ACH24b]. The directed landscape has also been shown to be the scaling limit of classical last passage percolation models [DOV22, DV21b], the KPZ equation [Wu23], and the stochastic six-vertex model [ACH24b]. The directed landscape is connected to the KPZ fixed point through a limiting version of (3). For an upper semi-continuous function $f:\mathbb{R}\to\mathbb{R}\cup\{-\infty\}$, the process
$\mathfrak{h}_{t}f(x)=\sup_{y\in\mathbb{R}}f(y)+\mathcal{L}(y,0;x,t)$ (4)
is a KPZ fixed point started from $f$ at time 00, and so the directed landscape gives a coupling of the KPZ fixed point, simultaneously from all times and all initial conditions.
1.1 Main results: characterization theorems
The directed landscape contains much more information than the KPZ fixed point. Indeed, through formula (4) it gives a coupling of the KPZ fixed point simultaneously from all initial conditions with the same driving noise. Nonetheless, one may hope that the law of the directed landscape can still be characterized through the KPZ point. In this paper we provide such a characterization. Our main theorem shows that the directed landscape is the unique directed metric on $\mathbb{R}^{2}$ with independent increments and KPZ fixed point marginals.
Setting some notation, for an upper semicontinuous function $f:\mathbb{R}\to\mathbb{R}\cup\{-\infty\}$, we write $(\mathfrak{h}_{t}f,t\geq 0)$ for the KPZ fixed point evolution started from the initial condition $f$. When we use this notation for random $f$, we always assume that the KPZ fixed point evolution is independent of $f$. Here and throughout the paper, we write $\mathbb{Q}^{4}_{\uparrow}=\mathbb{Q}^{4}\cap\mathbb{R}^{4}_{\uparrow}$.
Theorem 1.1.
Suppose that $\mathcal{M}:\mathbb{Q}^{4}_{\uparrow}\to\mathbb{R}$ is a random function satisfying:
  1. 1.
    (Triangle inequality) For any $o=(x,s)$, $p=(y,r)$, $q=(z,t)\in\mathbb{Q}^{2}$ with $s<r<t$, we have $\mathcal{M}(o;p)+\mathcal{M}(p;q)\leq\mathcal{M}(o;q)$.
  2. 2.
    (Independent increments) For any $s_{1}<s_{2}<\cdots<s_{k}\in\mathbb{Q}$, the increments $\mathcal{M}(\cdot,s_{i};\cdot,s_{i+1}):\mathbb{Q}^{2}\to\mathbb{R},i=1,\dots,k-1$ are independent.
  3. 3.
    (KPZ fixed point marginals) For any $p=(x,t)\in\mathbb{Q}^{2}$, and $y_{1},\dots,y_{k},s_{1},\dots,s_{k}\in\mathbb{Q}$ with each $s_{i}>0$, we have
    $\{\mathcal{M}(p;p+(y_{i},s_{i}))\}_{i=1}^{k}\stackrel{{\scriptstyle d}}{{=}}\{\mathfrak{h}_{s_{i}}\delta_{0}(y_{i})\}_{i=1}^{k}.$
    Here $\delta_{0}$ is the narrow wedge initial condition given by $\delta_{0}(0)=0$ and $\delta_{0}(x)=-\infty,x\neq 0$.
Then $\mathcal{M}$ has a continuous extension to $\mathbb{R}^{4}_{\uparrow}$ which is a directed landscape.
Remark 1.2.
  1. 1.
    The choice of $\mathbb{Q}$ is arbitrary, both in the above theorem and throughout the paper. We could replace it with any countable dense subset of $\mathbb{R}$. Note also that we could have replaced $\mathbb{Q}^{4}_{\uparrow}$ with $\mathbb{Q}^{4}_{\uparrow}\cap(\mathbb{R}\times I)^{2}$ for any open interval $I\subset\mathbb{R}$; the proof does not change and we would recover a directed landscape defined only on the strip ×$\mathbb{R}\times I$.
  2. 2.
    The KPZ fixed point marginal assumption can be alternately phrased as saying that for fixed $p\in\mathbb{Q}^{2}$, we have $\mathcal{M}(p;\cdot)\stackrel{{\scriptstyle d}}{{=}}\mathcal{L}(p;\cdot)$. Johansson and Rahman [JR21] and Liu [Liu22] derived different exact formulas for this law, neither of which is a special case of the KPZ fixed point formula. Theorem 1.1 can alternately be viewed as a uniqueness theorem for the directed landscape given either the Johansson-Rahman or Liu multi-time formulas.
  3. 3.
    The proof of Theorem 1.1 does not actually use much precise information about the directed landscape $\mathcal{L}$. Indeed, all we really need is that $\mathcal{L}$ satisfies the three conditions in Theorem 1.1 along with the following stationarity property:
    Property $\ast$:
    Let $\mathcal{B}_{1}$ and $\mathcal{B}_{2}$ be independent two-sided Brownian motions with drifts $a_{1}<a_{2}$. Let $\mathcal{R}_{1}(x)=\mathcal{B}_{1}(x)$, and $\mathcal{R}_{2}(x)=\mathcal{B}_{1}(x)+\max_{y\leq x}-\mathcal{B}_{1}(y)+\mathcal{B}_{2}(y)$. Fix $s<t$ and define $F_{i}(y)=\max_{x\in\mathbb{R}}\mathcal{R}_{i}(x)+\mathcal{L}(x,s;y,t)$. Then
    $(\mathcal{R}_{1},\mathcal{R}_{2})-(\mathcal{R}_{1}(0),\mathcal{R}_{2}(0))\stackrel{{\scriptstyle d}}{{=}}(F_{1},F_{2})-(F_{1}(0),F_{2}(0)).$
    Property $\ast$ was shown for the directed landscape in [BSS24]. However, in two-species TASEP the analogue much more classical, going back to Angel [Ang06] (see also [FM07] for $k$-species TASEP).
    This suggests that by combining Angel’s theorem, the KPZ fixed point convergence in [MQR21], and our present approach, one should be able to obtain an alternate construction the directed landscape as the scaling limit of colored TASEP. For brevity, we have not pursued this in the present paper.
We also give an alternate version of Theorem 1.1 that characterizes the directed landscape as the unique coupling of the KPZ fixed point satisfying certain natural conditions. This version, while not difficult to show given Theorem 1.1, may better demonstrate how the characterization links back to couplings of discrete models such as the AEP basic coupling described above.
Here and throughout this paper, we use $\operatorname{UC}$ to denote the space of all upper semicontinuous $f:\mathbb{R}\to\bar{\mathbb{R}}:=\mathbb{R}\cup\{\pm\infty\}$, and let $\operatorname{UC}_{0}\subset\operatorname{UC}$ be the subspace of all $f$, satisfying $f(x)<\infty$ for any $x\in\mathbb{R}$, and:
$\limsup_{x\to\pm\infty}\frac{f(x)}{x^{2}}\leq 0.$ (5)
Let $\operatorname{NW}\subset\operatorname{UC}_{0}$ be the set of functions $f:\mathbb{R}\to\mathbb{R}\cup\{-\infty\}$ such that $\{x:f(x)>-\infty\}$ is a finite subset of $\mathbb{Q}$, and $f(\mathbb{R})\subset\mathbb{Q}\cup\{-\infty\}$. In words, $\operatorname{NW}$ is the set of all finite rational combinations of narrow wedges.
Theorem 1.3.
Consider a family of random operators $\mathcal{K}_{s,t}:\operatorname{NW}\to\operatorname{UC}_{0},s\leq t\in\mathbb{Q}$ such that:
  1. 1.
    (KPZ fixed point marginals) For all $f\in\operatorname{NW}$ and $s\leq t\in\mathbb{Q}$, we have $\mathcal{K}_{s,t}f\stackrel{{\scriptstyle d}}{{=}}\mathfrak{h}_{t-s}f$.
  2. 2.
    (Independent increments) For any collection of disjoint intervals $[s_{i},t_{i}],i=1,\dots,k$, the operators $\mathcal{K}_{s_{i},t_{i}},i=1,\dots,k$ are independent.
  3. 3.
    (Monotonicity) For any $s\leq r<t\in\mathbb{Q}$ and any $f,g\in\operatorname{NW}$, if $g\leq\mathcal{K}_{s,r}f$ then $\mathcal{K}_{r,t}g\leq\mathcal{K}_{s,t}f$.
  4. 4.
    (Shift commutativity) For all $s<t\in\mathbb{Q},f\in\operatorname{NW},c\in\mathbb{Q}$, $\mathcal{K}_{s,t}(f+c)=\mathcal{K}_{s,t}f+c$.
Then there exists a directed landscape $\mathcal{L}$ such that almost surely, for any $s<t\in\mathbb{Q},f\in\operatorname{NW},x\in\mathbb{R}$ we have
$\mathcal{K}_{s,t}f(x)=\sup_{y\in\mathbb{R}}f(y)+\mathcal{L}(y,s;x,t).$ (6)
Remark 1.4.
  1. 1.
    As in the previous theorem, the time index set $\mathbb{Q}$ could be replaced with $P\cap I$ for any countable dense set $P$ and any open interval $I$. We could also replace $\mathbb{Q}$ with any countable dense set when defining $\operatorname{NW}$.
  2. 2.
    To check that assumptions $2-4$ above are natural, we can compare with properties in the basic coupling of AEPs. The independent increment property in this setting is immediate. Shift commutativity is also immediate, and comes from the flexibility we initially had in defining our height function at 00 given the particle description. Deterministic monotonicity can be readily checked for ASEP, but is not true for general AEPs. However, an approximate version still holds and passes to monotonicity in the limit (as we will show in Section 8.3).
  3. 3.
    The marginal condition $1$ in Theorem 1.3 is different from condition $3$ in Theorem 1.1: condition $3$ in Theorem 1.1 asks for matching multi-time marginals starting only from the narrow wedge initial condition, whereas $1$ in Theorem 1.3 asks only for fixed-time marginals, but from a dense set of initial conditions. In practice, these conditions are interchangeable; we have included the two versions for flexibility in applications.
  4. 4.
    Finally, we point out that the existence of a variational formula of the form (6) for some field $\mathcal{L}$ (not necessarily a directed landscape) is equivalent to saying that $\mathcal{K}_{s,t}$ satisfies max-preservation:
    $\mathcal{K}_{s,t}(f\vee g)=\mathcal{K}_{s,t}f\vee\mathcal{K}_{s,t}g.$
    Max-preservation fails for many prelimiting models (e.g. even for ASEP, AEP in the basic coupling), and we do not have a simple argument for how it arises in the limit. Rather, we prove this through utilizing the full power of Theorem 1.1 and the variational characterization (4) of the law of the KPZ fixed point.
In the statement of Theorem 1.3, we have restricted our initial conditions to the (essentially arbitrary) dense set $\operatorname{NW}$ mostly for convenience. Other initial conditions come along for the ride, as exhibited by the following corollary.
Corollary 1.5.
In the setting of Theorem 1.3, consider $f\in\operatorname{UC}_{0}$ satisfying
$\limsup_{z\to x,z\in\mathbb{Q}}f(z)=f(x),\qquad\text{ for all }x\in\mathbb{R}.$ (7)
Suppose that for some $s\leq t\in\mathbb{Q}$, there is a random variable $\mathcal{K}_{s,t}f\in\operatorname{UC}_{0}$ such that $\mathcal{K}_{s,t}f\stackrel{{\scriptstyle d}}{{=}}\mathfrak{h}_{t-s}f$ and $\mathcal{K}_{s,t}g\leq\mathcal{K}_{s,t}f$ whenever $g\in\operatorname{NW},g\leq f$. Then formula (6) holds almost surely for this $s,t$ and $f$.
Note that condition (7) will always hold with $\mathbb{Q}$ replaced by some countable dense set $Q$ which may depend on $f$. For example, letting $Q=\{\operatorname*{arg\,max}_{x\in[q,q^{\prime}]}f(x):q<q^{\prime}\in\mathbb{Q}\}$ gives such a set, where here we choose the rightmost $\operatorname*{arg\,max}$ if multiple points exist. Therefore in practice Corollary 1.5 can be applied to any function in $\operatorname{UC}_{0}$.
1.2 Main results: new framework for directed landscape convergence
The practical upshot of Theorems 1.1 and 1.3 is that if we can show that a model converges to the KPZ fixed point, then we can typically show that it converges to the directed landscape by checking a few straightforward conditions. This offers a new approach to proving convergence to the directed landscape. Note that all previous approaches to studying convergence rely on extra combinatorial structure coming from either some generalization of the Robinson-Schensted correspondence [DOV22, DZ24, DV21b, Wu23, DV24], or the Yang-Baxter equation [ACH24b].
Theorem 1.6.
The following models converge to the directed landscape after appropriate rescaling: the O’Connell-Yor polymer (Theorem 7.4), the KPZ equation (Theorem 7.2), the random walk web distance and Brownian web distance as conjectured in [VV23] (Theorem 6.5), AEPs under the basic coupling (Theorem 8.7), and ASEPs and TASEPs under various exotic couplings (Theorem 8.4).
For the ASEP exotic couplings (to be defined in Definition 8.2), we note that the particle coupling, where Poisson clocks are attached to labelled particles rather than sites, is included in this collection.
Remark 1.7.
  1. 1.
    The only models above where directed landscape convergence has been previously established are the KPZ equation [Wu23], and colored ASEP/TASEP [ACH24b] (which is a special case of both our AEP convergence theorem and our exotic coupling theorem for ASEP).
    For the many other models covered in Theorem 1.6, including the random walk/Brownian web distance, general AEPs under the basic coupling, and ASEP under general exotic couplings, such directed landscape convergence seems beyond the scope of all previous approaches, since the necessary precise combinatorial structures (such as line ensembles or the Yang-Baxter equation) are unknown or absent.
  2. 2.
    For brevity, we do not attempt to be comprehensive in applying our convergence framework. Beyond the models in Theorem 1.6, there are various other models where our convergence framework should apply straightforwardly. For example, for integrable last passage percolation (LPP) models, where directed landscape convergence was previously shown in [DOV22, DV21b], our Theorems 1.1 and 1.3 could be applied to get alternative proofs, using a priori known KPZ fixed point convergence (e.g. [MQR21] or [Liu22] for exponential LPP, [NQR20] for Brownian LPP, [JR21] for geometric LPP). Our framework could also be applied to obtain directed landscape convergence for variants of TASEP where the original method of [MQR21] has been extended to find formulas suitable for proving convergence to the KPZ fixed point, see [Ara20, MR23a, MR23b].
1.3 Structure of the paper and proof overview
After a short preliminary section, we prove Theorem 1.1 in Section 3, up to a key technical result that we leave for Section 4. These are the most important parts of the paper. We then prove Theorem 1.3 from Theorem 1.1 in Section 5. This section also contains a technical extension of Theorem 1.3 that we will use to study AEPs, where the known convergence to the KPZ fixed point is slightly more delicate. Finally, Sections 6, 7 and 8 prove the applications in Theorem 1.6. The proofs in the final sections are mostly straightforward, though there are some technical lemmas required to handle monotonicity of the more delicate models (in particular, general AEPs).
Broadly speaking, Theorem 1.1 can be viewed as a consequence of the fact that information flows along a countable set of coalescing geodesics in the directed landscape. Very roughly, because this countable set is lower-dimensional and the directed landscape is independent in far away regions, we can hope to characterize the directed landscape using only some lower-dimensional marginal information.
Setting some notations, we will write $\mathcal{M}_{s,t}(x,y)=\mathcal{M}(x,s;y,t)$ for the increment in $\mathcal{M}$ from time $s$ to time $t$, and similarly define $\mathcal{L}_{s,t}$ for the directed landscape. For functions $f:\mathbb{R}^{k}\to\mathbb{R},g:\mathbb{R}^{\ell}\to\mathbb{R},k,\ell\geq 1$, we define the metric composition $f\diamond g:\mathbb{R}^{k-1}\times\mathbb{R}^{\ell-1}\to\mathbb{R}$ by
$f\diamond g(x,y)=\sup_{z\in\mathbb{R}}f(x,z)+g(z,r).$ (8)
The directed landscape $\mathcal{L}$ is characterized by the law of the increment $\mathcal{S}:=\mathcal{L}_{0,1}$, known as the Airy sheet, independent increment and scale invariance properties, and the metric composition law
$\mathcal{L}_{s,t}=\mathcal{L}_{s,r}\diamond\mathcal{L}_{r,t},\qquad s<r<t.$
Indeed, we can view $\mathcal{S}$ as an analogue of the normal distribution, and $\mathcal{L}$ as an analogue of Brownian motion in the metric composition semigroup. We will prove Theorem 1.1 using a variant of the Lindeberg exchange method adapted to studying metric composition.
Call a function $\mathcal{M}$ satisfying the assumptions of Theorem 1.1 a pre-landscape. With the above framework in mind, we first show that any pre-landscape $\mathcal{M}$ can be extended to a continuous function on $\mathbb{R}^{4}_{\uparrow}$ satisfying the metric composition law, and with the property that for any $s<t$, and $f:\mathbb{R}\to\mathbb{R}$ we have
$f\diamond\mathcal{M}\stackrel{{\scriptstyle d}}{{=}}f\diamond\mathcal{L}.$ (9)
This is fairly quick, and we refer the reader to 3.1 for details. Given this structure, to show that =$\mathcal{M}\stackrel{{\scriptstyle d}}{{=}}\mathcal{L}$, it suffices to prove that $\mathcal{M}_{s,t}$ and $\mathcal{L}_{s,t}$ have the same finite dimensional distributions for all fixed $s<t$. We will outline how we check that for a set $F=\{(x_{1},y_{1}),(x_{2},y_{2})\}$, the two-point distribution $\mathcal{M}_{0,1}|_{F}$ matches that of $\mathcal{L}$. The general case uses the same idea. For $s\in[0,1]$ define $\mathcal{M}^{s}=\mathcal{L}_{0,s}\diamond\mathcal{M}_{s,1},$ so that $\mathcal{M}^{0}=\mathcal{M}_{0,1}$ and $\mathcal{M}^{1}=\mathcal{L}_{0,1}$. We will aim to show that with respect to the $L^{1}$-Wasserstein metric $d_{W}$, for $s<t=s+r$,
$d_{W}(\mathcal{M}^{s}|_{F},\mathcal{M}^{t}|_{F})=o(r).$ (10)
The equality in law between $\mathcal{M}^{0}|_{F}$ and $\mathcal{M}^{1}|_{F}$ will then follow by interpolating between $\mathcal{M}^{0}$ and $\mathcal{M}^{1}$ with $\mathcal{M}^{t}$ for $t$ in a fine mesh and applying the triangle inequality for $d_{W}$. To prove (10), we use the decomposition
$\mathcal{M}^{s}=\mathcal{L}_{0,s}\diamond\mathcal{M}_{s,t}\diamond\mathcal{M}_{t,1},\qquad\mathcal{M}^{t}=\mathcal{L}_{0,s}\diamond\mathcal{L}_{s,t}\diamond\mathcal{M}_{t,1}.$ (11)
We will construct a coupling of the increments $\mathcal{M}_{s,t},\mathcal{L}_{s,t}$ so that $\mathbb{P}[\mathcal{M}^{s}|_{F}=\mathcal{M}^{t}|_{F}]=1-O(r^{2/3+\delta})$ for some $\delta>0$. To see why this implies (10), observe that on a compact set $K$, KPZ scaling gives that $\|\mathcal{M}_{s,t}\|_{K},\|\mathcal{L}_{s,t}\|_{K}=O(r^{1/3})$, and so a crude estimate implies
$\mathbb{E}\|\mathcal{M}^{s}|_{F}-\mathcal{M}^{t}|_{F}\|_{2}=O(r^{1/3})\mathbb{P}[\mathcal{M}^{s}|_{F}\neq\mathcal{M}^{t}|_{F}]=O(r^{1+\delta}).$
Now to construct the coupling between $\mathcal{M}_{s,t},\mathcal{L}_{s,t}$, for $i=1,2$ let $A_{i}$ be the set of all $z\in\mathbb{R}$ where
$\mathcal{L}_{0,s}\diamond\mathcal{M}_{t,1}(x_{i},y_{i})\leq\mathcal{L}_{0,s}(x_{i},z)+\mathcal{M}_{t,1}(z,y_{i})+\log^{10}(r)r^{1/3}.$
The set $A_{i}$ is chosen so that in the candidate landscapes $\mathcal{M}^{s},\mathcal{M}^{t}$, the geodesic from $(x_{i},0)$ to $(y_{i},1)$ is in the set $A_{i}$ at times $s$ and $t$ with very high probability; this is guaranteed by the fact that the threshold $\log^{10}(r)r^{1/3}\gg r^{1/3}$. Therefore if we can couple $\mathcal{L}_{s,t},\mathcal{M}_{s,t}$ so that
$\max_{z\in A_{i}}\mathcal{L}_{0,s}(x_{i},z)+\mathcal{L}_{s,t}(z,y)=\max_{z\in A_{i}}\mathcal{L}_{0,s}(x_{i},z)+\mathcal{M}_{s,t}(z,y),\qquad\forall y\in A_{i},$ (12)
for $i=1,2$, then with high probability $\mathcal{M}^{s}|_{F}=\mathcal{M}^{t}|_{F}$. The existence of a coupling where (12) holds only for exactly one value of $i$ is immediate by the KPZ fixed point marginal property (9). These two couplings (for $i=1$ and $i=2$) are compatible if the sets $A_{1},A_{2}$ are well-separated from each other, since in this case the pieces $\mathcal{M}_{s,t}|_{A_{1}\times A_{1}},\mathcal{M}_{s,t}|_{A_{2}\times A_{2}}$ are approximately independent. Moreover, the couplings for $A_{1},A_{2}$ are still compatible even if $A_{1},A_{2}$ partially overlap, as long as whenever they overlap on an interval $I$, the functions $\mathcal{L}_{0,s}(x_{1},\cdot),\mathcal{L}_{0,s}(x_{2},\cdot)$ differ by a constant $C_{I}$ on $I$.
This type of compatibility can be viewed as roughly corresponding to the event where in a candidate landscape $\mathcal{M}^{t}$, if $\pi_{i},i=1,2$ are the geodesics from $(x_{i},0)$ and $(y_{i},1)$, then either $\pi_{1}(t),\pi_{2}(t)$ are far away from each other ($A_{1},A_{2}$ are separated), or else $\pi_{1}(t)=\pi_{2}(t)$ and the geodesics coalesce together well before time $s$ ($A_{1},A_{2}$ overlap and the functions $\mathcal{L}_{0,s}(x_{1},\cdot),\mathcal{L}_{0,s}(x_{2},\cdot)$ differ by a constant). In other words, the coalescence time for the geodesics $\pi_{1},\pi_{2}$ occurs far away from the interval $[s,t]$. If $\mathcal{M}^{s},\mathcal{M}^{t}$ were true directed landscapes, we could prove that this compatibility event holds with probability $1-r^{1-o(1)}$. For pre-landscapes, we will show the weaker bound $1-r^{2/3+\delta}$, which still gives (10). In practice, we will also work with a stationary version of $\mathcal{M}^{s}|_{F},\mathcal{M}^{t}|_{F}$ in order to facilitate technical computations.
On the Lindeberg exchange strategy and other related ideas.
Originating with Lindeberg’s proof of the central limit theorem [Lin22], the method of replacement or exchange has been generalized (see [Cha05, Cha06, Rot79, MOO10, CSZ17]), and has been applied to derive universality for many probabilistic models, including some in the KPZ class, e.g. the Sherrington-Kirkpatrick spin glass model [Cha05], random matrices [Cha05, TV11, TV10, KY17], last passage percolation in a thin rectangle [Sui06], and directed polymers [KQ18, Ran23, CSZ23, Tsa24].
There are some similarities between our proofs and some existing arguments; for example, the Tsai’s characterization of the critical stochastic heat flow [Tsa24], which appeared while we were finalizing this paper, also uses an exchange idea based on replacing time slices as in (11) to characterize a scaling limit. However, beyond the starting idea of ‘exchange’, our approach significantly differs from the usual Lindeberg method. In particular, the study of test functions and moments has no relevance in our proof. An explanation for this is that compared to the previous applications of the Lindeberg method to directed polymers, our setting of the directed landscape is in the strongest disorder regime, which is more sensitive to noise perturbations (see e.g. [GH24] and references therein), thereby requiring very different techniques.
Another related idea used at the limit shape level is that of constructive Euler hydrodynamics, e.g. see Theorem 3.1 in [BGRS02] and [BGRS19] for a survey. This strategy characterizes the entropy solution to Burgers’ equation using a set of axioms similar to some of those in our Theorem 1.3. For example, they require a version of monotonicity and a local dynamics assumption in their deterministic setting. The axioms are tailored to study hydrodynamic limit points of interacting particle systems.
Acknowledgments. The authors thank Jeremy Quastel for several useful discussions regarding convergence of general AEPs. The authors also thank Amol Aggarwal for pointing out the argument in [ACH24b, Appendix D.3], and the references [BGRS02, BGRS19], and Ivan Corwin, Milind Hegde, and Shalin Parekh for useful comments on an earlier version of the paper.
2 Preliminaries
Throughout this paper, we denote $\mathbb{Q}_{+}:=\mathbb{Q}\cap(0,\infty)$, $\mathbb{Z}_{+}:=\mathbb{Z}\cap(0,\infty)$, $\mathbb{Z}_{-}:=\mathbb{Z}\cap(-\infty,0)$, $\bar{\mathbb{R}}:=\mathbb{R}\cup\{\pm\infty\}$ and $\mathbb{R}^{4}_{\uparrow}:=\{(x,s;y,t)\in\mathbb{R}^{4}:s<t\}$ (as defined in the introduction), and $\llbracket a,b\rrbracket:=[a,b]\cap\mathbb{Z}$ for any $a<b\in\bar{\mathbb{R}}$. The metric composition notation $\diamond$ from (8) will also be frequently used.
2.1 The directed landscape and the KPZ fixed point
Recall that $\operatorname{UC}$ is the set of all upper semicontinuous functions from $\mathbb{R}\to\bar{\mathbb{R}}$. Then a function $f$ is in $\operatorname{UC}$ if and only if its hypograph
$\operatorname{hypo}(f)=\{(x,t)\in\mathbb{R}\times\bar{\mathbb{R}}:t\leq f(x)\}$
is closed. We equip $\operatorname{UC}$ with a local Hausdorff metric
$d_{H,\operatorname{loc}}(f,g):=d_{H}(E(\operatorname{hypo}(f_{n}),E(\operatorname{hypo}(f)).$
in the Hausdorff metric on closed subsets of ×[1,$\mathbb{R}\times[-1,1]$, where $E:\mathbb{R}\times\bar{\mathbb{R}}\to\mathbb{R}\times[-1,1]$ is the map $E(x,s)=(x,e^{-|x|}s/(1+|s|)),$ and the $d_{H}$ metric on the right-hand side above is the usual Hausdorff metric. It is straightforward to check that $\operatorname{UC}$ is a compact Polish space. Setting some notations, we write $\delta_{x}\in\operatorname{UC}$ for the narrow wedge at $x$, i.e., the function with $\delta_{x}(x)=0$ and $\delta_{x}(y)=-\infty$ for $y\neq x$. For $f\in\operatorname{UC}$ and $K\subset\mathbb{R}$ we also write $f|_{K}\in\operatorname{UC}$ for the function equal to $f$ on $K$, and $\infty$- ∞ on $K^{c}$.
In [MQR21], the KPZ fixed point is constructed as a time-homogeneous Markov process on an appropriate subset of $\operatorname{UC}$, taking a function $f$ and mapping it to $\mathfrak{h}_{t}f\in\operatorname{UC}$, for any $t\geq 0$. More precisely, in [MQR21] the KPZ fixed point is defined by specifying the translation probabilities
$\mathbb{P}[\mathfrak{h}_{t}f\leq g]$ (13)
[…]
An alternate characterization of the KPZ fixed point comes through the directed landscape, constructed in [DOV22]. The directed landscape is built from the Airy sheet, a random continuous function $\mathcal{S}:\mathbb{R}^{2}\to\mathbb{R}$. We call the rescaling $\mathcal{S}_{s}(x,y)=s\mathcal{S}(x/s^{2},y/s^{2})$ an Airy sheet of scale $s$. For the purposes of this paper, we will take the Airy sheet as a black box and only invoke its properties as necessary. Here recall that for $f:\mathbb{R}^{k}\to\mathbb{R},g:\mathbb{R}^{\ell}\to\mathbb{R}$ we define $f\diamond g:\mathbb{R}^{\ell-1}\times\mathbb{R}^{k-1}\to\mathbb{R}$ by $f\diamond g(x,y)=\sup_{z\in\mathbb{R}}f(x,z)+g(z,y)$.
Definition 2.1.
The directed landscape $\mathcal{L}:\mathbb{R}^{4}_{\uparrow}\to\mathbb{R},(x,s;y,t)\mapsto\mathcal{L}(x,s;y,t)=\mathcal{L}_{s,t}(x,y)$ is the unique random continuous function satisfying
  1. I.
    (Independent increments) For any disjoint time intervals $\{[t_{i},s_{i}]:i\in\{1,\dots k\}\}$, the random functions $\{\mathcal{L}_{s_{i},t_{i}}:i\in\{1,\dots,k\}\}$ are independent.
  2. II.
    (Metric composition law) Almost surely, for any $r<s<t$ we have $\mathcal{L}_{r,t}=\mathcal{L}_{r,s}\diamond\mathcal{L}_{s,t}.$
  3. III.
    (Airy sheet marginals) For any $t\in\mathbb{R}$ and $s>0$, $\mathcal{L}_{t,t+s^{3}}\stackrel{{\scriptstyle d}}{{=}}\mathcal{S}_{s}$, an Airy sheet of scale $s$.
Remark 2.2.
The uninitiated reader may wish to compare Definition 2.1 with the conditions in Theorem 1.1. The key difference is in the assumption on the marginals. In Theorem 1.1, once we prove metric composition for $\mathcal{M}$ it is not difficult to show that condition $3$ is equivalent to knowing that for any fixed $f,g$ we have the identity $f\diamond\mathcal{L}_{t+s^{3}}\diamond g\stackrel{{\scriptstyle d}}{{=}}f\diamond\mathcal{S}_{s}\diamond g$. Without any extra structure, this marginal information would not be enough to see that $\mathcal{L}_{t+s^{3}}\stackrel{{\scriptstyle d}}{{=}}\mathcal{S}_{s}$, and the bulk of the proof of Theorem 1.1 is devoted to proving this. The difference between the assumption of metric composition and the weaker triangle inequality is easier to overcome.
As already alluded to, we can alternately describe the KPZ fixed point directly in terms of the directed landscape. Indeed, for $f\in\operatorname{UC}$, define $\mathfrak{h}_{0}f=f$, and $\mathfrak{h}_{t}f$ for $t>0$ through (4) (i.e., $\mathfrak{h}_{t}f:=f\diamond\mathcal{L}_{0,t}$);
[…]
With definition (4) in mind, Sarkar and Virág [SV21] identified the domain on which the KPZ fixed point is finite at all times, which is the space $\operatorname{UC}_{0}$ from (5).
Proposition 2.3.
(Domain of the KPZ fixed point, [SV21, Proposition 6.1]) Let $f\in\operatorname{UC}$. Then $f\in\operatorname{UC}_{0}$ if and only if almost surely, $\mathfrak{h}_{t}f(x)<\infty$ all $x\in\mathbb{R},t>0$. Moreover, if $f\in\operatorname{UC}_{0}$ then $\mathfrak{h}_{t}f(x)$ is continuous on the set $x\in\mathbb{R},t>0$ and almost surely, $\mathfrak{h}_{t}f\in\operatorname{UC}_{0}$ for all $t>0$.
A close relative of Proposition 2.3 is a dominated convergence theorem for the KPZ fixed point.
Proposition 2.4.
(DCT for the KPZ fixed point) Suppose that $f_{n}\to f\in\operatorname{UC}$ and that for some $g\in\operatorname{UC}_{0}$ we have $f_{n}\leq g$ for all $n$. Then in the landscape coupling, $\mathfrak{h}_{t}f_{n}(x)\to\mathfrak{h}_{t}f(x)$ a.s., uniformly on compact subsets of $(x,t)\in\mathbb{R}\times(0,\infty)$. In particular, for fixed $t>0$, $\mathfrak{h}_{t}f_{n}\stackrel{{\scriptstyle d}}{{\to}}\mathfrak{h}_{t}f$ as random variables in $\operatorname{UC}$.
2.4 is almost immediate from the following shape theorem for the directed landscape, which we will refer to frequently to establish basic tail bounds in the paper.
For this proposition and throughout, we write $d(x,s;y,t)=(x-y)^{2}/(t-s)$.
Proposition 2.5.
(Shape theorem for $\mathcal{L}$, [DOV22, Corollary 10.7]) There exists a random constant $R>0$ satisfying $\mathbb{P}[R>a]<ce^{-da^{3/2}}$ for universal constants $c,d>0$ and all $a>0$, such that for all $u=(x,t;y,t+s)\in\mathbb{R}^{4}_{\uparrow}$ we have
$\Big{|}[\mathcal{L}+d](u)\Big{|}\leq Rs^{1/3}\log^{4/3}\left(\frac{2(\|u\|_{2}+2)}{s}\right)\log^{2/3}(\|u\|_{2}+2).$
Proofs of 2.4.
By the bounds in Proposition 2.4 and existence of the dominating function $g$, we can see that for any $t_{0}>0$, there exists a random variable $M>0$ such that for all $n\in\mathbb{N}$ and $L\geq M$ (and with $f$ in place of $f_{n}$) we have
$\mathfrak{h}_{t}f_{n}(x)=\max_{y\in[-L,L]}f_{n}(y)+\mathcal{L}(y,0;x,t)$
whenever $|x|\leq t_{0},t\leq t_{0}$. Continuity of $\mathcal{L}$ and the convergence $f_{n}\to f$ then implies that $\mathfrak{h}_{t}f_{n}(x)\to\mathfrak{h}_{t}f(x)$ uniformly on compact subsets of $(x,t)\in\mathbb{R}\times(0,\infty)$. ∎
2.2 Basic properties of the directed landscape
The directed landscape is best viewed as a random directed metric. Following [DV21b, Section 5], a directed metric of positive sign on a set $X$ is a function $d:X\times X\to\mathbb{R}\cup\{\infty\}$ satisfying $d(x,x)=0$ for all $x$, together with the triangle inequality
$d(v,w)+d(w,u)\geq d(u,v).$
We say that $d$ is a directed metric of negative sign if $-d$ is a directed metric of positive sign. Many natural KPZ models are directed metrics but not metrics, and moreover, unlike the class of metrics, the class of directed metrics is preserved under the kinds of rescaling typical seen in the KPZ class. If we extend the directed landscape to equal $\infty$- ∞ on $\mathbb{R}^{4}\setminus\mathbb{R}^{4}_{\uparrow}$ then it becomes a directed metric (of negative sign) on $\mathbb{R}^{2}$.
As with usual metrics, we can define path length and geodesics in the directed landscape. Paths will be continuous functions $\pi:[s,t]\to\mathbb{R}$ with length
$|\pi|_{\mathcal{L}}=\inf_{k\in\mathbb{N}}\inf_{r_{0}=s<r_{1}<\dots<r_{k-1}<r_{k}=t}\sum_{i=1}^{k}\mathcal{L}(\pi(r_{i-1}),r_{i-1};\pi(r_{i}),r_{i}).$
A path is a geodesic if $|\pi|_{\mathcal{L}}=\mathcal{L}(\pi(s),s;\pi(t),t)$, and geodesics exist between all points in the directed landscape [DOV22, Lemma 13.2]. The concept of geodesics is central to the proof of the main theorem. However, we will typically not use geodesics explicitly and instead think about them through the lens of metric composition. Indeed, a point $(y_{0},r)$ lies on a geodesic from $(x,s)$ to $(z,t)$ if it is an argmax for the following metric composition:
$y_{0}\in\operatorname*{arg\,max}_{y\in\mathbb{R}}\mathcal{L}_{s,r}(x,y)+\mathcal{L}_{r,t}(y,z).$ (14)
Geodesics between nearby points in the directed landscape coalesce [BGH22], which manifests itself through the fact that the argmax $y_{0}$ above is typically stable if we perturb $x$ or $z$. Coalescence also implies that for $x\neq x^{\prime}$, the difference profile $\mathcal{L}_{s,t}(x,\cdot)-\mathcal{L}_{s,t}(x^{\prime},\cdot)$ is constant off of a lower dimensional set [BGH21, BGH22, GZ22].
It will also be useful to note the following basic symmetries for $\mathcal{L}$, which we use frequently and without reference.
Lemma 2.6 (Lemma 10.2, [DOV22] and Proposition 1.23, [DV21b]).
We have the following equalities in distribution as continuous functions from $\mathbb{R}^{4}_{\uparrow}\to\mathbb{R}$. Here $r,c\in\mathbb{R}$, and $q>0$.
  1. 1.
    (Time stationarity)
    $\mathcal{L}(x,s;y,t)\stackrel{{\scriptstyle d}}{{=}}\mathcal{L}(x,s+r;y,t+r).$
  2. 2.
    (Spatial stationarity)
    $\mathcal{L}(x,s;y,t)\stackrel{{\scriptstyle d}}{{=}}\mathcal{L}(x+c,s;y+c,t).$
  3. 3.
    (Flip symmetries)
    $\mathcal{L}(x,s;y,t)\stackrel{{\scriptstyle d}}{{=}}\mathcal{L}(-x,s;-y,t)\stackrel{{\scriptstyle d}}{{=}}\mathcal{L}(y,-t;x,-s).$
  4. 4.
    (Skew stationarity)
    $\mathcal{L}(x,s;y,t)\stackrel{{\scriptstyle d}}{{=}}\mathcal{L}(x+cs,s;y+ct,t)+(t-s)^{-1}[(x-y)^{2}-(x-y-(t-s)c)^{2}].$
  5. 5.
    ($1:2:3$ rescaling)
    $\mathcal{L}(x,s;y,t)\stackrel{{\scriptstyle d}}{{=}}q\mathcal{L}(q^{-2}x,q^{-3}s;q^{-2}y,q^{-3}t).$
In particular, we have the following identity at the level of the one-point law. For fixed $x,s;y,t$,
$\mathcal{L}(x,s;y,t)\stackrel{{\scriptstyle d}}{{=}}(t-s)^{1/3}\mathcal{L}(0,0;0,1)-\frac{(x-y)^{2}}{t-s}.$
The random variable $T=\mathcal{L}(0,0;0,1)$ has the Tracy-Widom GUE distribution; this goes back to [BDJ99]. The identity above shows that all one-point marginals of $\mathcal{L}$ are affine shifts of this law. The symmetries $1$ and $3$ above imply that for any $f\in\operatorname{UC}_{0}$ and $s\in\mathbb{R}$,
$(f\diamond\mathcal{L}_{s,s+t})_{t\geq 0}\stackrel{{\scriptstyle d}}{{=}}(\mathcal{L}_{s-t,s}\diamond f)_{t\geq 0}.$
This equality is known as skew-time reversibility of the KPZ fixed point, and plays an important role in Section 3. It was first identified in [MQR21].
Again using Lemma 2.6, if we fix three of the coordinates of $\mathcal{L}$ and vary only one of the spatial coordinates $x$ or $y$ we end up with a rescaling of the parabolic Airy2 process $\mathcal{A}(x):=\mathcal{L}(0,0;x,1)$. We will need a few basic facts about the Airy2 process. In this next lemma, we record a standard comparison with a Brownian bridge with diffusion coefficient $\sqrt{2}$ (which, at least for the direction of Airy2 to Brownian bridge, has been known since [CH14, Proposition 4.1]). Throughout the paper, our convention for all Brownian motions and bridges is that they have diffusion coefficient $2$ unless otherwise stated.
Lemma 2.7.
Let $\mathcal{A}$ be the parabolic Airy2 process, fix a closed interval $[a,b]$, and let $L:[a,b]\to\mathbb{R}$ be the random affine function with $L(a)=\mathcal{A}(a),L(b)=\mathcal{A}(b)$. Then conditional on $\mathcal{A}|_{[a,b]^{c}}$, the laws of $\mathcal{A}|_{[a,b]}-L$ and a Brownian bridge $B:[a,b]\to\mathbb{R}$ with $B(a)=B(b)=0$ are mutually absolutely continuous.
Proof.
The process $\mathcal{A}=\mathcal{A}_{1}$ is the top curve in a sequence of continuous functions $\mathcal{A}_{i}:\mathbb{R}\to\mathbb{R},\mathcal{A}_{1}>\mathcal{A}_{2}>\dots$ known as the parabolic Airy line ensemble [PS02, CH14]. Corwin and Hammond [CH14] showed that the parabolic Airy line ensemble satisfies the following resampling property: if we condition on $\mathcal{A}_{i}(x)$ for all $(i,x)$ outside of some compact set $\{1,\dots,k\}\times[a,b]$ then $(\mathcal{A}_{1},\dots,\mathcal{A}_{k})|_{[a,b]}$ is simply given by $k$ independent Brownian bridges from $\mathcal{A}_{i}(a)$ to $\mathcal{A}_{i}(b)$ conditioned to not intersect. Applying this with $k=1$ implies that conditional on $\mathcal{A}|_{[a,b]^{c}}$, $\operatorname{Law}(\mathcal{A}|_{[a,b]}-L)\ll\operatorname{Law}(B)$.
On the other hand, it is not difficult to show that $\sup\mathcal{A}_{k}|_{[a,b]}\to-\infty$ almost surely as $k\to\infty$. This follows, for example, from the modulus of continuity in [DV21a, Theorem 1.4] and the one-point bound in [Sos00, (1.17), Theorem 1]. Therefore for any set $U$ with $\mathbb{P}[B\in U]>0$, there exists a random $K\in\mathbb{N}$ such that if we condition on $A_{i}(x)$ for $(i,x)$ outside of some compact set $\{1,\dots,k\}\times[a,b]$, then $\mathbb{P}[\mathcal{A}|_{[a,b]}-L\in U]>0$ as well. Hence conditional on $\mathcal{A}|_{[a,b]^{c}}$, $\operatorname{Law}(B)\ll\operatorname{Law}(\mathcal{A}|_{[a,b]}-L)$. ∎
We also require a quantitative Brownian comparison. This is much more difficult to achieve than the soft comparison in Lemma 2.7, but still uses the Gibbs resampling property in the previous proof as the key input.
Lemma 2.8 ([Dau24, Corollary 1.3]).
Fix $a<b\in\mathbb{R}$. The law $\nu$ of $\mathcal{A}(t)-\mathcal{A}(a),t\in[a,b]$ is absolutely continuous respect to the law $\mu$ of a Brownian motion in $[a,b]$ with $B(a)=0$, with Radon-Nikodym derivative $\tfrac{d\nu}{d\mu}\in L^{\infty}(\mu)$.
A quantitative comparison with Brownian motion for $\mathcal{A}$ was first achieved in [CHH23, Theorem 1.1] (see also [Ham22] for an earlier comparison to Brownian bridge). In that paper the authors showed that $\tfrac{d\nu}{d\mu}\in L^{p}(\mu)$ for all $p<\infty$, a result that would also suffice for our purposes.
2.3 The stationary horizon
We end the preliminary section by describing the stationary measure for the directed landscape. This stationary measure is the stationary horizon, constructed in [Bus24, SS23] and connected to the directed landscape in [BSS24]. The last-passage description we give below for the stationary horizon is from [DV24]. It is a version of the queuing-theoretic description given in the previously mentioned papers.
First, for a collection of functions $f=(f_{1},\dots,f_{k}):\mathbb{R}\to\mathbb{R}^{k}$, and a vector $\pi=(\pi_{1}\geq\dots\geq\pi_{k})\in\mathbb{R}^{k}$, define
$|\pi|_{f}=f_{k}(\pi_{k})+\sum_{j=1}^{k-1}f_{j}(\pi_{j})-f_{j}(\pi_{j+1}),\qquad f[k\to x]=\sup_{\pi\in\mathbb{R}^{k}_{\geq},\pi_{1}=x}|\pi|_{f}.$
Here $f[i\to x]$ is a last passage value from $(i,-\infty)$ to $(x,1)$ in the sense of [DV24], and $|\pi|_{f}$ can be thought of as the length of a path associated to $\pi$. Now let $a_{1}<\dots<a_{k}$ and let =$\mathcal{B}=\{\mathcal{B}_{i}\}_{i=1}^{k}$ be a family of independent two-sided Brownian motions where each $\mathcal{B}_{i}$ has slope $a_{i}$. Define =$\mathcal{R}=\{\mathcal{R}_{i}\}_{i=1}^{k}$ by
$\mathcal{R}_{i}(x)=\tilde{\mathcal{R}}_{i}(x)-\tilde{\mathcal{R}}_{i}(0),\qquad\tilde{\mathcal{R}}_{i}(x)=\mathcal{B}[i\to x].$ (15)
The following theorem generalizes Property * from the introduction.
Theorem 2.9.
(contained in [BSS24, Theorem 2.1]) For any $a_{1}<\dots<a_{k}\in\mathbb{R}$, if we let $\mathcal{R}$ be defined as above and independent of the directed landscape, then for any $s<t$,
$(\mathcal{R}_{1},\dots,\mathcal{R}_{k})\stackrel{{\scriptstyle d}}{{=}}(\mathcal{R}_{1}\diamond\mathcal{L}_{s,t},\dots\mathcal{R}_{k}\diamond\mathcal{L}_{s,t})-(\mathcal{R}_{1}\diamond\mathcal{L}_{s,t}(0),\dots\mathcal{R}_{k}\diamond\mathcal{L}_{s,t}(0)).$
Moreover, $(\mathcal{R}_{1},\dots,\mathcal{R}_{k})$ is the unique (in law) $k$-tuple of functions satisfying the stationarity property above and the asymptotic slope condition
$\lim_{t\to\pm\infty}\frac{\mathcal{R}_{i}(t)}{t}=a_{i},\qquad i=1,\dots,k.$
3 Main characterization: proof of Theorem 1.1
In this section we prove Theorem 1.1, up to the key technical input which we leave to the next section. Throughout the section, we call a function $\mathcal{M}$ satisfying the assumptions of Theorem 1.1 a pre-landscape.
3.1 Continuity and metric composition
The first step in the proof of Theorem 1.1 is extending any pre-landscape to a continuous function on all of $\mathbb{R}^{4}_{\uparrow}$. This continuity claim follows from skew-time reversibility and continuity properties of the KPZ fixed point. As a byproduct of the proof, we will show that the continuous extension of $\mathcal{M}$ has KPZ fixed point marginals in both the forward and backwards directions and satisfies metric composition, rather than just a triangle inequality.
Proposition 3.1.
Let $\mathcal{M}$ be a pre-landscape. Then almost surely, $\mathcal{M}$ has a continuous extension $\mathcal{M}:\mathbb{R}^{4}_{\uparrow}\to\mathbb{R}$. Moreover, for any fixed function $f\in\operatorname{UC}_{0}$ and any $t\in\mathbb{R}$, $s\geq 0$, we have that
$\mathfrak{h}_{s}f\stackrel{{\scriptstyle d}}{{=}}\mathcal{M}_{t-s,t}\diamond f\stackrel{{\scriptstyle d}}{{=}}f\diamond\mathcal{M}_{t-s,t}.$ (16)
where here the equality in law is as functions in $\operatorname{UC}_{0}$. Finally, almost surely, for any $s<r<t$ we have the metric composition law $\mathcal{M}_{s,r}\diamond\mathcal{M}_{r,t}=\mathcal{M}_{s,t}$.
We break the proof of 3.1 into a series of lemmas building up properties of $\mathcal{M}$.
Lemma 3.2.
Let $f\in\operatorname{UC}_{0}$ with $\text{supp}(f)\subset\mathbb{Q}$ and $t-s,t\in\mathbb{Q}$. Then $\mathfrak{h}_{s}f|_{\mathbb{Q}}\stackrel{{\scriptstyle d}}{{=}}(\mathcal{M}_{t-s,t}\diamond f)|_{\mathbb{Q}}.$
Proof.
Let $g\in\operatorname{UC}_{0}$ be any function supported on a finite set $F\subset\mathbb{Q}$. Take $r\in\mathbb{Q},r<t-s$, and $A\subset\{h\in\operatorname{UC}_{0}:h\geq g\}$ such that $\mathbb{P}[\mathcal{L}_{r,t-s}(0,\cdot)\in A]>0$. Then
$\mathbb{P}\left[\mathcal{M}_{t-s,t}\diamond f\leq-g\right]$ $=\mathbb{P}\left[g\diamond\mathcal{M}_{t-s,t}\leq-f\right]$
$\geq\mathbb{P}\left[\mathcal{M}_{r,t-s}(0,\cdot)\diamond_{\mathbb{Q}}\mathcal{M}_{t-s,t}\leq-f\;\big{\mid}\;\mathcal{M}_{r,t-s}(0,\cdot)\in A\right]$
$\geq\mathbb{P}\left[\mathcal{M}_{r,t}(0,\cdot)\leq-f\;\big{\mid}\;\mathcal{M}_{r,t-s}(0,\cdot)\in A\right].$
In the second line, in the metric composition law $\diamond_{\mathbb{Q}}$, we take a supremum over all $z\in\mathbb{Q}$ rather than over all $z\in\mathbb{R}$ as in the usual definition. This is required since $\mathcal{M}$ is not defined off of $\mathbb{Q}^{4}_{\uparrow}$. When we write $\mathcal{M}_{r,t-s}(0,\cdot)\in A$ we mean that $\mathcal{M}_{r,t-s}(0,\cdot)$ extends continuously to a function in $A$. This is a positive probability event since the same holds for $\mathcal{L}$ and we have the equality in law in condition $3$ of Theorem 1.1.
In the above computation, the first inequality uses the independent increment property of $\mathcal{M}$, and the second inequality uses the triangle inequality, which implies $\mathcal{M}_{r,t-s}\diamond_{\mathbb{Q}}\mathcal{M}_{t-s,t}\leq\mathcal{M}_{r,t}.$ Now, the final probability above is the same in both $\mathcal{L}$ and $\mathcal{M}$ by condition $3$ of Theorem 1.1, and so by the metric composition law and independent increment properties for $\mathcal{L}$, it is bounded below by
$\inf_{h\in A}\mathbb{P}[h\diamond\mathcal{L}_{t-s,t}\leq-f].$ (17)
Next, for any $\epsilon>0$, define the set $A_{\epsilon}=\{h\in\operatorname{UC}_{0}:g\leq h\leq\sup g+1,d_{H,\operatorname{loc}}(h,g)\leq\epsilon\}$. Lemma 2.7 and the almost sure boundedness of the Airy2 process $\mathcal{A}$ (see 2.5) shows that $\mathbb{P}[\mathcal{M}_{r,t-s}(0,\cdot)\in A_{\epsilon}]>0$. Moreover, since any $g^{\prime}\in A_{\epsilon}$ in bounded above by $\sup g+1$, 2.4 implies that
$\lim_{\epsilon\to 0}\inf_{g_{\epsilon}\in A_{\epsilon}}\mathbb{P}[g_{\epsilon}\diamond\mathcal{L}_{t-s,t}\leq-f]=\mathbb{P}[g\diamond\mathcal{L}_{t-s,t}\leq-f]=\mathbb{P}[f\diamond\mathcal{L}_{t-s,t}\leq-g].$
Therefore we conclude that $\mathbb{P}[\mathcal{M}_{t-s,t}\diamond f\leq-g]\geq\mathbb{P}[f\diamond\mathcal{L}_{t-s,t}\leq-g]$. In other words, $\mathfrak{h}_{s}f|_{\mathbb{Q}}\stackrel{{\scriptstyle d}}{{=}}(f\diamond\mathcal{L}_{t-s,t})|_{\mathbb{Q}}$ stochastically dominates $(\mathcal{M}_{t-s,t}\diamond f)|_{\mathbb{Q}}$. To prove the lemma from here, it suffices to observe that for a single point $q\in\mathbb{Q}$:
$\mathbb{P}\left[\mathcal{M}_{t-s,t}\diamond f(q)\leq a\right]=\mathbb{P}[\mathcal{M}_{t-s,t}(q,\cdot)\leq-f+a]=\mathbb{P}[\mathcal{L}_{t-s,t}(q,\cdot)\leq-f+a]=\mathbb{P}\left[f\diamond\mathcal{L}_{t-s,t}(q)\leq a\right],$
where the middle equality uses condition $3$ of Theorem 1.1. ∎
Note that the method for evaluating the infimum in (17) above by approximating deterministic initial conditions using random functions bears similarity to ideas from [QS23, Section 7] and [ACH24a, Proposition 5.1].
Lemma 3.3.
Any pre-landscape $\mathcal{M}$ satisfies the metric composition law on $\mathbb{Q}$: for all $s<r<t\in\mathbb{Q}$ we have $\mathcal{M}_{s,r}\diamond_{\mathbb{Q}}\mathcal{M}_{r,t}=\mathcal{M}_{s,t}$.
Proof.
Since $\mathbb{Q}$ is countable, it suffices to prove that for fixed $s<r<t\in\mathbb{Q}$ and $x,y\in\mathbb{Q}$ we have $\mathcal{M}_{s,r}\diamond_{\mathbb{Q}}\mathcal{M}_{r,t}(x,y)=\mathcal{M}_{s,t}(x,y)$. By the triangle inequality we have $\mathcal{M}_{s,r}\diamond_{\mathbb{Q}}\mathcal{M}_{r,t}(x,y)\leq\mathcal{M}_{s,t}(x,y)$ almost surely so it suffices to prove that equality holds in distribution. For this, observe that in the directed landscape $\mathcal{L}$ we have $\mathcal{L}_{s,r}\diamond_{\mathbb{Q}}\mathcal{L}_{r,t}(x,y)=\mathcal{L}_{s,t}(x,y)$ by the usual metric composition law for $\mathcal{L}$, continuity, and density of $\mathbb{Q}$. On the other hand, $\mathcal{L}_{s,t}(x,y)\stackrel{{\scriptstyle d}}{{=}}\mathcal{M}_{s,t}(x,y)$ (by condition $3$ of Theorem 1.1), and
$(\mathcal{L}_{s,r}(x,z),\mathcal{L}_{r,t}(z,y))_{z\in\mathbb{Q}}\stackrel{{\scriptstyle d}}{{=}}(\mathcal{M}_{s,r}(x,z),\mathcal{M}_{r,t}(z,y))_{z\in\mathbb{Q}}.$ (18)
This uses the independent increment property of $\mathcal{L},\mathcal{M}$, the equality $(\mathcal{L}_{s,r}(x,z))_{z\in\mathbb{Q}}\stackrel{{\scriptstyle d}}{{=}}(\mathcal{M}_{s,r}(x,z))_{z\in\mathbb{Q}}$ from condition $3$ of Theorem 1.1, and the equality $(\mathcal{L}_{r,t}(z,y))_{z\in\mathbb{Q}}\stackrel{{\scriptstyle d}}{{=}}(\mathcal{M}_{s,r}(z,y))_{z\in\mathbb{Q}}$ from Lemma 3.2. Therefore $\mathcal{M}_{s,r}\diamond_{\mathbb{Q}}\mathcal{M}_{r,t}(x,y)\stackrel{{\scriptstyle d}}{{=}}\mathcal{M}_{s,t}(x,y)$, as desired. ∎
Lemma 3.4.
Let $u_{1}=(x_{1},s_{1};y_{1},t_{1}),u_{2}=(x_{2},s_{2};y_{2},t_{2})\in\mathbb{Q}^{4}_{\uparrow}$, and $\hat{x}_{2},\hat{y}_{2}\in\mathbb{R}$ be chosen so that the four points $(x_{2},s_{2})$, $(\hat{x}_{2},s_{1})$, $(y_{2},t_{2})$, $(\hat{y}_{2},t_{1})$ are on the same straight line. Denote $\chi=|x_{1}-\hat{x}_{2}|+|y_{1}-\hat{y}_{2}|$, and $\tau=|t_{1}-t_{2}|+|s_{1}-s_{2}|$. Then letting $d(x,s;y,t)=(x-y)^{2}/(t-s)$, for any $M,L\geq 0$, we have
$\mathbb{P}[|[\mathcal{M}+d](u_{1})-[\mathcal{M}+d](u_{2})|>M\tau^{1/3}+L\chi^{1/2}]<ce^{-dM^{3/2}}+ce^{-dL^{2}},$
where $c,d>0$ are absolute constants.
Proof.
It suffices to prove this estimate assuming that precisely one of $x_{1}-x_{2}$, $s_{1}-s_{2}$, $y_{1}-y_{2}$, $t_{1}-t_{2}$ is non-zero. We claim that in all these cases we have
$(\mathcal{M}(u_{1}),\mathcal{M}(u_{2}))\stackrel{{\scriptstyle d}}{{=}}(\mathcal{L}(u_{1}),\mathcal{L}(u_{2})),$ (19)
in which case the conclusion follows from [DOV22, Lemma 10.4]. In the case when $y_{1}\neq y_{2}$ or $t_{1}\neq t_{2}$ this is immediate from condition $3$ of Theorem 1.1; and when $x_{1}\neq x_{2}$ this follows from Lemma 3.2. Now suppose $s_{1}\neq s_{2}$. By symmetry we may assume $s_{1}<s_{2}$. Using Lemma 3.3 we have
$\mathcal{M}(u_{1})=\mathcal{M}_{s_{1},s_{2}}\diamond_{\mathbb{Q}}\mathcal{M}_{s_{2},t_{1}}(x_{1},y_{1}),\qquad\mathcal{M}(u_{2})=\mathcal{M}_{s_{2},t_{1}}(x_{1},y_{1}).$
The same equalities hold with $\mathcal{L}$ in place of $\mathcal{M}$, and so (19) follows since the joint law of $\mathcal{M}_{s_{1},s_{2}}(x_{1},\cdot)|_{\mathbb{Q}}$, $\mathcal{M}_{s_{2},t_{1}}(\cdot,y_{1})|_{\mathbb{Q}})$ is the same as in $\mathcal{L}$ (this is an instance of (18)). ∎
Corollary 3.5.
Any pre-landscape $\mathcal{M}$ has a continuous extension which is locally Hölder-$1/2-$ in $x,y$, and locally Hölder-$1/3-$ in $s,t$.
Proof.
This is immediate from Lemma 3.4 and the Kolmogorov-Centsov criterion. ∎
Corollary 3.6.
Proposition 2.5 holds verbatim with a pre-landscape $\mathcal{M}$ in place of $\mathcal{L}$.
Proof.
The proof of Proposition 2.5 in [DOV22] only uses one-point and two-point tail bounds for $\mathcal{L}$, which are unchanged for $\mathcal{M}$ (the one-point marginals are equal and the two-point tail bound is given by Lemma 3.4). ∎
Proof of Proposition 3.1.
Corollary 3.5 provides the continuous extension of $\mathcal{M}$. Next, we can interpret Lemma 3.2 as saying that for any $f,g\in\operatorname{UC}_{0}$ supported inside $\mathbb{Q}$, we have
$\mathbb{P}[g\diamond\mathcal{L}_{t-s,t}\diamond f\leq 0]=\mathbb{P}[g\diamond\mathcal{M}_{t-s,t}\diamond f\leq 0],$
Since both $\mathcal{M},\mathcal{L}$ are continuous, by an approximation arguments this identity holds for arbitrary $f,g\in\operatorname{UC}_{0}$, giving (16) in full generality.
To prove the metric composition law, observe that continuity of $\mathcal{M}$ implies that for fixed $s<r<t\in\mathbb{Q}$ and $x,y\in\mathbb{Q}$ we have $\mathcal{M}_{s,r}\diamond_{\mathbb{Q}}\mathcal{M}_{r,t}(x,y)=\mathcal{M}_{s,r}\diamond\mathcal{M}_{r,t}(x,y)$. Therefore Lemma 3.3 shows that
$\mathcal{M}_{s,t}=\mathcal{M}_{s,r}\diamond\mathcal{M}_{r,t}$ (20)
on $\mathbb{Q}^{2}$. Using the continuity of $\mathcal{M}$, and Corollary 3.6 which controls the location of the argmax in the metric composition, we can deduce that both sides of (20) are continuous, and are also continuous in $s,r$, and $t$. Thus we can extend (20) to $\mathbb{R}^{2}$, and to hold for any $s<r<t$ simultaneously.
3.2 The law of $\mathcal{M}_{0,1}$ and an exchange strategy
By the metric composition law in 3.1 and independence of increments, via Definition 2.1, to prove Theorem 1.1 it suffices to show that $\mathcal{M}_{s,t}\stackrel{{\scriptstyle d}}{{=}}\mathcal{L}_{s,t}$ is a (rescaled) Airy sheet for any $s<t$. Without loss of generality, we prove this for $s=0$ and $t=1$. By continuity of $\mathcal{M}$, it suffices to prove the following.
Proposition 3.7.
Take any $k\in\mathbb{N}$, and $x_{1},\ldots,x_{k}\in\mathbb{R}$, and $y_{1},\ldots,y_{k}\in\mathbb{R}$. Then
$\{\mathcal{M}_{0,1}(x_{i},y_{i})\}_{i=1}^{k}\stackrel{{\scriptstyle d}}{{=}}\{\mathcal{L}_{0,1}(x_{i},y_{i})\}_{i=1}^{k}.$
The rest of this section is devoted to proving 3.7. We shall then use $C,c>0$ to denote large and small constants, whose values may change from line to line.
We will first prove a stationary version of 3.7. Take $\mathcal{R}=(\mathcal{R}_{1},\ldots,\mathcal{R}_{k})$ as in Theorem 2.9. Below the choice of slopes $a_{1}<\cdots<a_{k}$ does not matter, e.g., we can simply take each $a_{i}=i$.
Proposition 3.8.
Take any $k\in\mathbb{N}$, $y_{1},\ldots,y_{k}\in\mathbb{R}$, and $\mathcal{R}$ as above, independent of $\mathcal{L},\mathcal{M}$. Then,
(,$(\mathcal{R},\{\mathcal{R}_{i}\diamond\mathcal{M}_{0,1}(y_{i})\}_{i=1}^{k})\stackrel{{\scriptstyle d}}{{=}}(\mathcal{R},\{\mathcal{R}_{i}\diamond\mathcal{L}_{0,1}(y_{i})\}_{i=1}^{k}).$ (21)
The proof of Proposition 3.8 uses the strategy of the Lindeberg exchange method outlined in the introduction. To setup, for each $t\in[0,1]$, we denote [^1] Here and below, we use the convention that $\mathcal{L}(x,t;x,t)=\mathcal{M}(x,t;x,t)=0$ for any $x,t$, and $\mathcal{L}(x,t;y,t)=\mathcal{M}(x,t;y,t)=-\infty$ for any $t$ and $x\neq y$.
$\mathcal{M}^{t}_{i}=\mathcal{R}_{i}\diamond\mathcal{L}_{0,t}\diamond\mathcal{M}_{t,1}(y_{i}),\qquad\mathcal{M}^{t}=(\mathcal{M}^{t}_{1},\dots,\mathcal{M}^{t}_{k}),$ (22)
where $\mathcal{R},\mathcal{L},\mathcal{M}$ are all independent. The vectors $\mathcal{M}^{t},t\in[0,1]$ interpolate between the two sides of the equality in law in Proposition 3.8.
Our basic strategy will be to show that for $0\leq s<t\leq 1$ we can define a new coupling of the random variables $\mathcal{R},\mathcal{M}^{t}$ and $\mathcal{M}^{s}$ so that $\mathbb{E}\|\mathcal{M}^{t}-\mathcal{M}^{s}\|_{1}=o(t-s)$ for $t$ close to $s$. Note that this will typically not be true in the coupling given through (22) where $\mathcal{L},\mathcal{M}$, and $\mathcal{R}$ are all independent. The next proposition states this more precisely.
Below we allow all the constants $C,c>0$ to depend on $a_{1}<\cdots<a_{k}$ and $y_{1},\ldots,y_{k}$.
Proposition 3.9.
For each $\theta\in(0,1/4)$, there exists $\delta>0$ such that the following holds. For $s<t\in[\theta,1-\theta]$, we can find a coupling of $\mathcal{R},\mathcal{L},\mathcal{M}$ such that each of the pairs $(\mathcal{R},\mathcal{M}^{t})$ and $(\mathcal{R},\mathcal{M}^{s})$ have the same law as in (22), and
$\mathbb{P}[\mathcal{M}^{s}\neq\mathcal{M}^{t}]\leq C(t-s)^{2/3+\delta},$
where $C>0$ may also depend on $\theta$.
The proof of this proposition will be given shortly; we first finish the proof of 3.8 given Proposition 3.9. We will also need a tail estimate for the difference $\mathcal{M}^{s}_{i}-\mathcal{M}^{t}_{i}$ to convert the total variation bound in Proposition 3.9 to a bound on a Wasserstein distance.
Lemma 3.10.
For any $0\leq s<t\leq 1$, under any coupling between $\mathcal{M},\mathcal{L},$ and $\mathcal{R}$, we have the following: for any $i\in\llbracket 1,k\rrbracket$ and $M>0$,
$\mathbb{P}\left[|\mathcal{M}^{s}_{i}-\mathcal{M}^{t}_{i}|>M(1-\log(t-s))^{2}(t-s)^{1/3}\right]<C\exp(-cM).$
Proof.
We can assume that $M$ is large enough, since otherwise the conclusion is trivial. We first prove that
$\mathbb{P}\left[\mathcal{M}^{s}_{i}<\mathcal{M}^{t}_{i}-M(1-\log(t-s))^{2}(t-s)^{1/3}\right]<C\exp(-cM).$ (23)
By 2.5 and Corollary 3.6, with probability at least $1-C\exp(-cM)$, the bound in 2.5 holds for both $\mathcal{M},\mathcal{L}$ with some $R<M^{2/3}$. Also, with probability $>1-C\exp(-cM)$, we have $|\mathcal{R}_{i}(x)-\mathcal{R}_{i}(0)-a_{i}x|<|x|+M$ for all $x\in\mathbb{R},i=1,\dots,k$. Let $\mathcal{E}$ denote the intersection of the above two events. Then on $\mathcal{E}$, we have that
$\mathcal{R}_{i}(x)+\mathcal{L}_{0,s}(x,w)+\mathcal{L}_{s,t}(w,z)+\mathcal{M}_{t,1}(z,y_{i})<\mathcal{R}_{i}(0)+\mathcal{L}_{0,s}(0,0)+\mathcal{L}_{s,t}(0,y_{i})+\mathcal{M}_{t,1}(y_{i},1)\leq\mathcal{M}^{t}_{i},$
for any $x,w,z\in\mathbb{R}$ with $|x|\vee|w|\vee|z|>CM^{1/2}$. Then by the continuity of $\mathcal{M},\mathcal{L},\mathcal{R}$, there exist $x_{*},w_{*},z_{*}$ with $|x_{*}|\vee|w_{*}|\vee|z_{*}|\leq CM^{1/2}$, such that
$\mathcal{M}^{t}_{i}=\mathcal{R}_{i}(x_{*})+\mathcal{L}_{0,s}(x_{*},w_{*})+\mathcal{L}_{s,t}(w_{*},z_{*})+\mathcal{M}_{t,1}(z_{*},y_{i}).$
Also, on $\mathcal{E}$ we have
$\mathcal{M}_{s,t}(w_{*},z_{*})-\mathcal{L}_{s,t}(w_{*},z_{*})\geq-CM^{2/3}(t-s)^{1/3}(\log(M)+|\log(t-s)|)^{2}.$
On the other hand, we have
$\mathcal{M}^{s}_{i}\geq\mathcal{R}_{i}(x_{*})+\mathcal{L}_{0,s}(x_{*},w_{*})+\mathcal{M}_{s,t}(w_{*},z_{*})+\mathcal{M}_{t,1}(z_{*},y_{i}).$
These imply that on $\mathcal{E}$, $\mathcal{M}^{s}_{i}>\mathcal{M}^{t}_{i}-CM^{2/3}(t-s)^{1/3}(\log(M)+|\log(t-s)|)^{2}$, yielding (23). We can similarly prove this estimate with the roles of $\mathcal{M}^{s}_{i}$ and $\mathcal{M}^{t}_{i}$ switched, so the conclusion follows. ∎
Proof of 3.8.
It suffices to show that the equality in law holds with $\mathcal{R}$ replaced by $\mathcal{R}|_{F}$ where $F\subset\mathbb{R}$ is a finite subset of the domain of $\mathcal{R}:\mathbb{R}\to\mathbb{R}^{n}$. We think of $\mathcal{R}|_{F}$ as a random variable in $\mathbb{R}^{n|F|}$. Let $d_{W}(X,Y)=\inf_{\mu}\mathbb{E}_{\mu}\|X-Y\|_{1}$ be the $L^{1}$-Wasserstein distance between the laws of two $\mathbb{R}^{n}$-valued random variables $X$ and $Y$. Here the infimum is over all couplings of the random variables $X,Y$. Fix $\theta\in(0,1/4)$. By 3.9 and Lemma 3.10, for $s<t\in[\theta,1-\theta]$ in the coupling of $\mathcal{M}^{s},\mathcal{M}^{t}$ from 3.9 we have
$\mathbb{E}\|(\mathcal{R}|_{F},\mathcal{M}_{s})-(\mathcal{R}|_{F},\mathcal{M}_{t})\|_{1}=\mathbb{E}\|\mathcal{M}_{s}-\mathcal{M}_{t}\|_{1}<C(t-s)^{1+\delta/2}.$
This then bounds $d_{W}((\mathcal{R}|_{F},\mathcal{M}_{s}),(\mathcal{R}|_{F},\mathcal{M}_{t}))$. Also, Lemma 3.10 alone implies that in any coupling.
$\mathbb{E}|\mathcal{M}_{0}-\mathcal{M}_{\theta}|+\mathbb{E}|\mathcal{M}_{1}-\mathcal{M}_{1-\theta}|<C(\log(\theta))^{2}\theta^{1/3}.$
Therefore by the triangle inequality Wasserstein distance, we have
$d_{W}((\mathcal{R}|_{F},\mathcal{M}_{0}),(\mathcal{R}|_{F},\mathcal{M}_{1}))$ $<C(\log(\theta))^{2}\theta^{1/3}+\sum_{i=1}^{n}d_{W}((\mathcal{R}|_{F},\mathcal{M}_{\theta+(1-2\theta)(i-1)/n}),(\mathcal{R}|_{F},\mathcal{M}_{\theta+(1-2\theta)i/n}))$
$\leq C(\log(\theta))^{2}\theta^{1/3}+Cn^{-\delta/2}(1-2\theta)^{1+\delta/2}$
Taking $n\to\infty$ and then $\theta\to 0$ gives that $d_{W}((\mathcal{R}|_{F},\mathcal{M}_{0}),(\mathcal{R}|_{F},\mathcal{M}_{1}))=0$, and so $(\mathcal{R}|_{F},\mathcal{M}_{0})\stackrel{{\scriptstyle d}}{{=}}(\mathcal{R}|_{F},\mathcal{M}_{1})$. ∎
3.3 The coupling
We now define the coupling to be used in 3.9. Below we fix $s<t=s+r\in[\theta,1-\theta]$, and throughout assume $r=t-s$ is sufficiently small. We start with independent copies of $\mathcal{R}$, $\mathcal{L}_{0,s}$, and $\mathcal{M}_{t,1}$, and consider the following subsets of $\mathbb{R}$. For each $i\in\llbracket 1,k\rrbracket$, denote $H_{i}=\mathcal{R}_{i}\diamond\mathcal{L}_{0,s},$ and
$\tilde{A}_{i}=\{x\in\mathbb{R}:H_{i}(x)+\mathcal{M}_{t,1}(x,y_{i})>H_{i}\diamond\mathcal{M}_{t,1}(y_{i})-\log^{10}(r)r^{1/3}\}.$
We let $A_{i}$ be the $\log^{10}(r)r^{2/3}$-neighborhood of $\tilde{A}_{i}$. For each $i,j\in\llbracket 1,k\rrbracket$, $i\neq j$, we let $\tilde{D}_{i,j}$ be the set where $H_{i}-H_{j}$ is non-constant, i.e.,
$\tilde{D}_{i,j}=\{x\in\mathbb{R}:\text{ for all }\epsilon>0,\text{ there exists }y\in(x-\epsilon,x+\epsilon)\text{ with }H_{i}(x)-H_{j}(x)\neq H_{i}(y)-H_{j}(y)\}.$
We then let $D_{i,j}$ be the $\log^{10}(r)r^{2/3}$ neighborhood of $\tilde{D}_{i,j}$. The following estimate is the key to our coupling.
Proposition 3.11.
There is a constant $\delta>0$, such that for all sufficiently small $r>0$ and $i,j\in\llbracket 1,k\rrbracket$, $i\neq j$, we have $\mathbb{P}[A_{i}\cap A_{j}\cap D_{i,j}=\emptyset]>1-r^{2/3+\delta}$.
The proof of this estimate relies on careful analysis of the joint law of $\mathcal{R}_{i}$ and $\mathcal{R}_{j}$, together with Property $\ast$ from Remark 1.2. We postpone it to Section 4, and continue our construction of the coupling.
Let $\mathcal{E}_{*}$ be the event where $A_{i}\cap A_{j}\cap D_{i,j}=\emptyset$, for each $i,j\in\llbracket 1,k\rrbracket$, $i\neq j$. Note that for each $i\in\llbracket 1,k\rrbracket$, (using the shape bounds in 2.5, Corollary 3.6 on $\mathcal{L},\mathcal{M}$) $A_{i}$ consists of finitely many open intervals $A_{i,1},\dots,A_{i,\ell(i)}$, each of length at least $2\log^{10}(r)r^{2/3}$.
Lemma 3.12.
There exists a function $F:\mathbb{R}\to\mathbb{R}\cup\{-\infty\}$ that is $\sigma(\mathcal{R},\mathcal{L}_{0,s},\mathcal{M}_{t,1})$-measurable, and satisfies the following conditions:
  • On $\mathcal{E}_{*}$, for each $i\in\llbracket 1,k\rrbracket$, $F-H_{i}$ is constant on each interval $A_{i,j}$, $j=1,\dots,\ell(i)$.
  • $F=-\infty$ outside of $\bigcup_{i=1}^{k}A_{i}$,
  • For any $M>0$, with probability at least $1-C\exp(-cM)$ the following is true: $\mathcal{E}_{*}$ holds, and $|F(x)-F(y)|\leq M\log(|x-y|^{-1}+2)\sqrt{|x-y|}$ for all $x,y\in\bigcup_{i=1}^{k}A_{i}$, $|x-y|\leq\log^{10}(r)r^{2/3}$.
Proof.
On $\mathcal{E}_{*}^{c}$, set $F=0$ on $\bigcup_{i=1}^{k}A_{i}$ and $\infty$- ∞ off this set. Next, for any interval $A_{i,j}$ and a function $f:\mathbb{R}\to\mathbb{R}\cup\{-\infty\}$ , we say that $f$ is suitable on $A_{i,j}$, if $f-H_{i}$ is constant on $I$. Sort the intervals $A_{i,j},i=1,\dots,k,j=1,\dots,\ell(i)$ into a list $I_{1},\ldots,I_{n}$ so that the $I_{j}$ are ordered by their left endpoint (ordering arbitrarily if left endpoints agree). We define $F$ on $I_{1},\dots,I_{n}$ inductively.
For $I_{1}$, choosing $i$ so that $I_{1}=A_{i,j}$ for some $j$, and we let $F=H_{i}$ on $I_{1}$. Now assume that $F$ is already defined on $\bigcup_{\iota=1}^{\ell}I_{\iota}$ for some $\ell\in\llbracket 1,n-1\rrbracket$, and for each $\iota\in\llbracket 1,\ell\rrbracket$, $F$ is suitable on $I_{\iota}$. We then define $F$ on $I_{\ell+1}$, as follows. If $I_{\ell+1}$ intersects $\bigcup_{\iota=1}^{\ell}I_{\iota}$, there is a unique way of defining $F$ to make it suitable on $I_{\ell+1}$. Here we crucially use that under $\mathcal{E}_{*}$, for each $i<i^{\prime}$ and $j,j^{\prime}$, we have that $H_{i}-H_{i^{\prime}}$ is constant on $A_{i,j}\cap A_{i^{\prime},j^{\prime}}$, since $A_{i}\cap A_{j}$ is disjoint from $D_{i,j}$. Otherwise, we define $F$ to be suitable on $I_{\ell+1}$, so the right limit of $F$ at the left endpoint of $I_{\ell+1}$ equals the left limit of $F$ at the right endpoint of $\bigcup_{\iota=1}^{\ell}I_{\iota}$. Take $F=-\infty$ outside $\bigcup_{i=1}^{k}A_{i}$.
It remains to check the last condition. By Theorem 2.9 the law of each $H_{i}$ is a Brownian motion with drift $a_{i}$. Then for any $M>1$, with probability $>1-C\exp(-cM)$ we have that $|H_{i}(x)-H_{i}(y)|<M\log(|x-y|^{-1}+2)\sqrt{|x-y|}$ for each $i\in\llbracket 1,k\rrbracket$ and $x,y\in[-M,M]$. Also, by 2.5/Corollary 3.6, with probability $>1-C\exp(-cM)$, we have that $\bigcup_{i=1}^{k}A_{i}\subset[-M,M]$. Thus the last condition holds from the construction of $F$. The measurability is immediate from construction. ∎
Now, observe that if $\mathcal{M},\mathcal{L}$ and $\mathcal{R}$ are all independent, then conditional on $X:=(\mathcal{R},\mathcal{L}_{0,s},\mathcal{M}_{t,1})$ we have the equality in law
$F\diamond\mathcal{M}_{s,t}\stackrel{{\scriptstyle d}}{{=}}F\diamond\mathcal{L}_{s,t}.$ (24)
Indeed, by the independent increment property of $\mathcal{L},\mathcal{M}$, both $\mathcal{M}_{s,t},\mathcal{L}_{s,t}$ are independent of $X,F$, and so by 3.1, both sides of this equality are equal in law to $\mathfrak{h}_{r}F$. Therefore conditional on $X$, we can couple $\mathcal{M}_{s,t},\mathcal{L}_{s,t}$ so that this equality is realized almost surely. Observe that in this coupling, the pairs $(X,\mathcal{M}_{s,t})$ and $(X,\mathcal{L}_{s,t})$ are still respectively independent; it is only in the triple $(X,\mathcal{M}_{s,t},\mathcal{L}_{s,t})$ where we lose the independence. In particular, the individual laws of $(\mathcal{R},\mathcal{M}^{s})$ and $(\mathcal{R},\mathcal{M}^{t})$ (defined by (22)) are still the same as in the original independent coupling.
We use this new coupling to prove 3.9.
Proof of 3.9.
By Proposition 3.11, it is enough to show that
$\mathbb{P}\left[\mathcal{M}^{s}\neq\mathcal{M}^{t},\mathcal{E}_{*}\right]<C\exp(-c\log^{2}(r)).$ (25)
We let $\mathcal{E}$ be the event where $\mathcal{E}_{*}$ holds, and additionally:
  • $|F(x)-F(y)|<\log^{2}(r)\log(|x-y|^{-1}+2)\sqrt{|x-y|}$ for any $x,y\in\bigcup_{i=1}^{k}A_{i}$, $|x-y|\leq\log^{10}(r)r^{2/3}$.
  • $|H_{i}(x)-H_{i}(y)|<\log^{2}(r)\log(|x-y|^{-1}+2)\sqrt{|x-y|}$ for any $i\in\llbracket 1,k\rrbracket$ and $x,y\in[-\log^{2}(r),\log^{2}(r)]$.
  • $R<\log^{2}(r)$, for the $R$ in 2.5/Corollary 3.6 defined for both $\mathcal{L}$ and $\mathcal{M}$.
  • $|H_{i}(x)-H_{i}(0)|<C|x|+\log^{2}(r)$, for any $i\in\llbracket 1,k\rrbracket$ and $x\in\mathbb{R}$.
By Lemma 3.12, standard Brownian motion estimates, and 2.5/Corollary 3.6,
$\mathbb{P}[\mathcal{E}_{*}\setminus\mathcal{E}]<C\exp(-c\log^{2}(r)).$
It now remains to show that $\mathcal{M}^{s}=\mathcal{M}^{t}$ on $\mathcal{E}$.
Let $\hat{A}_{i}$ be the $\log^{10}(r)r^{2/3}/2$ neighborhood of $\tilde{A}_{i}$. We next show that on $\mathcal{E}$,
$H_{i}\diamond\mathcal{M}_{s,t}|_{\hat{A}_{i}}=H_{i}\diamond\mathcal{L}_{s,t}|_{\hat{A}_{i}}.$ (26)
First observe that on $\mathcal{E}$, the third bullet point (for $\mathcal{M}$) and the fourth bullet point above guarantee
$\bigcup_{i=1}^{k}A_{i}\subset[-\log^{2}(r)/2,\log^{2}(r)/2].$ (27)
Now take any $z\in[-\log^{2}(r)/2,\log^{2}(r)/2]$. On $\mathcal{E}$, for both of the functions
$F+\mathcal{M}_{s,t}(\cdot,z),\quad H_{i}+\mathcal{M}_{s,t}(\cdot,z),$
any argmax must be in the interval $I_{z}:=(z-\log^{10}(r)r^{2/3}/2,z+\log^{10}(r)r^{2/3}/2)$. This follows by invoking the first and third bullet point to bound the sum with $F$, and the second, third, and fourth bullet point to bound the sum with $H_{i}$, and using (27) to ensure that we can apply the bound in the second bullet point near $z$. Then for each $i\in\llbracket 1,k\rrbracket$, since $F-H_{i}$ is constant on $I_{z}$ when $z\in\hat{A}_{i}$ (by the construction of $F$), we must have that
$F\diamond\mathcal{M}_{s,t}(z)-H_{i}\diamond\mathcal{M}_{s,t}(z)=F(z)-H_{i}(z).$
This equality with $\mathcal{L}_{s,t}$ in place of $\mathcal{M}_{s,t}$ holds by the same reasoning. Hence (26) follows from (24).
Next, let $z_{*}=\operatorname*{arg\,max}_{z}H_{i}(z)+\mathcal{M}_{t,1}(z,y_{i})$ and let $Y_{i}=H_{i}\diamond\mathcal{M}_{t,1}(y_{i})$. On $\mathcal{E}$, (27) gives that $|z_{*}|\leq\log^{2}(r)$, and so by the third bullet point of $\mathcal{E}$ we have the bound
$H_{i}(z_{*})+\mathcal{M}_{s,t}(z_{*},z_{*})+\mathcal{M}_{t,1}(z_{*},y_{i})\geq Y_{i}-\log^{4}(r)r^{1/3}.$
On the other hand, for $z\notin\hat{A}_{i}$ and $x\in\mathbb{R}$ with $|x|,|z|\leq\log^{2}(r)$ we must have
$H_{i}(x)+\mathcal{M}_{s,t}(x,z)+\mathcal{M}_{t,1}(z,y_{i})$
[…]
Here the final inequality uses the second and third bullet points to bound $|H_{i}(x)-H_{i}(z)|+\mathcal{M}_{s,t}(x,z)$, and the definition of $\tilde{A}_{i}$ to bound $H_{i}(z)+\mathcal{M}_{t,1}(z,y_{i})$. An identical bound holds with $\mathcal{L}_{s,t}(x,z)$ in place of $\mathcal{M}_{s,t}(x,z)$. The same bound holds when either $|x|>\log^{2}(r)$ or $|z|>\log^{2}(r)$ by the shape bound from the third and fourth bullet points.
Combining the previous two displays, we see that in the metric composition laws
$\mathcal{M}^{s}_{i}$ $=\max_{x,z}H_{i}(x)+\mathcal{M}_{s,t}(x,z)+\mathcal{M}_{t,1}(z,y_{i}),$
$\mathcal{M}^{t}_{i}$ $=\max_{x,z}H_{i}(x)+\mathcal{L}_{s,t}(x,z)+\mathcal{M}_{t,1}(z,y_{i}),$
the maximum is always achieved at some $z\in\hat{A}_{i}$. Therefore by (26) we have $\mathcal{M}^{s}_{i}=\mathcal{M}^{t}_{i}$ on $\mathcal{E}$, as desired. ∎
3.4 Removing the stationary initial condition
In this subsection we deduce 3.7 from 3.8.
Proof of 3.7.
By the continuity of both $\mathcal{M}$ and $\mathcal{L}$, it suffices to consider the case when $x_{1},\ldots,x_{k}$ are all distinct. Without loss of generality, we further assume that $x_{1}<\cdots<x_{k}$.
We consider $\mathcal{R}=(\mathcal{R}_{1},\dots,\mathcal{R}_{k})$ constructed as in Theorem 2.9 from independent Brownian motions $\mathcal{B}=(\mathcal{B}_{1},\dots,\mathcal{B}_{k})$, with slopes $1,2,\dots,k$.
Now, since Proposition 3.8 makes a claim about the joint laws of $(\mathcal{R},\mathcal{R}\diamond\mathcal{M}_{0,1})$ and $(\mathcal{R},\mathcal{R}\diamond\mathcal{L}_{0,1})$, for any random function $\mathcal{Q}:\mathbb{R}\to\mathbb{R}^{n}$ whose law $\mu_{\mathcal{Q}}$ is absolutely continuous with respect to the law $\mu_{\mathcal{R}}$ of $\mathcal{R}$, we have that (21) holds with $\mathcal{Q}$ in place of $\mathcal{R}$ (where $\mathcal{Q}$ is independent of $\mathcal{L},\mathcal{M}$). With this in mind, we will construct a sequence of functions $\mathcal{R}^{\epsilon},\epsilon>0$ with $\mu_{\mathcal{R}^{\epsilon}}\ll\mu_{\mathcal{R}}$ such that as $\epsilon\to 0$,
$\sup_{x:|x-x_{i}|>\epsilon}\mathcal{R}_{i}^{\epsilon}(x)-2k|x|-\mathcal{R}_{i}^{\epsilon}(x_{i})\stackrel{{\scriptstyle\mathbb{P}}}{{\to}}-\infty,$ (28)
and
$\sup_{x:|x-x_{i}|\leq\epsilon}\mathcal{R}_{i}^{\epsilon}(x)-\mathcal{R}_{i}^{\epsilon}(x_{i})\stackrel{{\scriptstyle\mathbb{P}}}{{\to}}0.$ (29)
Then by (21) for $\mathcal{R}^{\epsilon}$, continuity of $\mathcal{M}$ and $\mathcal{L}$, and the tail bounds in 2.5/Corollary 3.6 on $\mathcal{M},\mathcal{L}$, the conclusion follows.
We now give the construction of $\mathcal{R}^{\epsilon}$. Recall that $\mathcal{R}=f(\mathcal{B})$ where $f$ is given by (15). Therefore if $\phi:\mathbb{R}\to\mathbb{R}^{n}$ is any deterministic function with $\phi(0)=0,\int\|\phi^{\prime}\|_{2}^{2}<\infty$, then by the Cameron-Martin theorem, $f(\mathcal{B}+\phi)$ has an absolutely continuous law with respect to $\mathcal{R}$. With this in mind, define $\phi_{\epsilon}$ to be the unique continuous function with $\phi_{\epsilon}(0)=(0,\dots,0)$ such that for all $i$:
  • $\phi_{\epsilon,i}^{\prime}(x)=0$ for $x<x_{i}-\epsilon,x>x_{k}+1$.
  • $\phi_{\epsilon,i}^{\prime}(x)=i\epsilon^{-2}$ for $x\in(x_{i}-\epsilon,x_{i})$.
  • $\phi_{\epsilon,i}^{\prime}(x)=-2i\epsilon^{-2}$ for $x\in(x_{i},x_{k}+1)$.
We set $\mathcal{R}^{\epsilon}=f(\mathcal{B}+\phi_{\epsilon})$ and check that (28), (29) hold. First, for all $x,i$, we have the inequalities
$\phi_{\epsilon}[i\to x]+|\pi^{i,x}|_{\mathcal{B}}\leq(\mathcal{B}+\phi_{\epsilon})[i\to x]\leq\mathcal{B}[i\to x]+\phi_{\epsilon}[i\to x]$
where $\pi^{i,x}$ is a vector achieving the supremum of $\phi_{\epsilon}[i\to x]$. Recognizing that $\pi^{i,x_{i}}=(x_{i},\dots,x_{i})$, these bounds imply that
$\mathcal{R}_{i}^{\epsilon}(x)-2k|x|-\mathcal{R}_{i}^{\epsilon}(x_{i})$ $\leq\phi_{\epsilon}[i\to x]-2k|x|-\phi_{\epsilon}[i\to x_{i}]+\tilde{\mathcal{R}}_{i}(x)-\mathcal{B}_{i}(x_{i})$
$=(\phi_{\epsilon}[i\to x]-\phi_{\epsilon}[i\to x_{i}])+(\mathcal{R}_{i}(x)-2k|x|+\tilde{\mathcal{R}}_{i}(0)-\mathcal{B}_{i}(x_{i})).$
As $\epsilon\to 0$, the second bracketed term above remains bounded even after we take a supremum over $x$, since $\mathcal{R}_{i}$ is a Brownian motion with drift $i\leq k$; while by straightforward algebra the first term goes to $\infty$- ∞ uniformly over $x$ with $|x-x_{i}|>\epsilon$. This yields (28). For the bound (29), observe that for any $f:\mathbb{R}\to\mathbb{R}^{n}$ and $x<x_{i}<y$ we have
$f[i\to x]-f[i\to x_{i}]$ $\leq f_{1}(x)-f_{1}(x_{i})$
$f[i\to y]-f[i\to x_{i}]$ $\leq\sum_{j=1}^{i}\sup_{z<z^{\prime}\in[x_{i},y]}f_{j}(z)-f_{j}(z^{\prime}).$
Using these inequalities with $f=\mathcal{B}+\phi_{\epsilon}$ yields (29) since $\mathcal{B}$ is continuous and each $\phi_{\epsilon,j}$ is non-decreasing on $[x_{i}-\epsilon,x_{i}]$ and non-increasing on $[x_{i},x_{i}+\epsilon]$. ∎
4 Branching point estimation: proof of Proposition 3.11
In this section, we prove 3.11, the final piece of Theorem 1.1. Without loss of generality we prove the bound for $i=1$ and $j=2$. In this section we again use $C,c>0$ to denote large and small constants, whose values may change from line to line, and they are allowed to depend on $\theta,y_{1},y_{2}$ and the slopes $a_{1},a_{2}$.
We start with a simple transversal fluctuation estimate.
Lemma 4.1.
We have $\mathbb{P}[A_{1}\not\subset[-|\log(r)|,|\log(r)|]]<\exp(-c|\log^{3}(r)|)$.
Proof.
This follows immediately from the fact that $H_{1}-H_{1}(0)$ is a Brownian motion with drift $1$ (Theorem 2.9), and Corollary 3.6. We omit the details. ∎
Now take any interval $I$ of length $r^{2/3}$, $I\subset[-|\log(r)|,|\log(r)|]$. Our proof will consist of two steps:
  1. 1.
    Bounding the probability that $A_{1}\cap D_{1,2}\cap I\neq\emptyset$;
  2. 2.
    Bounding the probability that $A_{2}\cap I\neq\emptyset$, conditional on the previous event.
We will show that (for some $\delta>0$) these two probabilities are $<r^{4/3-\delta}$ and $<r^{3\delta}$, respectively.
Let $\mathcal{B}_{1},\mathcal{B}_{2}$ be independent two-sided Brownian motions, with slopes $a_{1}$ and $a_{2}$ respectively. By Theorem 2.9, we can couple $(H_{1},H_{2})$ with $(\mathcal{B}_{1},\mathcal{B}_{2})$ so that
$\mathcal{B}_{1}=H_{1}-H_{1}(0),\quad\tilde{\mathcal{B}}_{2}-\tilde{\mathcal{B}}_{2}(0)=H_{2}-H_{2}(0),$
where
~$\tilde{\mathcal{B}}_{2}(x)=\mathcal{B}_{1}(x)+\sup_{z\leq x}\mathcal{B}_{2}(z)-\mathcal{B}_{1}(z).$
We also write $\mathcal{M}_{1}(x)=\mathcal{M}(x,t;y_{1},1)$ and $\mathcal{M}_{2}(x)=\mathcal{M}(x,t;y_{2},1)$. By 3.1 and the symmetries of $\mathcal{L}$, both of the processes $x\mapsto(1-t)^{-1/3}\mathcal{M}_{i}((1-t)^{2/3}(x-y_{i})),$ $i=1,2$ are parabolic Airy2 processes, so we can apply the comparison result Lemma 2.8 to $\mathcal{M}_{1},\mathcal{M}_{2}$.
We next bound the first probability.
Lemma 4.2.
We have $\mathbb{P}[A_{1}\cap D_{1,2}\cap I\neq\emptyset]<C\log^{44}(r)r^{4/3}$.
For this proof and throughout the remainder of the section, we define $I_{*}=[z_{-},z_{+}]$ to be the (closed) $\log^{10}(r)r^{2/3}$ neighbourhood of $I$.
Proof of Lemma 4.2.
Let $\mathcal{E}_{1}:=\{A_{1}\cap I\neq\emptyset\}$. This is the event where
$\max_{x\in I_{*}}\mathcal{B}_{1}(x)+\mathcal{M}_{1}(x)>\max_{x\in\mathbb{R}}\mathcal{B}_{1}(x)+\mathcal{M}_{1}(x)-\log^{10}(r)r^{1/3}.$
Similarly, $\mathcal{E}_{2}:=\{D_{1,2}\cap I\neq\emptyset\}$ is the event where
$\max_{x\in I_{*}}\mathcal{B}_{2}(x)-\mathcal{B}_{1}(x)>\sup_{x<z_{-}}\mathcal{B}_{2}(x)-\mathcal{B}_{1}(x).$
We define $\mathcal{E}_{1}^{\prime}$:
$\mathcal{B}_{1}(z_{+})+\mathcal{M}_{1}(z_{+})>\max_{z_{+}-1\leq x\leq z_{+}+1}\mathcal{B}_{1}(x)+\mathcal{M}_{1}(x)-2\log^{10}(r)r^{1/3},$
and $\mathcal{E}_{2}^{\prime}$:
$\mathcal{B}_{2}(z_{+})-\mathcal{B}_{1}(z_{+})>\max_{z_{+}-1\leq x\leq z_{+}}\mathcal{B}_{2}(x)-\mathcal{B}_{1}(x)-\log^{10}(r)r^{1/3}.$
Then $\mathcal{E}_{1}\setminus\mathcal{E}_{1}^{\prime}$ implies that
$\max_{x\in I_{*}}\mathcal{B}_{1}(x)+\mathcal{M}_{1}(x)>\mathcal{B}_{1}(z_{+})+\mathcal{M}_{1}(z_{+})+\log^{10}(r)r^{1/3},$
and $\mathcal{E}_{2}\setminus\mathcal{E}_{2}^{\prime}$ implies that
$\max_{x\in I_{*}}\mathcal{B}_{2}(x)-\mathcal{B}_{1}(x)>\mathcal{B}_{2}(z_{+})-\mathcal{B}_{1}(z_{+})+\log^{10}(r)r^{1/3}.$
Then $\mathbb{P}[\mathcal{E}_{1}\setminus\mathcal{E}_{1}^{\prime}],\mathbb{P}[\mathcal{E}_{2}\setminus\mathcal{E}_{2}^{\prime}]<C\exp(-c\log^{5}(r))$, using Lemma 2.8, the stationarity of $x\mapsto\mathcal{A}_{2}(x)+x^{2}$, standard Brownian estimates, and the fact that $I\subset[-|\log(r)|,|\log(r)|]$.
On the other hand, again using Lemma 2.8 and the stationarity of $x\mapsto\mathcal{A}_{2}(x)+x^{2}$, we can deduce that $\mathbb{P}[\mathcal{E}_{1}^{\prime}\cap\mathcal{E}_{2}^{\prime}]<Cr^{4/3}\log^{44}(r)$, by replacing $\mathcal{M}_{1}-\mathcal{M}_{1}(z_{+})$ on $[z_{+}-1,z_{+}+1]$ with a Brownian motion $\mathcal{B}_{3}$ with $\mathcal{B}_{3}(z_{+})=0$ with drift in $[-C|\log(r)|,C|\log(r)|]$, and using Lemma 4.3 below (with $X=\mathcal{B}_{3}(z_{+}-\cdot),Y=\mathcal{B}_{1}(z_{+}-\cdot)-\mathcal{B}_{1}(z_{+}),Z=\mathcal{B}_{2}(z_{+}-\cdot)-\mathcal{B}_{1}(z_{+}),\epsilon=2\log^{10}(r)r^{1/3}$ and $D=C|\log(r)|$).
By putting the above estimates together we get that $\mathbb{P}[\mathcal{E}_{1}\cap\mathcal{E}_{2}]<Cr^{4/3}\log^{44}(r)+C\exp(-c\log^{5}(r))$ and the conclusion follows. ∎
Lemma 4.3.
Take any $D>1$ and $10<\epsilon<$ 1. Let $X,Y,Z$ be three independent Brownian motions, each with drift in $[-D,D]$. Then we have
$\mathbb{P}\left[X(x)+Y(x)<\epsilon,\forall x\in[-1,1];\;Z(x)-Y(x)<\epsilon,\forall x\in[0,1]\right]<C\epsilon^{4}D^{4}.$
Proof.
It is straightforward to calculate that
$\mathbb{P}\left[X(x)+Y(x)<\epsilon,\forall x\in[-1,0]\right]<C\epsilon D.$
It now remains to show that
$\mathbb{P}\left[X(x)+Y(x),Z(x)-Y(x)<\epsilon,\forall x\in[0,1]\right]\leq C\epsilon^{3}D^{3}.$
We can rewrite this as
$\mathbb{P}\left[Z(x)<Y(x)+\epsilon<-X(x)+2\epsilon,\forall x\in[0,1]\right].$
Conditional on $X(1),Y(1),Z(1)$, the processes $Z(x)$, $Y(x)+\epsilon$, $2\epsilon-X(x)$ on $[0,1]$ are Brownian bridges, started at $0,\epsilon,2\epsilon$ respectively. Using the Karlin-McGregor formula, the probability that these processes stay ordered is an explicit Vandermonde determinant, which can be bounded by
$C\epsilon^{3}((Y(1)-Z(1)+\epsilon)\vee 0)((\epsilon-X(1)-Y(1))\vee 0)((2\epsilon-X(1)-Z(1))\vee 0).$
See e.g., [Ham22, Proposition 3.1] for details regarding this standard computation. Taking an expectation over $X(1),Y(1),Z(1)$ gives $C\epsilon^{3}D^{3}$, and the conclusion follows. ∎
We then bound the second probability.
Lemma 4.4.
There exists $\delta>0$ such that the following holds. Whenever $\mathbb{P}[A_{1}\cap D_{1,2}\cap I\neq\emptyset]>r^{10}$, we have $\mathbb{P}[A_{2}\cap I\neq\emptyset\mid A_{1}\cap D_{1,2}\cap I\neq\emptyset]<r^{\delta}$.
Proof.
The event $A_{2}\cap I\neq\emptyset$ is equivalent to
$\max_{x\in I_{*}}\tilde{\mathcal{B}}_{2}(x)+\mathcal{M}_{2}(x)>\max_{x\in\mathbb{R}}\tilde{\mathcal{B}}_{2}(x)+\mathcal{M}_{2}(x)-\log^{10}(r)r^{1/3}.$
This implies the following event, denoted by $\mathcal{E}_{3}$:
$\max_{x\in I_{*}}\tilde{\mathcal{B}}_{2}(x)+\mathcal{M}_{2}(x)>\max_{x\in[z_{+},z_{+}+1]}\mathcal{B}_{2}(x)+\mathcal{M}_{2}(x)-\log^{10}(r)r^{1/3},$
since ~$\tilde{\mathcal{B}}_{2}\geq\mathcal{B}_{2}$. Now, define the event $\mathcal{E}_{3}^{\prime}$ by
$\mathcal{B}_{2}(z_{+})+\mathcal{M}_{2}(z_{+})>\max_{x\in[z_{+},z_{+}+1]}\mathcal{B}_{2}(x)+\mathcal{M}_{2}(x)-2\log^{10}(r)r^{1/3}.$
Recalling the event $\mathcal{E}_{2}$ from the proof of Lemma 4.2, we see that on this event there is some point $z_{*}\in I_{*}$ with $\mathcal{B}_{2}(z_{*})=\tilde{\mathcal{B}}_{2}(z_{*})$. Therefore $\mathcal{E}_{3}\cap\mathcal{E}_{2}\setminus\mathcal{E}_{3}^{\prime}$ implies that
$\mathcal{B}_{2}(z_{*})-\mathcal{B}_{2}(z_{+})+\max_{x\in I_{*}}[\tilde{\mathcal{B}}_{2}(x)-\tilde{\mathcal{B}}_{2}(z_{*})+\mathcal{M}_{2}(x)-\mathcal{M}_{2}(z_{+})]>\log^{10}(r)r^{1/3}.$
Therefore $\mathbb{P}[\mathcal{E}_{3}\cap\mathcal{E}_{2}\setminus\mathcal{E}_{3}^{\prime}]<\exp(-c|\log^{5}(r)|)$ (using Lemma 2.8, the stationarity of Airy2 plus a parabola, standard Brownian motion estimates and the fact that $I\subset[-|\log(r)|,|\log(r)|]$). Now, additionally recall the event $\mathcal{E}_{1}$ from the proof of Lemma 4.2 chosen so that $\mathcal{E}_{1}\cap\mathcal{E}_{2}=\{A_{1}\cap D_{1,2}\cap I\neq\emptyset\}$. We then need to bound $\mathbb{P}[\mathcal{E}_{3}^{\prime}\mid\mathcal{E}_{1}\cap\mathcal{E}_{2}]$.
Observe that $\mathcal{B}_{2}(x)-\mathcal{B}_{2}(z_{+}),x\geq z_{+}$ is independent of $\mathcal{E}_{1}\cap\mathcal{E}_{2}$, and conditionally independent of $\mathcal{M}_{2}$ on $\mathcal{E}_{1}\cap\mathcal{E}_{2}$. Therefore letting $\mathcal{M}_{2}^{\prime}$ be $\mathcal{M}_{2}$ conditional on $\mathcal{E}_{1}\cap\mathcal{E}_{2}$, we have to estimate the probability of
$\mathbb{P}\left[\mathcal{B}_{2}(x)-\mathcal{B}_{2}(z_{+})<-\mathcal{M}_{2}^{\prime}(x)+\mathcal{M}_{2}^{\prime}(z_{+})+2\log^{10}(r)r^{1/3},\;\forall x\in[z_{+},z_{+}+1]\right].$
We again use Lemma 2.8 and the stationarity of Airy2 plus a parabola, to replace $\mathcal{M}_{2}^{\prime}$ with a Brownian motion with drift in $[-C|\log(r)|,C|\log(r)|]$, conditioned on an event of probability at least $cr^{10}$. Then by Lemma 4.5 below (with $D=C|\log(r)|,Y=\mathcal{M}_{2}^{\prime}(\cdot+z_{+})-\mathcal{M}_{2}^{\prime}(z_{+}),X=\mathcal{B}_{2}(\cdot+z_{+})-\mathcal{B}_{2}(z_{+})$, and $\epsilon=2\log^{10}(r)r^{1/3}$) the above probability is at most $Cr^{\delta}$ for some absolute constant $\delta>0$. Putting everything together gives that
$\mathbb{P}[A_{2}\cap I\neq\emptyset\mid A_{1}\cap D_{1,2}\cap I\neq\emptyset]$ $\leq\mathbb{P}[\mathcal{E}_{3}\mid\mathcal{E}_{1}\cap\mathcal{E}_{2}]$
$\leq\mathbb{P}[\mathcal{E}_{3}^{\prime}\mid\mathcal{E}_{1}\cap\mathcal{E}_{2}]+\mathbb{P}[\mathcal{E}_{3}\cap\mathcal{E}_{2}\setminus\mathcal{E}_{3}^{\prime}]\mathbb{P}[\mathcal{E}_{2}\cap\mathcal{E}_{1}]^{-1}$
$\leq Cr^{\delta}+r^{-10}\exp(-c|\log^{5}(r)|),$
and the conclusion follows. ∎
Lemma 4.5.
There exists $\delta>0$ such that the following holds. Take any $D>1$ and $10<\epsilon<$ 1, with $D<\epsilon^{-1/4}$. Let $X$ and $Y$ be two independent Brownian motions, each with drift in $[-D,D]$. Let $\mathcal{E}$ be an event measurable with respect to $Y$, such that $\mathbb{P}[\mathcal{E}]>\epsilon^{100}$. Then
$\mathbb{P}\left[X(x)<Y(x)+\epsilon,\;\forall x\in[0,1]\mid\mathcal{E}\right]<\epsilon^{\delta}.$
Proof.
Let $H$ be a constant whose value is to be determined.
Denote $m=\lfloor|\log_{2}(\epsilon)|\rfloor$. For each $i\in\mathbb{N}$, we let $\mathcal{E}_{i}$ be the event that $Y(2^{-i})\geq H2^{-i/2}$, and $\mathcal{E}_{*}$ be the event where $\mathcal{E}_{i}$ happens for at most half of all $i\in\llbracket 1,2m\rrbracket$. We will show that, by taking $H$ large enough, we have $\mathbb{P}[\mathcal{E}_{*}]>1-\epsilon^{101}$.
We need the following statement, which will be repeatedly used.
Claim: Let $Q\leq R$ be real numbers and let $f,g:[0,1]\to\mathbb{R}\cup\{-\infty\}$ be deterministic functions which are bounded above. Let $W$ be a Brownian motion with drift $R$, conditioned on the event $\{W(x)\geq g(x):x\in[0,1]\}$, and let $Z$ be a Brownian motion with drift $Q$ conditioned on the event $\{Z(x)\geq f(x):x\in[0,1]\}$.
Suppose that for some $s\geq 0$, $f(x)\leq g(x)+s$ for all $x\in[0,1]$ with $f(x)\neq-\infty$. Then $W+s$ stochastically dominates $Z$, in the sense that we can couple the two processes together so that $W(x)+s\geq Z(x)$ for any $x\geq 0$.
Proof: For each $y>1$, we let $W^{(y)}$ (resp. $Z^{(y)}$) be the Brownian bridge on $[0,y]$, with boundary values $W^{(y)}(0)=0$ and $W^{(y)}(y)=Ry$ (resp. $Z^{(y)}(0)=0$ and $Z^{(y)}(y)=Qy$), conditional on $W^{(y)}\geq g$ on $[0,1]$ (resp. $Z^{(y)}\geq f$ on $[0,1]$). The one line case of [CH14, Lemmas 2.6,2.7] shows that we can couple $W^{(y)}$ and $Z^{(y)}$ so that almost surely $W^{(y)}+s\geq Z^{(y)}$ on $[0,y]$. As $y\to\infty$, we have $W^{(y)}\to W$ and $Z^{(y)}\to Z$ weakly in the uniform topology in any compact set. Thus the claim follows. $\blacksquare$
Let $\Phi$ be the set of all length $m$ integer sequences $1\leq i_{1}<\cdots<i_{m}\leq 2m$. For any sequence $(i_{1},\dots,i_{m})$ in $\Phi$, we consider the probability of $\mathcal{E}_{i_{1}}\cap\cdots\cap\mathcal{E}_{i_{m}}$. For each $i\in\llbracket 1,2m\rrbracket$, let $W_{i}$ be a Brownian motion with drift $2^{-(i+1)/2}D$, conditional on $W_{i}(x)\geq x-1$ for any $x\in[0,1]$. By the claim above, for each $j\in\llbracket 1,m\rrbracket$, the process $W_{i_{j}}|_{[0,2]}+H$ stochastically dominates the process $x\mapsto 2^{(i_{j}+1)/2}Y(2^{-i_{j}-1}x)$ conditional on $\bigcap_{k\in\llbracket j+1,m\rrbracket}\mathcal{E}_{i_{k}}$. Therefore
$\mathbb{P}\left[\mathcal{E}_{i_{j}}\bigg{\mid}\bigcap_{k\in\llbracket j+1,m\rrbracket}\mathcal{E}_{i_{k}}\right]<\mathbb{P}[W_{i_{j}}(2)\geq(2^{1/2}-1)H].$
Then when $H>2$, $\mathbb{P}\left[\bigcap_{j\in\llbracket 1,m\rrbracket}\mathcal{E}_{i_{j}}\right]\leq\prod_{j=1}^{m}\mathbb{P}[W_{i_{j}}(2)\geq(2^{1/2}-1)H]$. By comparing $W_{i}$ and $W_{i^{\prime}}$ for $i<i^{\prime}$, using the claim again, we get that $W_{i}$ dominates $W_{i^{\prime}}$. Therefore the above probability is further bounded by $\mathbb{P}[W_{\lfloor m/2\rfloor}(2)\geq(2^{1/2}-1)H]^{m/2}$.
Now, the slope of $W_{\lfloor m/2\rfloor}$ (i.e., $2^{-(\lfloor m/2\rfloor+1)/2}D$) is at most of constant order by the bound $D<\epsilon^{-1/4}$. Therefore as $H\to\infty$, $\mathbb{P}[W_{\lfloor m/2\rfloor}(2)\geq(2^{1/2}-1)H]\to 0$, uniformly in $m$ and $D$.
Then by taking a union bound over all sequences in $\Phi$, we have
$\mathbb{P}[\mathcal{E}_{*}]>1-2^{2m}\mathbb{P}[W_{\lfloor m/2\rfloor}(2)\geq(2^{1/2}-1)H]^{m/2}.$
By taking $H$ large (independent of $m$ and $D$), we have that $\mathbb{P}[\mathcal{E}_{*}]>1-\epsilon^{101}$. Then since $\mathbb{P}[\mathcal{E}]>\epsilon^{100}$, we get $\mathbb{P}[\mathcal{E}_{*}\mid\mathcal{E}]>\frac{\mathbb{P}[\mathcal{E}]-\epsilon^{101}}{\mathbb{P}[\mathcal{E}]}>1-\epsilon$.
We next bound the probability of $X<Y+\epsilon$ in $[0,1]$, conditional on $\mathcal{E}_{*}\cap\mathcal{E}$. For each $i\in\llbracket 1,2m\rrbracket$, we let $\mathcal{E}_{i}^{\prime}$ be the event where $X(2^{-i})\leq H2^{1-i/2}$. We then have
$\mathbb{P}\left[X(x)<Y(x)+\epsilon,\;\forall x\in[0,1]\mid\mathcal{E}_{*}\cap\mathcal{E}\right]\leq\max_{(i_{1},\cdots,i_{m})\in\Phi}\mathbb{P}\left[\bigcap_{j=1}^{m}\mathcal{E}_{i_{j}}^{\prime}\right].$
For each $i\in\llbracket 1,2m\rrbracket$, we let $W^{\prime}_{i}$ be a Brownian motion with drift $-2^{-(i+1)/2}D$, conditional on $W^{\prime}_{i}\leq 0$ in $[0,1]$. Fix $(i_{1},\cdots,i_{m})\in\Phi$. By the claim above, for each $j\in\llbracket 1,m\rrbracket$, $W^{\prime}_{i_{j}}$ is stochastically dominated by the process $x\mapsto 2^{(i_{j}+1)/2}X(2^{-i_{j}-1}x)$ conditional on $\bigcap_{k\in\llbracket j+1,m\rrbracket}\mathcal{E}_{i_{k}}^{\prime}$. Therefore
$\mathbb{P}\left[\mathcal{E}_{i_{j}}^{\prime}\bigg{\mid}\bigcap_{k\in\llbracket j+1,m\rrbracket}\mathcal{E}_{i_{k}}^{\prime}\right]<\mathbb{P}[W_{i_{j}}^{\prime}(2)\leq 2^{3/2}H].$
Therefore
$\mathbb{P}\left[\bigcap_{j=1}^{m}\mathcal{E}_{i_{j}}^{\prime}\right]\leq\prod_{j=1}^{m}\mathbb{P}[W_{i_{j}}^{\prime}(2)\leq 2^{3/2}H].$
As above, by comparing $W_{i}^{\prime}$ and $W_{i^{\prime}}^{\prime}$ for $i<i^{\prime}$ and using the claim, this probability is further bounded by $\mathbb{P}[W_{\lfloor m/2\rfloor}^{\prime}(2)\leq 2^{3/2}H]^{m/2}$. As the slope of $W_{\lfloor m/2\rfloor}^{\prime}$ is at most of constant order, one can bound $\mathbb{P}[W_{\lfloor m/2\rfloor}^{\prime}(2)\leq 2^{3/2}H]$ away from $1$, independent of $m$ and $D$. Thus we conclude that
$\mathbb{P}\left[X(x)<Y(x)+\epsilon,\;\forall x\in[0,1]\mid\mathcal{E}_{*}\cap\mathcal{E}\right]<\epsilon^{\delta_{0}},$
for some $\delta_{0}>0$ small enough. This, together with the lower bound $\mathbb{P}[\mathcal{E}_{*}\mid\mathcal{E}]>1-\epsilon$, implies the conclusion. ∎
Proof of 3.11.
By Lemma 4.2 and Lemma 4.4, for some $\delta>0$ we have
$\mathbb{P}[A_{1}\cap A_{2}\cap D_{1,2}\cap I\neq\emptyset]<C\log^{44}(r)r^{4/3+\delta}+r^{10},$
for any $I$ of length $r^{2/3}$, $I\subset[-|\log(r)|,|\log(r)|]$. Taking a union over at most $3r^{-2/3}\log(r)$ covering $[-|\log(r)|,|\log(r)|]$ and using Lemma 4.1 yields the result. ∎
5 Variants of the characterization theorem
In this section we prove Theorem 1.3, Corollary 1.5, and a technical variant of Theorem 1.3 that will be more appropriate for studying general AEPs. Throughout this section, we will appeal to the following simple lemma.
Lemma 5.1.
  1. 1.
    If $X,Y$ are $\operatorname{UC}$-valued random variables with $X\stackrel{{\scriptstyle d}}{{=}}Y$ and $X\leq Y$ almost surely, then $X=Y$ almost surely.
  2. 2.
    Let $X_{n}\leq Y_{n}$ be sequences of $\operatorname{UC}$-valued random variables such that $X_{n}\stackrel{{\scriptstyle d}}{{\to}}Y$ and $Y_{n}\stackrel{{\scriptstyle\mathbb{P}}}{{\to}}Y$. Then $X_{n}\stackrel{{\scriptstyle\mathbb{P}}}{{\to}}Y$. The same is true if instead $X_{n}\geq Y_{n}$.
Proof.
Proof of 1. First, any upper semicontinuous function $f\in\operatorname{UC}$ is determined by the countable set of values $f_{q,n}:=\sup_{x\in[q-1/n,q+1/n]}f(q)$ where $q\in\mathbb{Q},n\in\mathbb{N}$. Therefore it suffices to show that $X_{q,n}=Y_{q,n}$ almost surely for all fixed $q,n$. Now, if $X_{q,n}<Y_{q,n}$ with positive probability, then we could find some $p$ in $\mathbb{R}$ such that
$\mathbb{P}(X_{q,n}<p)-\mathbb{P}(Y_{q,n}<p)=\mathbb{P}(X_{q,n}<p\leq Y_{q,n})>0)>0,$
contradicting that $X\stackrel{{\scriptstyle d}}{{=}}Y$. Here the equality in the display above uses the almost sure ordering $X\leq Y$ which implies that $X_{q,n}\leq Y_{q,n}$ almost surely as well.
Proof of 2. First observe that $(X_{n},Y_{n})\stackrel{{\scriptstyle d}}{{\to}}(Y,Y)$. Indeed, since $X_{n}\stackrel{{\scriptstyle d}}{{\to}}Y$ and $Y_{n}\stackrel{{\scriptstyle d}}{{\to}}Y$, the sequence $(X_{n},Y_{n})$ is tight. Let $(X,Y)$ be any subsequential limit of $(X_{n},Y_{n})$ in distribution. Since $X_{n}\leq Y_{n}$ (resp. $X_{n}\geq Y_{n}$) almost surely for all $n$ we have $X\leq Y$ (resp. $X\geq Y$) almost surely. Since $X\stackrel{{\scriptstyle d}}{{=}}Y$, part $1$ implies $X=Y$ almost surely. Next, by the continuous mapping theorem $d_{H,\operatorname{loc}}(X_{n},Y_{n})\stackrel{{\scriptstyle d}}{{\to}}d_{H,\operatorname{loc}}(Y,Y)=0$. Since the limit is constant, this convergence also holds in probability. By the triangle inequality,
$d_{H,\operatorname{loc}}(X_{n},Y)\leq d_{H,\operatorname{loc}}(X_{n},Y_{n})+d_{H,\operatorname{loc}}(Y_{n},Y),$
and so since the right-hand side converges to 00 in probability, so does the left-hand side, yielding the conclusion. ∎
Proof of Theorem 1.3.
We first show that for any $f\in\operatorname{NW}$ and $s\in\mathbb{Q}$,
$(\mathcal{K}_{s,s+t}f:t\in\mathbb{Q}\cap[0,\infty))\stackrel{{\scriptstyle d}}{{=}}(\mathfrak{h}_{t}f:t\in\mathbb{Q}\cap[0,\infty)),$ (30)
where here $\mathfrak{h}_{t}$ is the usual KPZ fixed point. To prove the equality, it is enough to show that if $\mathcal{F}_{t}=\sigma(\mathcal{K}_{s,s+r}f:0\leq r\leq t)$ and $t^{\prime}>t$, then conditional on $\mathcal{F}_{t}$ we have
$\mathcal{K}_{s,s+t^{\prime}}f\stackrel{{\scriptstyle d}}{{=}}\mathfrak{h}_{t^{\prime}-t}(\mathcal{K}_{s,s+t}f),$ (31)
where on the right-hand side of (31), the KPZ fixed point evolution $\mathfrak{h}_{t^{\prime}-t}$ is independent of $\mathcal{K}_{s,s+t}f$. To prove (31), let $g_{n}\in\operatorname{NW}$ be a (random) sequence such that $g_{n}\uparrow\mathcal{K}_{s,s+t}f$ almost surely in $\operatorname{UC}$. The existence of such a sequence $g_{n}$ uses that $\mathcal{K}_{s,s+t}f$ is continuous almost surely. Then by the independence of the increments of $\mathcal{K}$, conditional on $\mathcal{F}_{t}$ we have
$\mathfrak{h}_{t^{\prime}-t}g_{n}\stackrel{{\scriptstyle d}}{{=}}\mathcal{K}_{s+t,s+t^{\prime}}g_{n}$
for all $n$. On the other hand, monotonicity of $\mathcal{K}$ implies that $\mathcal{K}_{s+t,s+t^{\prime}}g_{n}$ is a monotone sequence in $\operatorname{UC}$, and so it converges to a limit $G$. By the previous display and the dominated convergence theorem for the KPZ fixed point (2.4) we have that conditional on $\mathcal{F}_{t}$, $G\stackrel{{\scriptstyle d}}{{=}}\mathfrak{h}_{t^{\prime}-t}(\mathcal{K}_{s,s+t}f)$. Finally, since $g_{n}\leq\mathcal{K}_{s,s+t}f$, another application of monotonicity for $\mathcal{K}$ implies that $\mathcal{K}_{s+t,s+t^{\prime}}g_{n}\leq\mathcal{K}_{s,s+t^{\prime}}f$ and so $G\leq\mathcal{K}_{s,s+t^{\prime}}f$ almost surely. On the other hand, unconditionally we have that $G\stackrel{{\scriptstyle d}}{{=}}\mathcal{K}_{s,s+t^{\prime}}f$. Therefore $G=\mathcal{K}_{s,s+t^{\prime}}f$ almost surely by Lemma 5.1, yielding (31) (and therefore (30)).
Now, for $(x,s;y,t)\in\mathbb{Q}^{4}_{\uparrow}$ define $\mathcal{M}(x,s;y,t)=\mathcal{K}_{s,t}\delta_{x}(y)$. Then $\mathcal{M}$ satisfies conditions $1-3$ of Theorem 1.1. Condition $2$ follows by the independent increment property of $\mathcal{K}$, condition $3$ follows from (30), and for the triangle inequality (condition $1$), observe that for any rationals $q,q^{\prime}\in\mathbb{Q}$ with $q^{\prime}\leq\mathcal{K}_{s,r}\delta_{x}(y)\leq q$ we have that
$\mathcal{M}(x,s;y,r)+\mathcal{M}(y,r;z,t)$ $=\mathcal{K}_{s,r}\delta_{x}(y)+\mathcal{K}_{r,t}\delta_{y}(z)$
$\leq q+\mathcal{K}_{r,t}\delta_{y}(z)=\mathcal{K}_{r,t}(\delta_{y}+q^{\prime})(z)+(q-q^{\prime})$
$\leq\mathcal{K}_{s,t}\delta_{x}+(q-q^{\prime})=\mathcal{M}(x,s;z,t)+(q-q^{\prime}).$
Here in the equality on the second line we have used shift commutativity of $\mathcal{K}$. In the inequality on the third line we have used monotonicity of $\mathcal{K}$ together with the fact that $\delta_{y}+q^{\prime}\leq\mathcal{K}_{s,t}\delta_{x}$ by the choice of $q^{\prime}$. Since $q,q^{\prime}$ were arbitrary, this yields the triangle inequality.
Applying Theorem 1.1, we see that $\mathcal{M}$ has a continuous extension to $\mathbb{R}^{4}_{\uparrow}$ which is a directed landscape. This implies (6) when $f=\delta_{x}$ for some $x\in\mathbb{R}$. For general $f\in\operatorname{NW}$, we can write $f=\max_{x\in F}(\delta_{x}+q_{x})$ for some finite rational set $F$ and rationals $q_{x},x\in F$. Therefore
$\mathcal{K}_{s,t}f\geq\max_{x\in F}\mathcal{K}_{s,t}(\delta_{x}+q_{x})=\max_{x\in F}[q_{x}+\mathcal{K}_{s,t}\delta_{x}]=\max_{x\in\mathbb{R}}f(x)+\mathcal{M}(x,s;\cdot,t).$ (32)
Here the inequality uses monotonicity of $\mathcal{K}$, the first equality uses shift commutativity and the second equality is by definition. Now, the left and right-hand sides above are equal in law since $\mathcal{M}$ is a directed landscape, so by Lemma 5.1 they are equal almost surely. ∎
Proof of Corollary 1.5.
We have that almost surely,
$\mathcal{K}_{s,t}f\geq\sup_{x\in\mathbb{Q}}\mathcal{K}_{s,t}(f|_{\{x\}})=\sup_{x\in\mathbb{Q}}f(x)+\mathcal{L}(x,s;\cdot,t)=\sup_{x\in\mathbb{R}}f(x)+\mathcal{L}(x,s;\cdot,t).$
Here the inequality uses monotonicity of $\mathcal{K}$, the first equality uses Theorem 1.3, and the final equality uses condition (7) and continuity of $\mathcal{L}$. Finally, the two sides above are equal in law and so by Lemma 5.1 they are equal almost surely. ∎
We now state and prove a technical variant of Theorem 1.3. First, let $\mathscr{H}$ be the set of continuous functions $h:\mathbb{R}\to\mathbb{R}$ satisfying
  • There exists a finite set $\{q_{1}<\dots<q_{k}\}\subset\mathbb{Q}$ such that $h$ is linear on each of the pieces $[q_{i},q_{i+1}]$, and $h(q_{i})\in\mathbb{Q}$ for all $i$.
  • $h^{\prime}=0$ off of the compact interval $[q_{1},q_{k}]$.
Note that the set $\mathscr{H}$ is countable.
Theorem 5.2.
Suppose that for every $s\in\mathbb{Q}_{+}$ and every $n\in\mathbb{N}$ we have a random continuous function $B_{s,n}:\mathbb{R}\to\mathbb{R}$ such that the tail $\sigma$-algebra $\mathcal{T}_{s}$ for each of the sequence $\{B_{s,n}\}_{n\in\mathbb{N}}$ is trivial.
Let $\mathscr{H}_{s,n}=\mathscr{H}+B_{s,n}$ and $\mathscr{H}_{s}=\bigcup_{n\in\mathbb{N}}\mathscr{H}_{s,n}$. Suppose that for every $s\leq t\in\mathbb{Q}_{+}$ there is a family of random operators $\mathcal{K}_{s,t}:\mathscr{H}_{s}\to\operatorname{UC},s\leq t\in\mathbb{Q}_{+}$ such that:
  1. 1.
    (KPZ fixed point marginals) For $s\leq t\in\mathbb{Q}_{+}$ and $f\in\mathscr{H}_{s}$, we have the equality in law $(\mathcal{K}_{s,t}f,f)\stackrel{{\scriptstyle d}}{{=}}(\mathfrak{h}_{t}f,f)$.
  2. 2.
    (Conditional independence) Fix $t\in\mathbb{Q}_{+}$, and define the $\sigma$-algebra $\mathcal{F}_{t}=\sigma\{B_{s,n},\mathcal{K}_{s,r}:s\leq r\leq t\in\mathbb{Q}_{+},n\in\mathbb{N}\}$. Then for any $t\leq r\in\mathbb{Q}_{+}$, $\mathcal{K}_{t,r}|_{\mathscr{H}_{t,n}}$ is conditionally independent of $\mathcal{F}_{t}$ given $B_{t,n}$.
  3. 3.
    (Comparability) For all $s\leq r\in\mathbb{Q}_{+}$ and $f\in\mathscr{H}_{s},g\in\mathscr{H}_{r}$, we have that $\sup(g-\mathcal{K}_{s,r}f)<\infty$.
  4. 4.
    (Monotonicity) For any $s\leq r<t\in\mathbb{Q}_{+}$ and $f\in\mathscr{H}_{s},g\in\mathscr{H}_{r}$, if $g\leq\mathcal{K}_{s,r}f$ then $\mathcal{K}_{r,t}g\leq\mathcal{K}_{s,t}f$.
  5. 5.
    (Shift commutativity) For all $s<t\in\mathbb{Q}_{+},f\in\mathscr{H}_{s},c\in\mathbb{Q}$, $\mathcal{K}_{s,t}(f+c)=\mathcal{K}_{s,t}f+c$.
Then there exists a directed landscape $\mathcal{L}$ such that almost surely, for any $s<t\in\mathbb{Q}_{+},f\in\mathscr{H}_{s}$,
$\mathcal{K}_{s,t}f=f\diamond\mathcal{L}_{s,t}.$ (33)
Proof.
For every $f\in\operatorname{NW},s\in\mathbb{Q}_{+}$ and $n\in\mathbb{N}$, define a (random) sequence $f_{m}=f_{m}(s,n),m\in\mathbb{N}$ in $\mathscr{H}$ as follows. Let $\{x_{1},\dots,x_{\ell}\}=\{x\in\mathbb{R}:f(x)>-\infty\}$. Define $f_{m}=\max_{i=1,\dots,\ell}(g_{m,i}+f(x_{i}))$ where $g_{m,i}$ is any function chosen that $g_{m,i}(x_{i})\in(-B_{s,n}(x_{i})+1/(m+1),-B_{s,n}(x_{i})+1/m]$, and
$g_{m,i}^{\prime}(x)=\begin{cases}m,\qquad&x\in(x_{i}-1,x_{i}),\\ -m,\qquad&x\in(x_{i},x_{i}+1),\\ 0,\qquad&x\notin[x_{i}-1,x_{i}+1].\end{cases}$
Then $B_{s,n}+f_{m}(s,n)\downarrow f$ in $\operatorname{UC}_{0}$. By monotonicity of $\mathcal{K}$, for any fixed $t\geq s\in\mathbb{Q}_{+}$ the functions $\mathcal{K}_{s,t}(B_{s,n}+f_{m}(s,n))$ form a monotone sequence almost surely. Let $\tilde{\mathcal{K}}_{s,t}^{n}f$ denote the almost sure limit of this monotone sequence. By 2.4, $\tilde{\mathcal{K}}_{s,t}^{n}f\stackrel{{\scriptstyle d}}{{=}}\mathfrak{h}_{t-s}f$.
We check that $\tilde{\mathcal{K}}_{s,t}^{n}f=\tilde{\mathcal{K}}_{s,t}^{n^{\prime}}f$ almost surely for all $n,n^{\prime}$. Indeed, by Condition $3$ (comparability) and our choice of approximation, there exists some constant $c>0$ such that $B_{s,n}\leq B_{s,n^{\prime}}+c$. Then for any fixed $m^{\prime}\in\mathbb{N}$ there exists $m_{0}(m^{\prime})$ such that for all $m\geq m_{0}(m^{\prime})$ we have that
$B_{s,n}+f_{m}(s,n)\leq B_{s,n^{\prime}}+f_{m^{\prime}}(s,n^{\prime}).$
Using this inequality and monotonicity of $\mathcal{K}$ implies that $\tilde{\mathcal{K}}_{s,t}^{n}f\leq\tilde{\mathcal{K}}_{s,t}^{n^{\prime}}f$. By a symmetric argument, we also have that $\tilde{\mathcal{K}}_{s,t}^{n}f\geq\tilde{\mathcal{K}}_{s,t}^{n^{\prime}}f$. From now on we drop the superscript $n$ and write $\tilde{\mathcal{K}}_{s,t}f=\tilde{\mathcal{K}}_{s,t}^{n}f$. We check that the family $\tilde{\mathcal{K}}_{s,t}f$ satisfies the conditions of Theorem 1.3.
The KPZ fixed point marginal condition was established above, and shift commutativity is immediate since we chose the prelimiting sequences so that $f_{m}+q=(f+q)_{m}$ for $q\in\mathbb{Q}$, and the original operators $\mathcal{K}_{s,t}:\mathscr{H}_{s}\to\operatorname{UC}$ satisfy shift commutativity.
We move to the independent increment property. Fix $t\leq r\in\mathbb{Q}_{+}$. First observe that by condition $2$ above (conditional independence), for any $\sigma$-algebra $\mathcal{A}$ with $\sigma(B_{t,n})\subset\mathcal{A}\subset\mathcal{F}_{t}$, $\mathcal{K}_{t,r}|_{\mathscr{H}_{t,n}}$ and $\mathcal{F}_{t}$ are conditionally independent given $\mathcal{A}$. We will use this with $\mathcal{A}=\mathcal{T}_{t,n}:=\sigma(B_{t,n^{\prime}}:n^{\prime}\geq n)$. Let $\mathcal{G}_{t}=\sigma(\tilde{\mathcal{K}}_{s,s^{\prime}}:s\leq s^{\prime}\leq t\in\mathbb{Q}_{+})$ and let $A\in\mathcal{G}_{t}\subset\mathcal{F}_{t}$. Let $B\in\sigma(\tilde{\mathcal{K}}_{t,r})$ and note that $B\in\sigma(\mathcal{K}_{t,r}|_{\mathscr{H}_{t,n}})$ for any fixed $n$ since $\tilde{\mathcal{K}}_{t,r}f=\tilde{\mathcal{K}}_{t,r}^{n}f$ for $f\in\operatorname{NW}$. Therefore
$\mathbb{P}[A\cap B\mid\mathcal{T}_{t,n}]=\mathbb{P}[A\mid\mathcal{T}_{t,n}]\mathbb{P}[B\mid\mathcal{T}_{t,n}].$ (34)
Taking $n\to\infty$ in (34) gives that $\mathbb{P}[A\cap B\mid\mathcal{T}_{t}]=\mathbb{P}[A\mid\mathcal{T}_{t}]\mathbb{P}[B\mid\mathcal{T}_{t}]$ where $\mathcal{T}_{t}=\bigcap_{n\in\mathbb{N}}\mathcal{T}_{t,n}$ is the tail $\sigma$-algebra. The existence and evaluation of this limit uses backward martingale convergence. Then, since the tail $\sigma$-algebra $\mathcal{T}_{t}$ is trivial, $\mathbb{P}[A\cap B]=\mathbb{P}[A]\mathbb{P}[B]$, and so $\tilde{\mathcal{K}}_{t,r}$ is independent of $\mathcal{G}_{t}$. This implies the independent increment property.
Finally we prove monotonicity. Fix $s\leq r<t\in\mathbb{Q}_{+}$ and suppose that $g\leq\tilde{\mathcal{K}}_{s,r}f$ for some $g,f\in\operatorname{NW}$. Since we constructed $\tilde{\mathcal{K}}_{s,t}f$ as the decreasing limit of $\mathcal{K}_{s,r}[B_{s,1}+f_{m}(s,1)]$, for any $m\in\mathbb{N}$ we have that $g\leq\mathcal{K}_{s,r}[B_{s,1}+f_{m}(s,1)].$ The construction of $g$ from $B_{r,1}+g_{m}(r,1)$ and the fact that $\sup(B_{r,1}-\mathcal{K}_{s,t}[B_{s,1}+f_{m}(s,1)])<\infty$ for all fixed $m$ (condition $3$ above) implies that for any fixed $m\in\mathbb{N}$ and $q\in\mathbb{Q}_{+}$ we can find some $m_{0}(m)\in\mathbb{N}$ such that for $m^{\prime}\geq m_{0}(m)$ we have
$B_{r,1}+g_{m^{\prime}}(r,1)\leq\mathcal{K}_{s,r}[B_{s,1}+f_{m}(s,1)]+q=\mathcal{K}_{s,r}[B_{s,1}+f_{m}(s,1)+q].$
Therefore using monotonicity of $\mathcal{K}$ we have that
$\mathcal{K}_{r,t}[B_{r,1}+g_{m^{\prime}}(r,1)]\leq\mathcal{K}_{s,t}[B_{s,1}+f_{m}(s,1)+q]=\mathcal{K}_{s,t}[B_{s,1}+f_{m}(s,1)]+q.$
Thus taking $m^{\prime}\to\infty$, then $m\to\infty$, and then $q\to 0$ gives that $\tilde{\mathcal{K}}_{r,t}g\leq\tilde{\mathcal{K}}_{s,t}f$ almost surely, as desired.
Therefore appealing to Theorem 1.3 we have that formula (33) holds for some directed landscape $\mathcal{L}$, with $f\in\operatorname{NW}$ and $\tilde{\mathcal{K}}$ in place of $\mathcal{K}$. To extend this to $\mathcal{K}$ and $f\in\mathscr{H}_{s}$, we can apply Corollary 1.5. The only thing to check is that for $g\in\operatorname{NW}$ and $f\in\mathscr{H}_{s}$ with $g\leq f$ we have that $\tilde{\mathcal{K}}_{s,t}g\leq\mathcal{K}_{s,t}f$. For this, observe that as in the previous paragraph, we can argue that for any fixed rational $q\in\mathbb{Q}_{+}$, for all large enough $m$ we have $B_{s,1}+g_{m}(s,1)\leq f+q$, and so $\tilde{\mathcal{K}}_{s,t}g\leq\mathcal{K}_{s,t}[f+q]=\mathcal{K}_{s,t}f+q$, which gives the result since $q\in\mathbb{Q}_{+}$ is arbitrary. ∎
In the remaining sections, we apply our general framework to prove convergence of specific models to the directed landscape. All convergence results are new, except for convergence of colored TASEP, colored ASEP and the KPZ equation. In those cases, our method offers a shorter route to proving convergence that bypasses a technical analysis of the underlying line ensembles.
6 Web distances
For coalescing random walks, a classical probability model and the dual of the voter model, the trajectories form an infinite tree. In the discrete-time setting with state space $\mathbb{Z}$, this structure gives rise to a $(1+1)$-dimensional random walk web. Under diffusive scaling limits, it converges to the Brownian web, which can be understood as coalescing Brownian motions. The Brownian web was first constructed by Tóth and Werner [TW98], building on the work of Arratia [Arr79]. The convergence of the random walk web to the Brownian web was established in [FINR04], where the term Brownian web was introduced. More background and results about this model can be found in the related survey [SSS17].
In the random walk web and the Brownian web, Vető and Virág [VV23] introduced two models of random directed metrics. Through a correspondence with classical models of last passage percolation, they showed that certain two-parameter marginals of these directed metrics converge to the KPZ fixed point run from the narrow wedge initial condition. The correspondence with classical last passage percolation models does not extend to the full four-parameter field in these models and so convergence to the directed landscape was not clear. They left this as a conjecture in their paper. We can resolve this conjecture with an application of Theorem 1.1.
We next define these two models precisely. Let $\mathbb{Z}^{2}_{e}=\{(i,n)\in\mathbb{Z}^{2}:i+n\text{ is even}\}$ be the even lattice, and from every point $v=(i,n)$ assign two outgoing directed edges to $(i-1,n+1)$ and $(i+1,n+1)$. Let $\{\zeta(v),v\in\mathbb{Z}^{2}_{e}\}$ be a field of i.i.d. Rademacher random variables. The random walk web $Y$ is a family of coalescing random walks $Y_{(i,n)}:\llbracket n,\infty\rrbracket\to\mathbb{Z}$ started at each point $(i,n)\in\mathbb{Z}^{2}_{e}$, given by setting $Y_{(i,n)}(n)=i$ and
$Y_{(i,n)}(j+1)=Y_{(i,n)}(j)+\zeta(Y_{(i,n)}(j),j),\qquad\text{ for }j\in\llbracket n,\infty\rrbracket.$
We can now define the random walk web distance $D^{\operatorname{RW}}$.
Definition 6.1.
For any $(i,n;j,m)\in(\mathbb{Z}^{2}_{e})^{2}$ let $D^{\operatorname{RW}}(i,n;j,m)$ be the smallest non-negative integer $k$ such that there are points $(i_{0},n_{0})=(i,n),\dots,(i_{k},n_{k}),(i_{k+1},n_{k+1})=(j,m)\in\mathbb{Z}^{2}_{e}$ with the following property. There are random walk paths in $Y$ from $(i_{0},n_{0})$ to $(i_{1},n_{1})$ and from $(i_{j}-\zeta(i_{j},n_{j}),n_{j}+1)$ to $(i_{j+1},n_{j+1})$ for all $1\leq j\leq k$. We set $D^{\operatorname{RW}}(i,n;j,m)=\infty$ if no such $k$ exists.
In the scaling limit where the system of random walks $D^{\operatorname{RW}}$ converges to a system of coalescing Brownian motions (the Brownian web), the random walk web distance converges to a limit known as the Brownian web distance $D^{\operatorname{Br}}$ which has a description similar to Definition 6.1. To define this distance, we first need a precise definition of the Brownian web.
We can give an explicit recursive construction of the rational Brownian web $(B_{q})_{q\in\mathbb{Q}^{2}}$:
  1. 1.
    Let $q_{i}=(x_{i},s_{i}),i\in\mathbb{N}$ be any enumeration of $\mathbb{Q}^{2}$, and sample independent standard Brownian paths $W_{q_{i}},i\in\mathbb{N}$, where $W_{q_{i}}:[s_{i},\infty)\to\mathbb{R}$ has initial condition $W_{q_{i}}(s_{i})=x_{i}$. Set $B_{q_{1}}=W_{q_{1}}.$
  2. 2.
    For all $i\geq 2$, given $B_{q_{1}},\dots,B_{q_{i-1}}$, let $T$ be the first time in $[s_{i},\infty)$ when $W_{q_{i}}(t)=B_{q_{j}}(t)$ for some $j<i$, and let $J$ be the minimal index such that $W_{q_{i}}(T)=B_{q_{J}}(T)$. Then set $B_{q_{i}}(t)=W_{q_{i}}(t)$ for $t\leq T$ and $B_{q_{i}}(t)=B_{q_{J}}(t)$ for $t\geq T$.
Intuitively, this construction gives a countable collection of Brownian motions that run independently until they hit, after which point they run together. The law of this construction is independent of the choice of the enumeration of $\mathbb{Q}$. To construct the full Brownian web, we take the closure of the set of paths $(B_{q})_{q\in\mathbb{Q}}$ is an appropriate topology.
Following [FINR04], we put a topology on the space $\Pi$ of all continuous functions $f:[t,\infty)\to\mathbb{R}$, where $t\in\mathbb{R}$. First, for such a function $f$, let $\hat{f}:\mathbb{R}\to\mathbb{R}$ be the function equal to $f$ on $[t,\infty)$ and equal to $f(t)$ on $(-\infty,t)$. We say that a sequence of continuous functions $f_{n}:[t_{n},\infty)\to\mathbb{R}$ converges to a limit $f:[t,\infty)\to\mathbb{R}$ if $t_{n}\to t$ and $\hat{f}_{n}\to\hat{f}$ uniformly on compact sets. This defines a Polish space. The Brownian web is the closure $\mathcal{B}$ of the rational Brownian web $\{B_{q}:q\in\mathbb{Q}^{2}\}$ in this space. Given the Brownian web, we can now define the Brownian web distance.
Definition 6.2.
For $(x,s;y,t)\in\mathbb{R}^{4}$, define the Brownian web distance $D^{\operatorname{Br}}(x,s;y,t)$ to be the infimum over all non-negative integers $k$ for which there exist points
$(x_{0},t_{0})=(x,s),(x_{1},t_{1}),\dots,(x_{k},t_{k}),(x_{k+1},t_{k+1})=(y,t)\in\mathbb{R}^{2}$
such that $s<t_{1}<\cdots<t_{k}<t$, and there is a continuous path $\pi:[s,t]\to\mathbb{R}$ such that $\pi(t_{i})=x_{i}$ for each $i\in\llbracket 0,k+1,\rrbracket$, and for each $i\in\llbracket 0,k\rrbracket$ there exists $B_{i}\in\mathcal{B}$ with $\pi|_{[t_{i},t_{i+1}]}=B_{i}|_{[t_{i},t_{i+1}]}$. The infimum is equal to +$\infty$+ ∞ if there is no such $k\in\mathbb{N}$.
Note that in Definition 6.2, the points $(t_{i},x_{i}),i=1,\dots,k$ are special: these are points that both lie on the interior of some Brownian path $B\in\mathcal{B}$, and at the beginning of a different Brownian path $B^{\prime}\in\mathcal{B}$ which is not contained in $B$. Such points are not visible if we restrict our attention to the rational Brownian web, since when restricted to the rationals, there is a unique Brownian path emanating from each point.
Neither of the distances $D^{\operatorname{RW}}$ or $D^{\operatorname{Br}}$ are true metrics, because they are not symmetric or finite everywhere, but both are directed metrics on $\mathbb{Z}^{2}_{e}$ and $\mathbb{R}^{2}$, respectively.
One of the main results of [VV23] is a two-parameter convergence result relating $D^{\operatorname{RW}},D^{\operatorname{Br}}$ to the directed landscape. To set up this theorem, for $\eta\in(0,1)$ define positive constants
$a_{\eta}=\frac{(1-\eta^{2})^{1/6}}{(\eta/2)^{2/3}},\quad b_{\eta}=\frac{1-\sqrt{1-\eta^{2}}}{2},\quad c_{\eta}=\frac{\eta^{1/3}(1-\eta^{2})^{1/6}}{2^{1/3}},\quad d_{\eta}=2^{2/3}\eta^{1/3}(1-\eta^{2})^{2/3},$
and for $n\in\mathbb{N}$ define rescalings
$\mathcal{M}^{\eta}_{n}(x,s;y,t)$ $=-a_{\eta}n^{-1/3}\left(D^{\operatorname{RW}}(\eta nt+d_{\eta}n^{2/3}y,-nt;\eta ns+d_{\eta}n^{2/3}x,-ns)-b_{\eta}n(t-s)-c_{\eta}n^{2/3}(y-x)\right),$
$\mathcal{K}_{n}(x,s;y,t)$ $=-n^{1/3}\Big{(}D^{\operatorname{Br}}(2yn^{2/3}+2tn,-tn;2xn^{2/3}+2sn,-sn)-n(t-s)-2n^{2/3}(y-x)\Big{)}.$
Here in the rescaling of $D^{\operatorname{RW}}$ we actually need to take integer parts in the argument in order to put the points on the lattice $\mathbb{Z}^{2}_{e}$; throughout this section we ignore this minor point, for the convenience of notations. Note also that while we consider rescalings of $D^{\operatorname{RW}}$ in all directions $\eta$, we only consider the rescaling of $D^{\operatorname{Br}}$ in one direction; other directions are equivalent by Brownian scaling. Finally, it is worth pointing out the argument exchange on $(x,s)$ and $(y,t)$ when we move between $D^{\operatorname{Br}},D^{\operatorname{RW}}$ and the limiting versions. We do this so that the main theorems of [VV23] translate to KPZ fixed point marginals moving forwards in time, rather than backwards. The directed landscape is time-reversible, so we would ultimately end up with the same convergence theorem without this step.
At Lebesgue almost every $u=(x,s)\in\mathbb{R}^{2}$, $\mathcal{K}_{n}(u;\cdot)\equiv-\infty$ almost surely [VV23, Proposition 3.3], and a similar phenomenon happens in the limit of $\mathcal{M}^{\eta}_{n}$. Therefore in order to see the directed landscape in the limit, we need to modify the definition slightly. Define $f_{n},g_{n}^{\eta}:\mathbb{R}\to\mathbb{R}\cup\{-\infty\}$ by letting $f_{n}(\eta)=g_{n}^{\eta}(x)=-\infty$ for $x>0$ and setting $f_{n}(x)=2n^{1/3}x$ and $g_{n}^{\eta}(x)=c_{\eta}a_{\eta}n^{1/3}x$ for $x\leq 0$. Then set
$\begin{split}\widetilde{\mathcal{K}}_{n}(x,s;y,t)=\max_{z\in\mathbb{R}}f_{n}(z-x)+\mathcal{K}_{n}(z,s;y,t),\\ \widetilde{\mathcal{M}}_{n}^{\eta}(x,s;y,t)=\max_{z\in\mathbb{R}}g_{n}^{\eta}(z-x)+\mathcal{M}_{n}^{\eta}(x,s;y,t).\end{split}$ (35)
In other words, if we view $\mathcal{K}_{n},\mathcal{M}_{n}^{\eta}$ as describing the driving noise for a growth process, $\widetilde{\mathcal{K}}_{n},\widetilde{\mathcal{M}}_{n}^{\eta}$ represents that growth process started at soft narrow wedges, either $f_{n}$ or $g_{n}^{\eta}$.
Theorem 6.3 (Theorem 1.6 and Theorem 1.9, [VV23]).
As $n\to\infty$,
$\widetilde{\mathcal{K}}_{n}(0,0;y,t)\stackrel{{\scriptstyle d}}{{\to}}\mathcal{L}(0,0;y,t)$
with respect to the topology of uniform convergence on compact subsets of $(y,t)\in\mathbb{R}\times(0,\infty)$. Similarly, for any fixed $\eta\in(0,1)$, as $n\to\infty$, for any finite set $F\subset(0,\infty)$,
~$\widetilde{\mathcal{M}}_{n}^{\eta}(0,0;y,t)\stackrel{{\scriptstyle d}}{{\to}}\mathcal{L}(0,0;y,t)$
with respect to the topology of uniform convergence on compact subsets of $(y,t)\in\mathbb{R}\times F$.
Remark 6.4.
The notations we use here are slight different from those in [VV23], but they are easily checked to be equivalent. Moreover, Theorem 1.9 in [VV23] (which concerns ~$\widetilde{\mathcal{M}}_{n}^{\eta}$) is stated only for $F=\{1\}$. However, the proof goes through for any finite set $F$ because of the coupling with the Seppäläinen-Johannson model/reflected simple random walks proven in that paper (see Proposition 3.7 therein). Vető and Virág chose to state their theorem only when $F=\{1\}$ only because they did not have access to the upgraded uniform convergence theorem (see the discussion after Theorem 1.9 in that paper).
Given the two-parameter convergence in Theorem 6.3, one may expect that $\widetilde{\mathcal{K}}_{n},\widetilde{\mathcal{M}}_{n}^{\eta}$ converge to $\mathcal{L}$ as a four-parameter fields. This is Conjecture 1.8 in [VV23] for $\widetilde{\mathcal{K}}_{n}$. Even the convergence $\widetilde{\mathcal{K}}_{n}(\cdot,-1;0,0)\stackrel{{\scriptstyle d}}{{\to}}\mathcal{L}(\cdot,-1;0,0)$ is far from clear, because there is no natural symmetry in $x,y$ in the definition of $D^{\operatorname{Br}}(x,s;y,t)$. This was also conjectured in [VV23] (Conjecture 1.7 therein). As an application of our first characterization Theorem 1.1, we can establish full convergence of $\widetilde{\mathcal{K}}_{n},\widetilde{\mathcal{M}}_{n}^{\eta}\stackrel{{\scriptstyle d}}{{\to}}\mathcal{L}$. We restrict our attention to proving convergence of finite-dimensional distributions. We expect no conceptual barrier to extending to uniform in compact convergence (especially for $\widetilde{\mathcal{K}}_{n}$), but this would require extra technical work.
Theorem 6.5.
Let (~$(\widetilde{\mathcal{M}}_{n},n\in\mathbb{N})$ equal $(\mathcal{M}_{n}^{\eta},n\in\mathbb{N})$ for some fixed $\eta$, or $(\tilde{\mathcal{K}}_{n},n\in\mathbb{N})$. Then for any finite set $F\subset\mathbb{R}^{4}_{\uparrow}$, as $n\to\infty$, ~$\widetilde{\mathcal{M}}_{n}|_{F}\stackrel{{\scriptstyle d}}{{\to}}\mathcal{L}|_{F}.$
Given our Theorem 1.1, and Theorem 6.3, the only non-trivial task is to check that the triangle inequality passes to the limit. This is not obvious, since while $\mathcal{M}_{n}^{\eta},\mathcal{K}_{n}$ satisfy the triangle inequality, ~$\widetilde{\mathcal{M}}_{n}^{\eta},\widetilde{\mathcal{K}}_{n}$ do not.
Lemma 6.6.
Fix $o=(x,s),p=(y,r),q=(z,t)\in\mathbb{R}^{2}$ with $s<r<t$. Let ~$\widetilde{\mathcal{M}}_{n}$ be as in Theorem 6.5. For any $\delta>0$, we have
$\lim_{n\to\infty}\mathbb{P}\left[\widetilde{\mathcal{M}}_{n}(o,p)+\widetilde{\mathcal{M}}_{n}(p,q)\leq\widetilde{\mathcal{M}}_{n}(o,q)+\delta\right]=1.$ (36)
Proof.
Let $\mathcal{M}_{n}$ equal $\mathcal{K}_{n}$ or $\mathcal{M}_{n}^{\eta}$, depending on whether ~$\widetilde{\mathcal{M}}_{n}$ equals $\widetilde{\mathcal{K}}_{n}$ or ~$\widetilde{\mathcal{M}}_{n}^{\eta}$. Then we have the metric composition law
$\max_{w\in\mathbb{R}}\widetilde{\mathcal{M}}_{n}(o;w,r)+\mathcal{M}_{n}(w,r;q)=\widetilde{\mathcal{M}}_{n}(o;q).$ (37)
Let $g_{n}$ equal $g_{n}^{\eta}$ or $f_{n}$, to be narrow wedge used in the definition of ~$\widetilde{\mathcal{M}}_{n}$. Then if we let $y_{n}$ be the argmax of the map
$u\mapsto g_{n}(u-y)+\mathcal{M}_{n}(u,r;q),$
we have that ~$\widetilde{\mathcal{M}}_{n}(p,q)=g_{n}(y_{n}-y)+\mathcal{M}_{n}(y_{n},r;q)\leq\mathcal{M}_{n}(y_{n},r;q),$ and so by (37) we have that
~$\widetilde{\mathcal{M}}_{n}(o,p)+\widetilde{\mathcal{M}}_{n}(p,q)\leq\widetilde{\mathcal{M}}_{n}(o,p)+\mathcal{M}_{n}(y_{n},r;q)\leq\widetilde{\mathcal{M}}_{n}(o,q)+|\widetilde{\mathcal{M}}_{n}(o;p)-\widetilde{\mathcal{M}}_{n}(o;y_{n},r)|.$
Now, ~$\widetilde{\mathcal{M}}_{n}(o;\cdot,r)$ converges in law to a random continuous function uniformly on compact sets. This uses Theorem 6.3 and translation invariance of ~$\widetilde{\mathcal{M}}_{n}$, which follows from translation invariance of $D^{\operatorname{Br}}$ and $D^{\operatorname{RW}}$. Therefore to prove (36), it suffices to show that $y_{n}\stackrel{{\scriptstyle\mathbb{P}}}{{\to}}y$ as $n\to\infty$.
To see why $y_{n}\stackrel{{\scriptstyle\mathbb{P}}}{{\to}}y$, first note that $y_{n}\leq y$ by the definition of $g_{n}$. Now fix $\delta>0$ and observe that on the event $y_{n}\leq y-\delta$, we have the equality
$\widetilde{M}_{n}(p;q)=\widetilde{M}_{n}(y-\delta,r;q)-g_{n}(-\delta).$
Since both $\widetilde{M}_{n}(p;q),\widetilde{M}_{n}(y-\delta,r;q)$ converge in law to finite random variables as $n\to\infty$ by Theorem 6.3 and translation invariance, and $g_{n}(-\delta)\to-\infty$, this implies that $\mathbb{P}[y_{n}\leq y-\delta]\to 0$ as $n\to\infty$, as desired. ∎
Proof of Theorem 6.5.
Let $Q\subset\mathbb{R}$ be a countable dense set such that $F\subset Q^{4}_{\uparrow}:=Q^{4}\cap\mathbb{R}^{4}_{\uparrow}$. The sequence ~$\widetilde{\mathcal{M}}_{n}|_{Q^{4}_{\uparrow}}$ is tight, since its one-point laws ~$\widetilde{\mathcal{M}}_{n}(u)$ are all tight by Theorem 6.3. Let ~$\widetilde{\mathcal{M}}:Q^{4}_{\uparrow}\to\mathbb{R}$ be any subsequential limit in law of ~$\widetilde{\mathcal{M}}_{n}|_{Q^{4}_{\uparrow}}$. Then $\widetilde{\mathcal{M}}$ has independent increments, since the same was true of ~$\widetilde{\mathcal{M}}_{n}$. It also satisfies the triangle inequality by Lemma 6.6, and the marginals of ~$\widetilde{\mathcal{M}}(o;\cdot)$ match those of the directed landscape for any fixed $o\in Q^{2}$ by Theorem 6.3 and translation invariance of $\mathcal{M}_{n}$. Therefore by Theorem 1.1, ~$\widetilde{\mathcal{M}}=\mathcal{L}|_{Q^{4}_{\uparrow}}$ for some directed landscape $\mathcal{L}$, yielding the result. ∎
7 Polymer models
We next consider two polymer models: the KPZ equation and the O’Connell-Yor polymer. For the KPZ equation, convergence to the directed landscape was shown in [Wu23], and so our result here is simply an alternate proof. For the O’Connell-Yor polymer, the techniques used by [Wu23] to establish convergence of the KPZ equation should also be applicable, but the required technical work has not been carried out.
7.1 The KPZ equation
Introduced by Kardar, Parisi, and Zhang [KPZ86], the KPZ equation is the following stochastic PDE:
$\partial_{t}h=\frac{\lambda}{2}(\partial_{x}h)^{2}+\nu\partial^{2}_{x}h+\xi,$
where $\xi$ is a space-time Gaussian white noise, and $\lambda,\nu$ are parameters. Since the seminal work [BG97], a solution for the KPZ equation in terms of a Cole-Hopf transform has been widely used, and a more precise and rigorous solution theory was developed in [Hai13].
Over the past 15 years, exact formulas and long-time scaling limits have been proven for the KPZ equation, see e.g., [ACQ11, Dot13, SS10, IS11, BCFV15]. More recently, KPZ fixed point convergence was independently established in [QS23] and [Vir20], and directed landscape convergence was shown in [Wu23]. Here we give a short proof of directed landscape convergence, using the results of [QS23, Vir20] as the key inputs. Compared to [Wu23], our proof does not go through the KPZ line ensemble.
We start by stating KPZ fixed point convergence, following [QS23]. For $\delta>0$, consider the following KPZδ equation:
$\partial_{t}h=\frac{1}{4}(\partial_{x}h)^{2}+\frac{\delta}{4}\partial^{2}_{x}h+\delta^{1/2}\xi.$ (38)
We shall consider its Cole-Hopf solution $h=4\delta\log Z$, where $Z$ is the solution to the (rescaled) stochastic heat equation (SHE) with multiplicative noise:
$\partial_{t}Z=\delta\partial^{2}_{x}Z+\frac{1}{4}\delta^{-1/2}\xi Z.$ (39)
The solution for the SHE can be thought of as the partition function for a Brownian polymer weighted by the white noise $\xi$ [AKQ14]. More precisely, given the white noise $\xi$ there is a random field $Z^{\delta}:\mathbb{R}^{4}_{\uparrow}\to\mathbb{R}$ such that $Z^{\delta}(x,s;\cdot,\cdot)$ solves the SHE started from the Dirac $\delta$-measure at location $x$ and time $s$, where formally we have
[…]
Here the expectation is taken over a Brownian motion with $B(x)=s$ and variance $2\delta$. To make sense of the previous display rigorously and connect this with prelimiting models (i.e. the O’Connell-Yor polymer) one can use a chaos expansion. The SHE solution $Z$ for more general measure-valued initial conditions can be built from the field $Z^{\delta}$ using the fact that the SHE is linear in the initial condition. Applying the Cole-Hopf transform then solves the KPZ equation; here Dirac $\delta$-initial conditions for the SHE become narrow wedge initial conditions for the KPZ equation.
Theorem 7.1 ([QS23, Theorem 2.2(3)]).
Consider $f\in\operatorname{UC}$ which is either a sum of finitely many narrow wedge initial conditions, or is the maximum of such a function with a continuous function satisfying that $\sup_{x}f(x)/(1+|x|)<\infty$. Take the Cole-Hopf solution $h$ to (38) with initial condition $h(0,\cdot)=f$. As $\delta\to 0$ we have $h(t,\cdot)+\frac{t}{12}\stackrel{{\scriptstyle d}}{{\to}}\mathfrak{h}_{t}f$, in the uniform-on-compact topology, jointly for finitely many $t>0$. In particular, this holds for $h(t,\cdot)=4\delta\log Z^{\delta}(x,s;\cdot,s+t)$ (the narrow wedge initial condition at location $x$, time $s$).
Using Theorem 1.1 and Theorem 7.1, we can recover the convergence of the KPZ equation to the directed landscape from [Wu23].
Theorem 7.2.
The function $(x,s;y,t)\mapsto\mathcal{H}^{\delta}(x,s;y,t)+\frac{t-s}{12}$ converges in law to $\mathcal{L}$ as $\delta\to 0$, in the sense of finite dimensional distributions.
We remark that this finite dimensional convergence can be upgraded to (1) uniform-on-compact convergence, or (2) the convergence of multiple KPZ equations with initial conditions as in Theorem 7.1 coupled via the same space-time white noise. For these stronger convergence statements, see [Wu23, Theorem 1.6, Corollary 1.9]. To derive these upgrades from Theorem 7.2, it suffices to prove the tightness of $\mathcal{H}^{\delta}$ as a continuous function on $\mathbb{R}^{4}_{\uparrow}$. This follows from modulus of continuity bounds of $\mathcal{H}^{\delta}$ from e.g., [CGH21, DG23], and has been detailed in [Wu23, Proposition 6.2].
Proof of Theorem 7.2.
Fix a finite set $F\subset\mathbb{R}^{4}_{\uparrow}$, and let $Q\subset\mathbb{R}$ be a countable dense set such that $F\subset Q^{4}\cap\mathbb{R}^{4}_{\uparrow}$. Let $\mathcal{M}:Q^{4}\cap\mathbb{R}^{4}_{\uparrow}\to\mathbb{R}$ be a joint subsequential limit in law as $\delta\to 0$ of $(x,s;y,t)\mapsto\mathcal{H}^{\delta}(x,s;y,t)+\frac{t-s}{12},(x,s;y,t)\in Q^{4}\cap\mathbb{R}^{4}_{\uparrow}$.
We will apply Theorem 1.1 to $\mathcal{M}$. We check that the three conditions of that theorem are satisfied. Independence of increments is obvious, and the KPZ fixed point marginal condition follow from Theorem 7.1. We next prove the triangle inequality. Take any $o=(x,s)$, $p=(y,r)$, $q=(z,t)\in Q^{2}$ with $s<r<t$. The
$\exp\left((4\delta)^{-1}\mathcal{H}^{\delta}(o;q)\right)$ $=\int\exp\left((4\delta)^{-1}(\mathcal{H}^{\delta}(o;w,r)+\mathcal{H}^{\delta}(w,r;q))\right)\mathrm{d}w$
$\geq\int_{y-\delta}^{y+\delta}\exp\left((4\delta)^{-1}(\mathcal{H}^{\delta}(o;w,r)+\mathcal{H}^{\delta}(w,r;q))\right)\mathrm{d}w.$
Here the first line is the Chapman-Kolmogorov equation for the SHE (e.g. see [AKQ14, Theorem 3.1]); this should be viewed as a positive temperature version of the metric composition law. In the KPZ scaling limit as $\delta\to 0$, it becomes the usual metric composition law. In particular, the above computation implies that
$\mathcal{H}^{\delta}(o;q)\geq$ $\mathcal{H}^{\delta}(o;p)+\mathcal{H}^{\delta}(p;q)+4\delta\log(2\delta)$
$+\min_{w:|w-y|\leq\delta}\mathcal{H}^{\delta}(o;w,r)-\mathcal{H}^{\delta}(o;p)+\min_{w:|w-y|\leq\delta}\mathcal{H}^{\delta}(w,r;q)-\mathcal{H}^{\delta}(p;q).$
By Theorem 7.1, we have
$\min_{w:|w-y|\leq\delta}\mathcal{H}^{\delta}(o;w,r)-\mathcal{H}^{\delta}(o;p)\to 0,\quad\min_{w:|w-y|\leq\delta}\mathcal{H}^{\delta}(w,r;q)-\mathcal{H}^{\delta}(p;q)\to 0$
in probability as $\delta\to 0$. For the second convergence, we used that $\min_{w:|w-y|\leq\delta}\mathcal{H}^{\delta}(w,r;q)-\mathcal{H}^{\delta}(p;q)$ has the same law as $\min_{w:|w-y|\leq\delta}\mathcal{H}^{\delta}(-q;-w,-r)-\mathcal{H}^{\delta}(-q;-p)$, due to the reflection symmetry of $\mathcal{H}^{\delta}$ (see e.g., [AJRAS22, Proposition 2.3]). Also note that $4\delta\log(2\delta)\to 0$ as $\delta\to 0$. Therefore we have $\mathcal{M}(o;q)\geq\mathcal{M}(o;p)+\mathcal{M}(p;q)$, i.e.  the triangle inequality holds. Then by Theorem 1.1, $\mathcal{M}$ has the same law as $\mathcal{L}$ restricted to $Q^{4}\cap\mathbb{R}^{4}_{\uparrow}$, so the conclusion follows. ∎
7.2 O’Connell-Yor polymer
The O’Connell-Yor polymer [OY01] is an integrable semi-discrete directed polymer model defined through a family of independent Brownian motions. See e.g., [O’C12, BCF14, IS16, Nic21, Vir20, LNS23] and references therein for a selection of different perspectives on the model. In particular, Borodin, Corwin, and Ferrari [BCF14] proved convergence of the one-point distributions of the O’Connell-Yor polymer starting from a certain family of initial conditions corresponding to the Baik-Ben Arous-Peché statistics from random matrix theory. Virág [Vir20] built on these results to show that the O’Connell-Yor polymer converges to the KPZ fixed point from an arbitrary initial condition. We use Theorem 1.3 to upgrade Virág’s theorem to give directed landscape convergence.
To define the O’Connell-Yor polymer, consider a collection $B=\{B_{i}\}_{i\in\mathbb{Z}}$ of independent two-sided Brownian motions. In this section only, our Brownian motions have variance $1$ rather than variance $2$ to match the conventions in the literature. For $n<m\in\mathbb{Z}$ and $x<y\in\mathbb{R}$, define the polymer free energy at inverse temperature $\beta>0$ by
$H^{\beta}_{n,m}(x,y)=\beta^{-1}\log\int_{x=x_{n}<x_{n+1}<\dots<x_{m}=y}\mathrm{d}x_{n+1}\cdots\mathrm{d}x_{m-1}\exp\left(\sum_{i=n+1}^{m}\beta[B_{i}(x_{i})-B_{i}(x_{i-1})]\right).$
We also let $H^{\beta}_{n,m}(x,y)=-\infty$ when $n<m\in\mathbb{Z}$ and $x\geq y$ or when $n=m\in\mathbb{Z}$ and $x\neq y$, and set $H^{\beta}_{n,n}(x,x)=0$ for any $n\in\mathbb{Z}$ and $x\in\mathbb{R}$.
We will consider limits of the free energy as the distance between $(x,n),(y,m)\to\infty$ along a fixed diagonal in the plane. From now on we will assume $\beta=1$ as the corresponding results for other $\beta$ can be recovered by applying a shear transformation of the model and Brownian scaling. Following [BCF14], let $\Psi(\theta)=(\log\Gamma(\theta))^{\prime}$ be the digamma function and define functions $a=a(\theta),b=b(\theta),c=c(\theta)$ by
$a=-2/\Psi^{\prime\prime},\qquad b=a\Psi^{\prime},\qquad c=\theta b-a\Psi.$
Below we fix $\theta>0$, and for $s\leq t\in(an^{3})^{-1}\mathbb{Z}$ and $x,y\in\mathbb{R}$, set
$\mathcal{H}^{n}(x,s;y,t)=n^{-1}\left(H^{1}_{asn^{3},atn^{3}}(bsn^{3}+2xn^{2},btn^{3}+2yn^{2})-c(t-s)n^{3}+2(x-y)n^{2}\Psi^{\prime}\right).$
For any other $s\leq t\in\mathbb{R}$, we let
$\mathcal{H}^{n}(x,s;y,t)=\mathcal{H}^{n}(x,(an^{3})^{-1}\lfloor asn^{3}\rfloor;y,(an^{3})^{-1}\lfloor atn^{3}\rfloor),$
and define $\mathcal{H}^{n}_{s,t}=\mathcal{H}^{n}(\cdot,s;\cdot,t)$. As with the KPZ equation, the field $\mathcal{H}^{n}_{s,t},s\leq t\in\mathbb{R}$ satisfies a positive temperature version of the metric composition law:
$\mathcal{H}^{n}_{s,t}(x,y)=n^{-1}\log\int\exp\left(n\mathcal{H}^{n}_{s,r}(x,z)+n\mathcal{H}^{n}_{r,t}(z,y)\right)\mathrm{d}z,$ (40)
for any $s\leq r\leq t\in\mathbb{R}$ and $x,y\in\mathbb{R}$.
Theorem 7.3 ([Vir20], Corollary 12).
For any fixed $s<t$, $\mathcal{H}^{n}_{s,t}$ is tight in the uniform-on-compact topology as $n\to\infty$. Moreover, any subsequential limit $\mathcal{H}^{\infty}_{s,t}$ is continuous with KPZ fixed point marginals: for any $f,g\in\operatorname{UC}$, $f\diamond\mathcal{H}^{\infty}_{s,t}\diamond g\stackrel{{\scriptstyle d}}{{=}}f\diamond\mathcal{L}_{s,t}\diamond g$.
We can upgrade this KPZ fixed point convergence to the following directed landscape convergence.
Theorem 7.4.
We have $\mathcal{H}^{n}\to\mathcal{L}$ as $n\to\infty$, in the sense that for any fixed $s_{1}<t_{1},\ldots,s_{k}<t_{k}$, we have $\{\mathcal{H}^{n}_{s_{i},t_{i}}\}_{i=1}^{k}\stackrel{{\scriptstyle d}}{{\to}}\{\mathcal{L}_{s_{i},t_{i}}\}_{i=1}^{k}$ in the uniform in compact topology.
Proof.
Let $Q\subset\mathbb{R}$ be a countable dense set containing $s_{i},t_{i}$ for each $i\in\llbracket 1,k\rrbracket$. By Theorem 7.3, $\{\mathcal{H}^{n}_{s,t}\}_{s<t\in Q}$ is tight in the uniform-on-compact topology. To prove the theorem, it suffices to show that any limit point $\{\mathcal{H}^{\infty}_{s,t}\}_{s<t\in Q}$ has the same distribution as $\{\mathcal{L}_{s,t}\}_{s<t\in Q}$. For the convenience of notations, below we denote $\mathcal{H}^{\infty}_{s,s}(x,x)=0$ and $\mathcal{H}^{\infty}_{s,s}(x,y)=-\infty$ for any $s\in Q$ and $x\neq y$.
We will apply Theorem 1.3. For this, define a family of random operators $\{\mathcal{K}_{s,t}\}_{s\leq t\in Q}$ on $\operatorname{NW}$, as follows. For $s\leq t\in Q$, and $f\in\operatorname{NW}$, we let $\mathcal{K}_{s,t}f=f\diamond\mathcal{H}^{\infty}_{s,t}$.
The operators $\mathcal{K}_{s,t}$ have KPZ fixed point marginals by Theorem 7.3, the independent increment property is inherited from the O’Connell-Yor polymer, and shift commutativity are inherited from the corresponding properties of the O’Connell-Yor polymer. It then remains to check monotonicity, i.e., condition 3 of Theorem 1.3.
From (40) and Theorem 7.3, and using the same arguments as in the proof of Theorem 7.2, we have that for any $o=(x,s)$, $p=(y,r)$, $q=(z,t)\in\mathbb{R}\times Q$ with $s<r<t$,
$\mathcal{H}^{\infty}_{s,r}(x,y)+\mathcal{H}^{\infty}_{r,t}(y,z)\leq\mathcal{H}^{\infty}_{s,t}(x,z).$ (41)
Then for any $f,g\in\operatorname{NW}$ and $s\leq r<t\in Q$ with $g\leq\mathcal{K}_{s,r}f$, we have
$\mathcal{K}_{r,t}g=g\diamond\mathcal{H}^{\infty}_{r,t}\leq\mathcal{K}_{s,r}f\diamond\mathcal{H}^{\infty}_{r,t}=f\diamond\mathcal{H}^{\infty}_{s,r}\diamond\mathcal{H}^{\infty}_{r,t}\leq f\diamond\mathcal{H}^{\infty}_{s,t}=\mathcal{K}_{s,t}f,$
where the second inequality is due to the triangle inequality (41). Thereby we get monotonicity, as stated in condition 3 of Theorem 1.3.
Now by Theorem 1.3, we conclude that there exists a directed landscape $\mathcal{L}$, such that for any $s<t\in Q$ and $x\in Q$, we have
$\mathcal{H}^{\infty}_{s,t}(x,\cdot)=\mathcal{L}_{s,t}(x,\cdot).$
As both $\mathcal{H}^{\infty}_{s,t}$ and $\mathcal{L}_{s,t}$ are continuous, we necessarily have that $\mathcal{H}^{\infty}_{s,t}=\mathcal{L}_{s,t}$. The conclusion follows. ∎
8 Exclusion processes
Finally we consider general finite range asymmetric exclusion processes (AEPs). As in the introduction, we consider a jump distribution $p$ on $\mathbb{Z}$ such that $\{v:p(v)+p(-v)>0\}$ is finite and additively generates $\mathbb{Z}$, and whose expectation satisfies $\sum_{v}vp(v)=1$. In this section, $C,c>0$ denote constants that are allowed to depend on $p$, and whose values can change from line to line.
We encode particle configurations on $\mathbb{Z}$ by functions $\eta:\mathbb{Z}\to\{0,1\}$. For each $x\in\mathbb{Z}$, $\eta(x)=1$ means there is a particle at $x$, and $\eta(x)=0$ means there is no particle at $x$. For any particle configuration $\eta$, we can write down its height function $h:\mathbb{Z}\to\mathbb{Z}$ through the rule that $h(x+1)-h(x)=2\eta(x)-1$ for any $x\in\mathbb{Z}$. For an AEP evolution $h_{t}$, as discussed in the introduction we have a choice as to the anchor point $h_{0}(0)$. We always assume that $h_{0}(0)\in 2\mathbb{Z}$.
As in [QS23], we let $\operatorname{SRW}_{\epsilon}\subset\operatorname{UC}$ be the set of functions $f:\mathbb{R}\to\mathbb{R}$ such that for each $x\in\mathbb{Z}$, $f|_{[x\epsilon/2,(x+1)\epsilon/2]}$ is linear with slope ±$\pm 2\epsilon^{-1/2}$ and $f(0)\in 2\epsilon^{1/2}\mathbb{Z}$. We call $\operatorname{SRW}_{\epsilon}$ the set of simple random walk paths at scale $\epsilon$. Then $\operatorname{SRW}:=\operatorname{SRW}_{1}$ is the state space for an AEP if we extend our AEP height functions to be linear on each of the pieces $[x,x+1],x\in\mathbb{Z}$. Let $A_{\epsilon}:\operatorname{SRW}\to\operatorname{SRW}_{\epsilon}$ be the natural rescaling map given by $A_{\epsilon}f(x)=\epsilon^{1/2}f(2\epsilon^{-1}x)$.
To put things into the framework of Theorem 1.3 and Theorem 5.2, we will encode the dynamics of a collection of coupled AEPs using a collection of random operators $\mathcal{K}^{\epsilon}=\{\mathcal{K}^{\epsilon}_{s,t}:\operatorname{SRW}_{\epsilon}\to\operatorname{SRW}_{\epsilon},s\leq t\in\mathbb{R}\}$. Specifically, suppose that we have a collection of coupled AEPs $h_{s,t}f$, $s\leq t$, defined for all $f\in\operatorname{SRW}$. For $f\in\operatorname{SRW}_{\epsilon}$ we write
$\mathcal{K}_{s,t}^{\epsilon}f=A_{\epsilon}\circ h_{2\epsilon^{-3/2}s,2\epsilon^{-3/2}t}\circ(A_{\epsilon})^{-1}f+2\epsilon^{1/2}\lfloor\epsilon^{-3/2}(t-s)/2\rfloor.$ (42)
8.1 ASEPs under various couplings
As discussed in the introduction, the nearest neighbor case where $p(v)\neq 0$ only when $v\in\{-1,1\}$ is also referred to as the asymmetric simple exclusion process (ASEP). The case where $p(1)=1,p(-1)=0$ is TASEP. For ASEP, Quastel and Sarkar [QS23] proved convergence to the KPZ fixed point.
Theorem 8.1 ([QS23, Theorem 2.2(2)]).
Consider a collection of functions $f^{\epsilon}\in\operatorname{SRW}_{\epsilon},\epsilon>0$, and $f\in\operatorname{UC}$ such that $f^{\epsilon}\to f$ in $\operatorname{UC}$. Suppose that there exists a constant $A>0$ with $f^{\epsilon}(x)\leq A(|x|+1)$ for each $\epsilon>0$ and $x\in\mathbb{R}$. Consider ASEP with jump distribution $p$. Then for any fixed $t>0$, $\mathcal{K}^{\epsilon}_{0,t}f^{\epsilon}\stackrel{{\scriptstyle d}}{{\to}}\mathfrak{h}_{t}f$ in the topology of uniform convergence on compact sets.
To describe different couplings of ASEP, we will think of the dynamics as being generated by a collection of Poisson point processes on $\mathbb{R}$, which we call Poisson clocks. A Poisson point at time $t\in\mathbb{R}$ will encode the attempt of a certain particle to make a certain jump at time $t$. The following natural couplings can be described in this way:
  • Basic coupling (colored ASEP). This coupling was previously described in the introduction. To each $x\in\mathbb{Z}$, we associate two Poisson clocks, with rate $p(1)$ and $p(-1)$ respectively. All Poisson clocks are jointly independent. The rate $p(1)$ and $p(-1)$ clocks associated to $x$ encode the jumps of particles from site $x$ to $x+1$ and $x-1$, respectively, simultaneously for all instances of ASEP.
  • Particle coupling. In each of the ASEPs, we label all the particles by integers, such that if a particle is labeled $\ell$, then the next particle to its right is labeled $\ell-1$.
    For each $\ell\in\mathbb{Z}$, we associate two Poisson clocks, with rate $p(1)$ and $p(-1)$ respectively. All such Poisson clocks are jointly independent. The rate $p(1)$ (resp. $p(-1)$) clock associated to $\ell$ encodes the right (resp. left) jumps of the particle with label $\ell$, simultaneously for all instances of ASEP.
  • Hole coupling. Instead of thinking of ASEP as particles attempting to move to the right with rate $p(1)$ and to the left with rate $p(-1)$, the model can equivalently be described as each hole attempting to move to the left with rate $p(1)$ and to the right with rate $p(-1)$, independently for all the holes. In each of the instances of ASEP, we label all the holes by integers, such that if a hole is labeled $\ell$, then the next hole to its right is labeled $\ell+1$.
    For each $\ell\in\mathbb{Z}$, we have two Poisson clocks, with rate $p(1)$ and $p(-1)$ respectively. All such Poisson clocks are jointly independent. The rate $p(1)$ (resp. $p(-1)$) clock associated to hole $\ell$ encodes the left (resp. right) jumps of this hole.
These three couplings are example of more general exotic couplings.
To define a general exotic coupling, we first introduce the following labeling of particles and holes. For any ASEP configuration, we label all particles and holes by integers via the height function. More precisely, for a particle configuration with height function $h:\mathbb{Z}\to\mathbb{Z}$, if a particle is at location $x\in\mathbb{Z}$, it is given the label $-(h(x)+x)/2$; if a hole is at location $x\in\mathbb{Z}$, it is given the label $(-h(x)+x)/2$. It is straightforward to check that this labeling rule is compatible with ASEP dynamics.
Definition 8.2.
For any $(a,b)\in\mathbb{N}^{2}\setminus\{(0,0)\}$, we define the $(a,b)$-exotic coupling as follows. Let $\mathbb{Z}^{2}/(a,b)$ be the quotient of $\mathbb{Z}^{2}$ under addition by $(a,b)$. We associate each element $x\in\mathbb{Z}^{2}/(a,b)$ to two Poisson clocks with rate $p(1)$ and $p(-1)$, respectively. Let $Q:\mathbb{Z}^{2}\to\mathbb{Z}^{2}/(a,b)$ denote the projection map. Any ASEP under the $(a,b)$-exotic coupling is generated by these Poisson clocks as follows. For any $\ell,k\in\mathbb{Z}$, if the particle with $\ell$ sits immediate to the left (resp. right) of the hole with label $k$, then their swap is encoded by the rate $p(1)$ (resp. $p(-1)$) clock associated to $Q(\ell,k)$.
The basic, particle, and hole couplings are the $(1,1)$-, $(1,0)$-, and $(0,1)$-exotic couplings, respectively (see Figure 1).
-4-4-3-3-2-2-1-1001122334455660/1/2/3/4/5/6/7/8/9/10/1234567891011-1/-2/-3/-4/-5/-6/-7/-8/-9/-10/-2-3-4-5-6-7-8-9-10-11particle labelhole labelinteraction site ($x/x+1$)
Figure 1: An illustration of the Poisson clocks in exotic couplings: under the (1,0)-exotic coupling, Poisson clocks in the same row are taken to be the same, giving the particle coupling; under the (0,1)-exotic coupling, Poisson clocks in the same column are taken to be the same, giving the hole coupling; under the (1,1)-exotic coupling, Poisson clocks in the same diagonal (corresponding to the same interaction site) are taken to be the same, giving the basic coupling.
Lemma 8.3.
For $\mathcal{K}^{\epsilon}$ coming from any exotic coupling, any $s<t$, and any fixed initial conditions $f\leq g\in\operatorname{SRW}_{\epsilon}$, we have $\mathcal{K}^{\epsilon}_{s,t}f\leq\mathcal{K}^{\epsilon}_{s,t}g$.
Proof.
It is straight forward to check that, for any two particle configurations under any exotic coupling, if the associated height functions $h^{+},h^{-}:\mathbb{Z}\to\mathbb{Z}$ are ordered, i.e., $h^{+}(x)\geq h^{-}(x)$ for any $x\in\mathbb{Z}$, at some time, then after any possible jump the same ordering holds. The conclusion follows by induction. ∎
Our main result on ASEP is the following.
Theorem 8.4.
Take $f_{1},\ldots,f_{k}\in\operatorname{UC}$ and $f_{1}^{\epsilon},\ldots,f_{k}^{\epsilon}\in\operatorname{SRW}_{\epsilon}$ for each $\epsilon>0$, such that $f_{i}^{\epsilon}\to f_{i}$ in $\operatorname{UC}$ for each $i\in\llbracket 1,k\rrbracket$. Suppose also that for some $A>0$, we have $f_{i}^{\epsilon}(x)\leq A(1+|x|)$ for all $i\in\llbracket 1,k\rrbracket$, $\epsilon>0$, and $x\in\mathbb{R}$.
Consider operators $\mathcal{K}^{\epsilon}=\{\mathcal{K}^{\epsilon}_{s,t}:\operatorname{SRW}_{\epsilon}\to\operatorname{SRW}_{\epsilon},s<t\in\mathbb{R}\}$ defined through $(a_{\epsilon},b_{\epsilon})$-exotic couplings of ASEP with a common jump distribution $p$. Suppose $(a_{\epsilon},b_{\epsilon})\in\mathbb{N}^{2}\setminus\{(0,0)\}$ satisfies $a_{\epsilon}\epsilon^{1/2}\to 0$ and $b_{\epsilon}\epsilon^{1/2}\to 0$ as $\epsilon\to 0$. Then for any finite $T\subset(0,\infty)$, as $\epsilon\to 0$,
$\Big{\{}\mathcal{K}^{\epsilon}_{0,t}f_{i}^{\epsilon}\Big{\}}_{i\in\llbracket 1,k\rrbracket,t\in T}\stackrel{{\scriptstyle d}}{{\to}}\Big{\{}f_{i}\diamond\mathcal{L}_{0,t}\Big{\}}_{i\in\llbracket 1,k\rrbracket,t\in T},$
in the topology of uniform convergence on compact sets.
To prove Theorem 8.4 using Theorem 1.3, we need to construct a limiting family of random operators on $\operatorname{NW}$.
For any $f\in\operatorname{NW}$ and $\epsilon>0$, we define $f^{\epsilon}$ to be the minimal function in $\operatorname{SRW}_{\epsilon}$ satisfying $f\leq f^{\epsilon}$. Such a minimal function exists since if $g_{1},g_{2}\in\operatorname{SRW}_{\epsilon}$ then so is $g_{1}\wedge g_{2}$. Observe that $f^{\epsilon}\downarrow f$ in $\operatorname{UC}$ as $\epsilon\to 0$ and each of $f^{\epsilon}$, $\epsilon\in(0,1)$, is bounded above by $\sup f+1$, and so satisfies the convergence conditions of Theorem 8.1.
Since $\operatorname{NW}$ is countable, Theorem 8.1 guarantees that the collection $\{\mathcal{K}^{\epsilon}_{s,t}f^{\epsilon}\}_{s\leq t\in\mathbb{Q},f\in\operatorname{NW}}$ is tight as $\epsilon\to 0$, in the product of uniform-on-compact topologies.
Lemma 8.5.
Let $\{\mathcal{K}_{s,t}f\}_{s\leq t\in\mathbb{Q},f\in\operatorname{NW}}$ be any subsequential limit in law of $\{\mathcal{K}^{\epsilon}_{s,t}f^{\epsilon}\}_{0\leq s\leq t\in\mathbb{Q},f\in\operatorname{NW}}$ as $\epsilon\to 0$. Then $\mathcal{K}$ satisfies shift commutativity, i.e., for all $s\leq t\in\mathbb{Q},f\in\operatorname{NW},c\in\mathbb{Q}$, $\mathcal{K}_{s,t}(f+c)=\mathcal{K}_{s,t}f+c$.
Proof.
Fix any $f\in\operatorname{NW},c\in\mathbb{Q}$, and $s\leq t\in\mathbb{Q}$. For any $\epsilon>0$, from the definition of the $(a_{\epsilon},b_{\epsilon})$-exotic coupling and the construction of $f^{\epsilon}$ we have that
$\mathcal{K}^{\epsilon}_{s,t}(f^{\epsilon}+c_{\epsilon})(x)=\mathcal{K}^{\epsilon}_{s,t}f^{\epsilon}(x-c_{\epsilon}(a_{\epsilon}+b_{\epsilon})^{-1}(a_{\epsilon}-b_{\epsilon})\epsilon^{1/2}/2)+c_{\epsilon},$ (43)
for any $c_{\epsilon}\in(a_{\epsilon}+b_{\epsilon})\epsilon^{1/2}\mathbb{Z}$ and $x\in\mathbb{R}$. We now take $c_{\epsilon}^{+}=(a_{\epsilon}+b_{\epsilon})\epsilon^{1/2}\lceil c(a_{\epsilon}+b_{\epsilon})^{-1}\epsilon^{-1/2}\rceil$ and $c_{\epsilon}^{-}=(a_{\epsilon}+b_{\epsilon})\epsilon^{1/2}\lfloor c(a_{\epsilon}+b_{\epsilon})^{-1}\epsilon^{-1/2}\rfloor$. By Lemma 8.3, we have that
$\mathcal{K}^{\epsilon}_{s,t}(f^{\epsilon}+c_{\epsilon}^{-})\leq\mathcal{K}^{\epsilon}_{s,t}(f+c)^{\epsilon}\leq\mathcal{K}^{\epsilon}_{s,t}(f^{\epsilon}+c_{\epsilon}^{+}).$
Now, using (43) we have
$\mathcal{K}^{\epsilon}_{s,t}f^{\epsilon}(x-c_{\epsilon}^{-}(a_{\epsilon}+b_{\epsilon})^{-1}(a_{\epsilon}-b_{\epsilon})\epsilon^{1/2}/2)+c_{\epsilon}^{-}\leq\mathcal{K}^{\epsilon}_{s,t}(f+c)^{\epsilon}(x)\leq\mathcal{K}^{\epsilon}_{s,t}f^{\epsilon}(x-c_{\epsilon}^{+}(a_{\epsilon}+b_{\epsilon})^{-1}(a_{\epsilon}-b_{\epsilon})\epsilon^{1/2}/2)+c_{\epsilon}^{+},$
for any $x\in\mathbb{R}$. Taking $\epsilon\to 0$ along the given subsequence, we have that $\mathcal{K}^{\epsilon}_{s,t}f^{\epsilon}\to\mathcal{K}_{s,t}f$ in the uniform-on-compact topology, $(a_{\epsilon}+b_{\epsilon})^{-1}(a_{\epsilon}-b_{\epsilon})\epsilon^{1/2}/2\to 0$, and $c_{\epsilon}^{+},c^{\epsilon}_{-}\to c$. Therefore since $\mathcal{K}_{s,t}f$ is continuous, the function $x\mapsto\mathcal{K}^{\epsilon}_{s,t}f^{\epsilon}(x-c_{\epsilon}^{\pm}(a_{\epsilon}+b_{\epsilon})^{-1}(a_{\epsilon}-b_{\epsilon})\epsilon^{1/2}/2)+c_{\epsilon}^{\pm}$ converges to $\mathcal{K}_{s,t}f+c$ (in distribution, uniformly on compact sets as $\epsilon\to 0$ along the subsequence). Therefore $\mathcal{K}_{s,t}(f+c)=\mathcal{K}_{s,t}f+c$. ∎
Proof of Theorem 8.4.
Without loss of generality, we assume that condition (7) holds for each $f_{i}$. Otherwise we can replace $\mathbb{Q}$ in our definition of $\operatorname{NW}$ with a countable dense set $Q\subset\mathbb{R}$ tailored so that (7) holds for each $f_{i}$, as remarked below Corollary 1.5.
From the convergence of $\mathcal{K}^{\epsilon}_{0,t}f_{i}^{\epsilon}$ for each $i\in\llbracket 1,k\rrbracket$, given by Theorem 8.1, we have that $\{\mathcal{K}^{\epsilon}_{0,t}f_{i}^{\epsilon}\}_{i\in\llbracket 1,k\rrbracket,t\in T}$ is tight as $\epsilon\to 0$, in the product of uniform-on-compact topologies. We then take a joint subsequential limit of $\{\mathcal{K}^{\epsilon}_{0,t}f_{i}^{\epsilon}\}_{i\in\llbracket 1,k\rrbracket,t\in T}$ and $\{\mathcal{K}^{\epsilon}_{s,t}f^{\epsilon}\}_{s\leq t\in\mathbb{Q},f\in\operatorname{NW}}$ as $\epsilon\to 0$, denoted by $\{\tilde{\mathcal{K}}_{0,t}f_{i}\}_{i\in\llbracket 1,k\rrbracket,t\in T}$ and $\{\mathcal{K}_{s,t}f:s\leq t\in\mathbb{Q},f\in\operatorname{NW}\}$.
By Theorem 8.1, the independence of the underlying Poisson clocks, Lemma 8.3, and Lemma 8.5, $\{\mathcal{K}_{s,t}f\}_{0\leq s\leq t\in\mathbb{Q}}$ satisfies four conditions specified in Theorem 1.3. Therefore there exists a directed landscape $\mathcal{L}$ such that almost surely $\mathcal{K}_{s,t}f=f\diamond\mathcal{L}_{s,t}$ for any $s\leq t\in\mathbb{Q}$ and $f\in\operatorname{NW}$.
Now for each $f_{i}$ and $f\in\operatorname{NW}$, such that $f_{i}(x)>f(x)$ for all $x\in\mathbb{R}$, from the convergence of $f^{\epsilon}_{i}\to f_{i}$ and $f^{\epsilon}\to f$ and the minimal choice of the approximators $f^{\epsilon}$, we have $f^{\epsilon}_{i}\geq f^{\epsilon}$ for all small enough $\epsilon$. Then for each $t\in T$, by Lemma 8.3 we have $\mathcal{K}^{\epsilon}_{0,t}f^{\epsilon}_{i}\geq\mathcal{K}^{\epsilon}_{0,t}f^{\epsilon}$ for all small enough $\epsilon$, thus $\mathcal{K}_{0,t}f^{\epsilon}_{i}\geq\mathcal{K}_{0,t}f^{\epsilon}$. By Corollary 1.5 the conclusion follows. ∎
8.2 AEPs under the basic coupling
In the remainder of this section, we consider general AEPs with the basic coupling. The setup is described in the introduction and at the beginning of Section 8. In this subsection we will prove joint convergence of AEPs to the directed landscape by appealing to Theorem 5.2. One technical claim verifying approximate monotonicity for the model is left to Section 8.3.
Throughout the section, we fix a valid jump distribution $p$, and let $\mathcal{K}^{\epsilon}=\{\mathcal{K}^{\epsilon}_{s,t}:\operatorname{SRW}_{\epsilon}\to\operatorname{SRW}_{\epsilon},s\leq t\in\mathbb{R}\}$ denote the rescaled basic coupling of AEP with this jump distribution, via (42).
As with ASEPs, KPZ fixed point convergence for general AEPs was shown in [QS23]. The convergence result applies to initial conditions which are not too far from the stationary Bernoulli product measure. We let $\nu$ be the law on $\operatorname{SRW}$ of a simple symmetric random walk path $X$ with $X(0)=0$. We let $\nu_{\epsilon}$ be the pushforward of $\nu$ under the natural map $A_{\epsilon}:\operatorname{SRW}\to\operatorname{SRW}_{\epsilon}$.
Theorem 8.6.
Fix $s<t\in\mathbb{R}$, and suppose that for all $\epsilon>0$, we have a random initial condition $f^{\epsilon}$ with law $g_{\epsilon}d\nu_{\epsilon}$, such that $f^{\epsilon}$ is independent of $\mathcal{K}^{\epsilon}_{s,t}$. Additionally assume that:
  • As $\epsilon\to 0$, $g_{\epsilon}d\nu_{\epsilon}$ converges weakly in $\operatorname{UC}$ to a limit law $\mu$ on $\operatorname{UC}_{0}$.
  • $\lim_{\epsilon\to 0}\epsilon^{1/4}\|g_{\epsilon}d\nu_{\epsilon}\|_{L^{2}(\nu_{\epsilon})}=0$.
  • Define random variables
    $M_{\epsilon}=\inf\{a\in\mathbb{R}:f^{\epsilon}(x)<a(1+|x|)\text{ for all }x\in\mathbb{R}\}.$
    Then $M_{\epsilon},\epsilon>0$ is a tight sequence of random variables.
Then $(f^{\epsilon},\mathcal{K}^{\epsilon}_{s,t}f^{\epsilon})\stackrel{{\scriptstyle d}}{{\to}}(f,\mathfrak{h}_{t-s}f)$, where $f\sim\mu$ is independent of the KPZ fixed point evolution. Here the underlying topology is $\operatorname{UC}$-convergence of $f^{\epsilon}$ to $f$ and uniform-on-compact convergence of $\mathcal{K}^{\epsilon}_{s,t}f^{\epsilon}$ to $\mathfrak{h}_{t-s}f$.
Proof of Theorem 8.6.
The convergence of $\mathcal{K}^{\epsilon}_{s,t}f^{\epsilon}\stackrel{{\scriptstyle d}}{{\to}}\mathfrak{h}_{t-s}f$ is Theorem 2.2(1) in [QS23]. For the joint convergence, let $(f,\mathcal{K}_{s,t}f)$ be any joint subsequential limit in distribution of $(f^{\epsilon},\mathcal{K}^{\epsilon}_{s,t}f^{\epsilon})$. Let $U$ be any continuity set for the law of $f$ such that $\mathbb{P}[f\in U]>0$. Then $\mathbb{P}[f^{\epsilon}\in U]\to\mathbb{P}[f\in U]$ as $n\to\infty$. Moreover, since $\mathbb{P}[f\in U]>0$, if $f^{\epsilon}_{U}$ has the law of $f^{\epsilon}$ given $f^{\epsilon}\in U$, then $f^{\epsilon}_{U}$ still satisfies the three bullets above and hence the conclusion of Theorem 2.2(1) in [QS23]. Therefore for any measurable $A\subset\operatorname{UC}$, we have $\mathbb{P}[(f,\mathcal{K}_{s,t}f)\in U\times A]=\mathbb{P}[(f,\mathfrak{h}_{t-s}f)\in U\times A]$, and so $(f,\mathcal{K}_{s,t}f)\stackrel{{\scriptstyle d}}{{=}}(f,\mathfrak{h}_{t-s}f)$ as desired. ∎
We upgrade Theorem 8.6 to directed landscape convergence.
Theorem 8.7.
Take (potentially random) initial conditions $f_{0},\ldots,f_{k}\in\operatorname{UC}$ and $f_{0}^{\epsilon},\ldots,f_{k}^{\epsilon}\in\operatorname{SRW}_{\epsilon}$ for each $\epsilon>0$, such that $f_{i}^{\epsilon}\stackrel{{\scriptstyle d}}{{\to}}f_{i}$ for each $i\in\llbracket 0,k\rrbracket$. Suppose that $f_{0}^{\epsilon}\sim\nu_{\epsilon}$ and for every $i\in\llbracket 1,k\rrbracket$, the initial conditions $f_{i}$ satisfy the first two bullet points of Theorem 8.6 and that the random variables
$\sup(f_{i}^{\epsilon}-f_{0}^{\epsilon}),\epsilon>0$
are tight. Then for any finite $T\subset(0,\infty)$, as $\epsilon\to 0$,
$\{\mathcal{K}^{\epsilon}_{0,t}f_{i}^{\epsilon}\}_{t\in T,i\in\llbracket 1,k\rrbracket}\stackrel{{\scriptstyle d}}{{\to}}\{f_{i}\diamond\mathcal{L}_{0,t}\}_{t\in T,i\in\llbracket 1,k\rrbracket},$
in the topology of uniform convergence on compact sets.
We prove Theorem 8.7 by appealing to the framework of Theorem 5.2. For this, we first need to construct prelimiting initial conditions for each of the sets $\mathscr{H}_{s}$. Fix $\epsilon>0$. We start by constructing the prelimiting versions $B_{s,n}^{\epsilon}$ of the $B_{s,n}$, which will be Brownian motions in the limit.
For all $s\in\mathbb{Q}_{+}$, define $B_{s,1}^{\epsilon}=\mathcal{K}_{0,s}^{\epsilon}f_{0}^{\epsilon}-\mathcal{K}_{0,s}^{\epsilon}f_{0}^{\epsilon}(0)$. Each function $B_{s,1}^{\epsilon}$ is thus distributed according to the rescaled simple random walk measure $\nu_{\epsilon}$. Then, we define $B_{s,n}^{\epsilon}$, $n\geq 2$ to be a sequence of simple random walks at scale $\epsilon$, by letting $B_{s,n}^{\epsilon}\sim\nu_{\epsilon}$ conditional on $B_{s,n}^{\epsilon}|_{[-n,n]^{c}}=B_{s,1}^{\epsilon}|_{[-n,n]^{c}}$. In other words, (conditional on $B_{s,1}^{\epsilon}$) $B_{s,n}^{\epsilon}|_{[-n,0]}$ and $B_{s,n}^{\epsilon}|_{[0,n]}$ are independent random walk bridges at scale $\epsilon$.
Next, recall the set of piecewise linear functions $\mathscr{H}$ defined prior to Theorem 5.2. Let $\{U_{i}\}_{i\in\mathbb{Z}}$ be a collection of i.i.d. uniform random variables on $[0,1]$. For a function $h\in\mathscr{H}$, we define an approximation to $B_{s,n}^{\epsilon}+h$ using the random variables $U_{i}$ as follows. First, since $B_{s,n}^{\epsilon}\sim\nu_{\epsilon}$ we can extract a sequence of i.i.d. Rademacher random variables $e=\{e_{i}\}_{i\in\mathbb{Z}}$ by letting
$e_{i}=\epsilon^{-1/2}(B_{s,n}^{\epsilon}(i\epsilon/2)-B_{s,n}^{\epsilon}((i-1)\epsilon/2)).$
The map $B_{s,n}^{\epsilon}\mapsto e$ is bijective, so we can think of $B_{s,n}^{\epsilon}=g(e)$ for some map $g:\{\pm 1\}^{\mathbb{Z}}\to\operatorname{SRW}_{\epsilon}$. For $h\in\mathscr{H}$ with $h(0)=0$, we will define a sequence of random variables $v(h)=\{v(h)_{i}\}_{i\in\mathbb{Z}}$, and let $B_{s,n}^{\epsilon}[h]:=g(v(h))$.
Let $\{q_{1}<\cdots<q_{k}\}$ be the set of points where $h^{\prime}$ does not exist, so that $h^{\prime}=0$ off of $[q_{1},q_{k}]$ and is constant on each interval $(q_{j},q_{j+1})$. For $i\epsilon/2\notin(q_{1},q_{k}]$ we set $v(h)_{i}=e_{i}$. Now, for each $j\in\{1,\dots,k-1\}$, let $s_{j}$ be the value of $h^{\prime}$ on $(q_{j},q_{j+1})$. For $i\epsilon/2\in(q_{j},q_{j+1}]$ we let
$v(h)_{i}=\begin{cases}&\max(e_{i},2\cdot\mathbf{1}(U_{i}\in[0,s_{j}\epsilon/2])-1),\qquad s_{j}\geq 0,\\ &\min(e_{i},1-2\cdot\mathbf{1}(U_{i}\in[0,|s_{j}|\epsilon/2])),\qquad s_{j}<0.\end{cases}$
It is straightforward to check (i.e. using the law of large numbers) that, for $B_{s,n}^{\epsilon}[h]:=g(v(h))$, we have
$B_{s,n}^{\epsilon}[h]-B_{s,n}^{\epsilon}\to h$ (44)
uniformly in probability as $\epsilon\to 0$. For functions $h\in\mathscr{H}$ with $h(0)\neq 0$, we define
$B_{s,n}^{\epsilon}[h]:=B_{s,n}^{\epsilon}[h-h(0)]+2\epsilon^{1/2}\lfloor\epsilon^{-1/2}h(0)/2\rfloor.$
With this setup, we have the following straightforward tightness lemma.
Lemma 8.8.
For any fixed $s<t\in\mathbb{Q}_{+}$, $n\in\mathbb{N}$ and $h\in\mathscr{H}$, we have that
$(B_{s,n}^{\epsilon}[h],\mathcal{K}_{s,t}^{\epsilon}B_{s,n}^{\epsilon}[h])\stackrel{{\scriptstyle d}}{{\to}}(B+h,\mathfrak{h}_{t-s}(B+h)),$
in the uniform-on-compact topology, where $B$ is a Brownian motion (of variance $2$), and the KPZ fixed point evolution $\mathfrak{h}$ is independent of $B$. Moreover, if we let
$(B_{s,n,h}+h,\mathcal{K}_{s,t}(B_{s,n,h}+h))_{s<t\in\mathbb{Q}_{+},n\in\mathbb{N},h\in\mathscr{H}}$ (45)
be any subsequential limit in law of $(B_{s,n}^{\epsilon}[h],\mathcal{K}_{s,t}^{\epsilon}B_{s,n}^{\epsilon}[h])_{s<t\in\mathbb{Q}_{+},n\in\mathbb{N},h\in\mathscr{H}}$, then $B_{s,n,h}=B_{s,n,h^{\prime}}$ for all $h,h^{\prime}$. Henceforth we drop the subscript $h$. This collection satisfies conditions $1,2,$ and $5$ of Theorem 5.2, along with the tail-triviality of the sequences $\{B_{s,n}\}_{n\in\mathbb{N}}$.
Proof.
For the first claim in the lemma, we have that $(h+B_{s,n})^{\epsilon}\stackrel{{\scriptstyle d}}{{\to}}h+B$ by construction. To check the joint limit claim, it suffices to consider the case when $h(0)=0$, since both $\mathfrak{h}_{t}$ and $\mathcal{K}_{s,t}^{\epsilon}$ are shift commutative (the latter by the definition of the basic coupling for AEPs). We will apply Theorem 8.6. The first and third bullet points are straightforward using standard bounds on random walks, convergence of random walks to Brownian motion, and the law of large numbers to handle the augmentation by $h$. For the second bullet point, we claim that if $g_{\epsilon}d\nu_{\epsilon}$ is the law of $B_{s,n}^{\epsilon}[h]$, then $\|g_{\epsilon}d\nu_{\epsilon}\|_{L^{2}(\nu_{\epsilon})}$ stays bounded as $\epsilon\to 0$.
Consider $F_{\epsilon}=A_{\epsilon}^{-1}(B_{s,n}^{\epsilon}[h])\in\operatorname{SRW}$, and let $h_{\epsilon}$ be the Radon-Nikodym derivative of its law with respect to $\nu$. Equivalently, we will show that $\|h_{\epsilon}d\nu\|_{L^{2}(\nu)}$ stays bounded as $\epsilon\to 0$. By construction, there is an integer set $Q_{\epsilon}=\{q_{1,\epsilon},\dots,q_{k,\epsilon}\}$ (up to $O(\epsilon)$-errors, this set is given by $2\epsilon^{-1}\{q_{1},\dots,q_{k}\}$ with the $q_{i}$ as above) such that conditional on $F_{\epsilon}|_{Q_{\epsilon}}$, $F_{\epsilon}$ has the same law as a simple symmetric random walk path conditioned on the values on $Q_{\epsilon}$. This follows from the fact that if we condition a parameter-$p$ Bernoulli process on $\{1,\dots,n\}$ to have exactly $k$ $1$s, then under this conditioning the $k$ $1$s are uniformly distributed amongst all $k$-element subsets of $\{1,\dots,n\}$; in particular, the parameter $p$ plays no role.
Therefore $\|h_{\epsilon}d\nu\|_{L^{2}(\nu)}=\|h_{\epsilon}^{\prime}d\nu|_{F}\|_{L^{2}(\nu|_{Q_{\epsilon}})}$, where $\nu|_{Q_{\epsilon}}$ is the pushforward of $\nu$ under the projection map $f\mapsto f|_{Q_{\epsilon}}$, and $h_{\epsilon}^{\prime}$ is the Radon-Nikodym derivative of $F_{\epsilon}|_{Q_{\epsilon}}$ against this law. A uniform in $\epsilon$ bound on $\|h_{\epsilon}^{\prime}d\nu|_{F}\|_{L^{2}(\nu|_{Q_{\epsilon}})}$ can then be checked by a computation comparing the laws of binomial random variables whose means differ on the order of their standard deviations.
For the claims regarding the joint limit (45), the fact that $B_{s,n,h}=B_{s,n,h^{\prime}}$ for all $h,h^{\prime}\in\mathscr{H}$ follows from (44). Condition $1$ in Theorem 5.2 (KPZ fixed point marginals) is established above. The conditional independence (condition $2$) and shift commutativity (condition $5$) are immediate from the corresponding properties in the prelimit, which follow from the construction of the basic coupling. Tail-triviality of the sequences $\{B_{s,n}\}_{n\in\mathbb{N}}$ is immediate from the construction of $B_{s,n}^{\epsilon}$, which implies that the collection $\{B_{s,n}\}_{n\in\mathbb{N}}$ consists of Brownian motions which are conditionally independent given the equalities $B_{s,n}(x)=B_{s,1}(x),n\in\mathbb{N},|x|\geq n$. ∎
The final two assumptions in Theorem 5.2 require a monotonicity claim for general AEPs. This monotonicity claim will also allow us to connect the field in (45) to the evolutions in Theorem 8.7.
For this next lemma, consider two AEPs with height functions $(h^{-}_{t})_{t\geq 0}$ and $(h^{+}_{t})_{t\geq 0}$, started at time 00 in the basic coupling. Assume that $h^{-}_{0}\leq h^{+}_{0}$. For any $t,w>0$, we denote
$Y_{t,w}=\max_{x\in\llbracket-w,w\rrbracket}h^{-}_{t}(x)-h^{+}_{t}(x).$
Lemma 8.9.
For any $m,t,w>0$, we have $\mathbb{P}[Y_{t,w}>m]<C(t+w)\exp(-cm)+C\exp(-ct)$.
We postpone the proof of Lemma 8.9 to the next section. To pass it to the limit, we need the following simple estimate on AEPs in the basic coupling.
Lemma 8.10.
Suppose that $\eta^{1}_{t},\eta^{2}_{t}$ are two AEPs with the basic coupling started from initial conditions $\eta^{1}_{0},\eta^{1}_{0}$ that agree off of a set $\{-n,\dots,n\}$. Then there exists $C_{1},C_{2}>0$ such that for all $t>0$,
$\mathbb{P}\left[\eta^{1}_{t}(x)=\eta^{2}_{t}(x)\text{ for all }x\notin[-C_{1}t-n,C_{1}t+n]\right]\geq 1-\exp(-C_{2}t).$
Proof.
Let $M_{t}=\sup\{m\in\mathbb{N}:(\eta^{1}_{t}(-m),\eta^{1}_{t}(m))\neq(\eta^{2}_{t}(-m),\eta^{2}_{t}(m))\}$. Then the process $(M_{t},\eta^{1}_{t},\eta^{2}_{t})$ is jointly Markov, and since the jump distribution $p$ has finite support, the process $M_{t}$ is stochastically dominated by $N_{t}+n$, where $N_{t},t\geq 0$ is a Poisson counting process with some $p$-dependent rate $C_{1}/2>0$. The bound in the lemma then follows from a standard tail estimate on Poisson random variables. ∎
With these two lemmas, we can verify the remaining two conditions required by Theorem 5.2.
Lemma 8.11.
Let all notations be as in Theorem 8.7 and Lemma 8.8.
  1. 1.
    (Comparability) Let $s\leq r<t\in\mathbb{Q}_{+}$, and $n,n^{\prime}\in\mathbb{N}$, $h,h^{\prime}\in\mathscr{H}$. Then for $(f,g)=(B_{s,n}+h,B_{r,n^{\prime}}+h^{\prime})$ from (45), we have that
    $\sup(g-\mathcal{K}_{s,r}f)<\infty.$
  2. 2.
    (Monotonicity) In the setup of part $1$, if $g\leq\mathcal{K}_{s,r}f$ then $\mathcal{K}_{r,t}g\leq\mathcal{K}_{s,t}f.$
  3. 3.
    For any finite set $T_{*}\subset\mathbb{Q}_{+}$, consider a joint subsequential limit $(\tilde{\mathcal{K}}_{0,t}f_{i})_{t\in T_{*},i\in\llbracket 1,k\rrbracket}$ of $(\mathcal{K}_{0,t}^{\epsilon}f^{\epsilon}_{i})_{t\in T_{*},i\in\llbracket 1,k\rrbracket}$ along with (45). Then for each $i\in\llbracket 1,k\rrbracket$ and $t_{*}=\min T_{*}$, we can find a (random) sequence $\{h_{m}\}_{m\in\mathbb{N}}\in\mathscr{H}$ such that
    $B_{t_{*},1}+h_{m}\downarrow\tilde{\mathcal{K}}_{0,t_{*}}f_{i},\qquad$ as $m\to\infty$, almost surely in $\operatorname{UC}$,
    $\mathcal{K}_{t_{*},t}(B_{t_{*},1}+h_{m})\geq\tilde{\mathcal{K}}_{0,t}f_{i}\qquad$ for $\text{for each }m\in\mathbb{N},t\in T_{*},t>t_{*}.$
Proof.
For part $1$, since $\sup(g-B_{r,1})<\infty$, it suffices to show that $\sup(B_{r,1}-\mathcal{K}_{s,r}f)<\infty$. Moreover, $B_{r,1}=\mathcal{K}_{s,r}B_{s,1}-\mathcal{K}_{s,r}B_{s,1}(0)$ by definition, so it suffices to show that $\sup(\mathcal{K}_{s,r}B_{s,1}-\mathcal{K}_{s,r}f)<\infty$.
For $\epsilon>0$, let $Q_{\epsilon}=1+\lceil\sup B^{\epsilon}_{s,1}-B^{\epsilon}_{s,n}[h]\rceil$. The sequence $Q_{\epsilon},\epsilon>0$ is tight. Let $Q$ be a subsequential limit of $Q_{\epsilon}$ taken jointly with (45). We will use Lemma 8.9 and Lemma 8.10 to show that
$\mathcal{K}_{s,r}B_{s,1}\leq\mathcal{K}_{s,r}(f+Q)=\mathcal{K}_{s,r}f+Q,$ (46)
from which the claim follows. Here the equality above uses the shift commutativity established in Lemma 8.8. Observe that $B^{\epsilon}_{s,1}\leq B^{\epsilon}_{s,n}[h+Q_{\epsilon}]$. Therefore
$\sup\mathcal{K}_{s,r}^{\epsilon}B_{s,1}^{\epsilon}-\mathcal{K}_{s,r}^{\epsilon}B_{s,n}^{\epsilon}[h+Q_{\epsilon}]$ $=N_{\epsilon}+\max_{x\in[\epsilon^{-2},\epsilon^{-2}]}\mathcal{K}_{s,r}^{\epsilon}B_{s,1}^{\epsilon}(x)-\mathcal{K}_{s,r}^{\epsilon}B_{s,n}^{\epsilon}[h+Q_{\epsilon}](x)$
$\leq N_{\epsilon}+\epsilon^{1/2}\log(\epsilon^{-1})M_{\epsilon},$
where $M_{\epsilon},\epsilon>0$ is a tight collection of random variables and $N_{\epsilon}\stackrel{{\scriptstyle\mathbb{P}}}{{\to}}0$ with $\epsilon$. Here the convergence of $N_{\epsilon}$ to 00 uses Lemma 8.10 and the tightness of $M_{\epsilon}$ uses Lemma 8.9. The inequality (46) follows by taking $\epsilon\to 0$.
For part $2$, note that the above arguments actually imply the stronger claim that the collection of random variables
$(S^{\epsilon})_{\epsilon>0}:=(\sup(B_{r,n^{\prime}}^{\epsilon}[h^{\prime}]-\mathcal{K}_{s,r}^{\epsilon}B_{s,n}^{\epsilon}[h]))_{\epsilon>0}$
is tight. If we take any subsequential limit $S$, jointly with (45), then by applying Lemmas 8.9 and 8.10 as before we would have that $\mathcal{K}_{r,t}g\leq\mathcal{K}_{s,t}f+S$. This would imply part $2$ if we could show that $S=\sup(g-\mathcal{K}_{s,r}f)$. However, this is not obvious and so we work around this difficulty with a truncation argument.
For each $m\in\mathbb{N}$, define $w_{m}\in\mathscr{H}$ by letting $w_{m}(x)=-1/m$ when $x\in[-m,m]$, and $w_{m}(x)=-m$ when $x\in[-m-1,m+1]^{c}$, and letting $w_{m}$ be linear on $[-m-1,-m]$ and $[m,m+1]$. Then similarly, the collection of random variables
$(S^{\epsilon}_{m})_{\epsilon>0}:=(\sup(B_{r,n^{\prime}}^{\epsilon}[h^{\prime}+w_{m}]-\mathcal{K}_{s,r}^{\epsilon}B_{s,n}^{\epsilon}[h]))_{\epsilon>0}$
is also tight for fixed $m$. Let $(S_{m})_{m\in\mathbb{N}}$ be a subsequential limit of $(S^{\epsilon}_{m})_{m\in\mathbb{N}}$ as $\epsilon\to 0$, jointly with (45).
Using the tightness of $(S^{\epsilon})_{\epsilon>0}$, almost surely $S_{m}=\sup(g+w_{m}-\mathcal{K}_{s,r}f)$ for all $m$ large enough, and thus $S_{m}\to\sup(g-\mathcal{K}_{s,r}f)$ as $m\to\infty$.
Our approximations ensure that for all $m\in\mathbb{N}$ and any $\gamma\in\mathbb{Q}$, $\gamma>S_{m}$, we have
$B_{r,n^{\prime}}^{\epsilon}[h^{\prime}+w_{m}]\leq B_{r,n^{\prime}}^{\epsilon}[h^{\prime}],\qquad B_{r,n^{\prime}}^{\epsilon}[h^{\prime}+w_{m}]\leq\mathcal{K}_{s,r}^{\epsilon}B_{s,n}^{\epsilon}[h+\gamma],$
with probability $\to 1$ as $\epsilon\to 0$. Then by using Lemmas 8.9 and 8.10 as before we have that
$\mathcal{K}_{r,t}(g+w_{m})\leq\mathcal{K}_{r,t}g,\qquad\mathcal{K}_{r,t}(g+w_{m})\leq\mathcal{K}_{s,t}(f+\gamma).$ (47)
Now, as $m\to\infty$, $g+w_{m}\stackrel{{\scriptstyle d}}{{\to}}g$ and so by 2.4, $\mathcal{K}_{r,t}(g+w_{m})\stackrel{{\scriptstyle d}}{{\to}}\mathcal{K}_{r,t}g$. By the first inequality in (47) and Lemma 5.1, this convergence must also hold in probability. On the other hand, using shift commutativity of $\mathcal{K}_{s,t}$, the second inequality in (47) implies that $\mathcal{K}_{r,t}(g+w_{m})\leq\mathcal{K}_{s,t}f+S_{m}$. By sending $m\to\infty$ we have that
$\mathcal{K}_{r,t}g\leq\mathcal{K}_{s,t}f+\sup(g-\mathcal{K}_{s,r}f),$
which yields part $2$.
For part $3$, we let $R^{\epsilon}=\sup(f_{i}^{\epsilon}-f_{0}^{\epsilon})$ and note that $(R^{\epsilon})_{\epsilon>0}$ is again tight, as assumed in Theorem 8.7. Then using Lemmas 8.9 and 8.10 as in previous parts,
$\mathcal{K}_{0,t_{*}}^{\epsilon}f_{i}^{\epsilon}-B_{t_{*},1}^{\epsilon}=\mathcal{K}_{0,t_{*}}^{\epsilon}f_{i}^{\epsilon}-\mathcal{K}_{0,t_{*}}^{\epsilon}f_{0}^{\epsilon}+\mathcal{K}_{0,t_{*}}^{\epsilon}f_{0}^{\epsilon}(0)\leq R^{\epsilon}+\mathcal{K}_{0,t_{*}}^{\epsilon}f_{0}^{\epsilon}(0)+E_{\epsilon},$
where $E_{\epsilon}\stackrel{{\scriptstyle\mathbb{P}}}{{\to}}0$ as $\epsilon\to 0$. Note that $\mathcal{K}_{0,t_{*}}^{\epsilon}f_{0}^{\epsilon}(0)$ is also tight according to Theorem 8.6. Now let $R$ be a joint subsequential limit of $R^{\epsilon}+\mathcal{K}_{0,t_{*}}^{\epsilon}f_{0}^{\epsilon}(0)$ taken with the other random variables. The above display together with the monotonicity argument from previous parts (again using Lemmas 8.9 and 8.10) implies that if $h\in\mathscr{H}$ is any function with $h(x)\geq R+1$ for all large enough $|x|$, and $B_{t_{*},1}(x)+h(x)>\tilde{\mathcal{K}}_{0,t_{*}}f_{i}(x)$ for all $x$, then
$\mathcal{K}_{t_{*},t}(B_{t_{*},1}+h)\geq\tilde{\mathcal{K}}_{0,t}f_{i},$
for all $t\in T_{*}$, $t>t_{*}$. Finally, by choosing any decreasing sequence $\{h_{m}\}_{m\in\mathbb{N}}$ satisfying $B_{t_{*},1}+h_{m}\downarrow\tilde{\mathcal{K}}_{0,t_{*}}f_{i}$ yields the conclusion. (Here the existence of such a sequence is guaranteed by the fact that both $B_{t_{*},1}$ and $\mathcal{K}_{0,t_{*}}f_{i}$ are continuous, due to the fact that $\mathcal{K}_{0,t_{*}}f_{i}\stackrel{{\scriptstyle d}}{{=}}\mathfrak{h}_{t_{*}}f_{i}$ by Theorem 8.6, and that $B_{t_{*},1}$ is a Brownian motion.) ∎
Proof of Theorem 8.7.
Without loss of generality, below we assume that $T\subset\mathbb{Q}_{+}$ and is finite.
Take a joint ($\epsilon\to 0$) subsequential limit $\{\tilde{\mathcal{K}}_{0,t}f_{i}\}_{t\in\mathbb{Q}_{+},i\in\llbracket 1,k\rrbracket}$ of $\{\mathcal{K}^{\epsilon}_{0,t}f_{i}^{\epsilon}\}_{t\in\mathbb{Q}_{+},i\in\llbracket 1,k\rrbracket}$, together with the random variables leading to (45). By Lemma 8.8 and parts $1,2$ of Lemma 8.11, the collection of random variables in (45) satisfies the conditions of Theorem 5.2, and hence can be given by a directed landscape $\mathcal{L}$ as in that theorem.
Take any $t_{*}\in\mathbb{Q}_{+}$ and $t_{*}<\min T$. Now, using part 3 of Lemma 8.11 (with $T_{*}=T\cup\{t_{*}\}$), for each $i\in\llbracket 1,k\rrbracket$, we can find random sequences $\{h_{m}\}_{m\in\mathbb{N}}$ such that for any $t\in T$,
$B_{t_{*},1}+h_{m}\downarrow\tilde{\mathcal{K}}_{0,t_{*}}f_{i},\quad(B_{t_{*},1}+h_{m})\diamond\mathcal{L}_{t_{*},t}\geq\tilde{\mathcal{K}}_{0,t}f_{i}.$
Noting that $(B_{t_{*},1}+h_{m})\diamond\mathcal{L}_{t_{*},t}\stackrel{{\scriptstyle d}}{{\to}}\tilde{\mathcal{K}}_{0,t}f_{i}$ as $m\to\infty$, we can apply Lemma 5.1 to conclude that $(B_{t_{*},1}+h_{m})\diamond\mathcal{L}_{t_{*},t}\stackrel{{\scriptstyle\mathbb{P}}}{{\to}}\tilde{\mathcal{K}}_{0,t}f_{i}$ as $m\to\infty$, and so
$\tilde{\mathcal{K}}_{0,t}f_{i}=(\tilde{\mathcal{K}}_{0,t_{*}}f_{i})\diamond\mathcal{L}_{t_{*},t}.$
Finally, taking $t_{*}\to 0$ and using that $(\tilde{\mathcal{K}}_{0,t_{*}}f_{i})\diamond\mathcal{L}_{t_{*},t}\stackrel{{\scriptstyle\mathbb{P}}}{{\to}}f_{i}\diamond\mathcal{L}_{0,t}$ in this limit yields the result. ∎
8.3 Monotonicity in general AEPs
In the final subsection we prove approximate monotonicity for general AEPs in the basic coupling, Lemma 8.9. The proof is based on an idea from [ACH24b, Appendix D.3] for the stochastic six-vertex model.
The basic coupling of $(\eta^{-}_{t})_{t\geq 0}$ and $(\eta^{+}_{t})_{t\geq 0}$ in Lemma 8.9 can be encoded by a multi-type exclusion process $(\gamma_{t})_{t\geq 0}$, where for each $t\geq 0$ and $x\in\mathbb{Z}$, we let
$\gamma_{t}(x)=\begin{cases}1,\quad&\text{if }\eta^{-}_{t}(x)=\eta^{+}_{t}(x)=1,\\ 0,\quad&\text{if }\eta^{-}_{t}(x)=\eta^{+}_{t}(x)=0,\\ 2,\quad&\text{if }\eta^{-}_{t}(x)=0,\;\eta^{+}_{t}(x)=1,\\ 3,\quad&\text{if }\eta^{-}_{t}(x)=1,\;\eta^{+}_{t}(x)=0.\end{cases}$
We interpret the values $1$, 00, $2$, $3$ as representing an (ordinary) particle, a hole, a type-2 particle’, and a type-3 particle’. Each particle, type-2 particle, and type-3 particle jumps according to $p$, whenever the target location is empty. In addition, for any $x\in\mathbb{Z}$ and $v\in\mathbb{Z}\setminus\{0\}$ with $p(v)>0$,
  • if there is a particle at $x$, and a type-2 or type-3 particle at $x+v$, they swap with rate $p(v)$;
  • if there is a type-2 (resp. type-3) particle at $x$ and a type-3 (resp. type-2) particle at $x+v$, with rate $p(v)$ the type-2 (resp. type-3) particle moves to $x+v$ to annihilate with the type-3 (resp. type-2) particle, i.e., after the jump there is a hole at site $x$ and an ordinary particle at $x+v$.
Below we use the notion of a move to denote one operation under these dynamics, i.e. a jump, a swap, or an annihilation. We denote $r=\max\{|v|:p(v)>0\}$ for the maximum range of movement and let $R=10r^{r}$.
55-5- 511-1- 122223333444488881010101022-2- 2002222333355-5- 511-1- 122223333444488881010101022-2- 20022223333
55-5- 511-1- 12222333344441010101022-2- 200333355-5- 511-1- 122223333444488881010101022-2- 20022223333
Figure 2: Top panel: an illustration of the evolution of labels under one jump of a type-2 particle. Bottom panel: an illustration of the evolution of labels under one annihilation, where we randomly choose a type-2 label ($8$) and a type-3 label ($2$) to be removed. (In both panels, black/red/blue/white balls denote ordinary particles/type-2 particles/type-3 particles/holes.)
Labels. For technical reasons to handle non-nearest neighbor moves, we make an important augmentation to the above dynamics, via labels.
We assign each type-2 and each type-3 particle a distinct real number, which we call its label. [^2] As will be seen shortly, we only care about the relative ordering of the labels, rather than the precise values. We require that labels on type-$2$ particles increase from left to right, as must labels on type-$3$ particles.
Whenever a move happens, the labels evolve under the following rules:
  • Under any move, labels for all but finitely many particles remain unchanged.
  • The type-2 (resp. type-3) labels are always increasing from left to right.
  • The set of all type-2 (resp. type-3) labels either remains the same, or has one label removed.
These rules completely determine the evolution of the labels under either a jump or swap (see the top panel of Figure 2). However, when an annihilation happens, we have some freedom in choosing which pair of labels (one type-2 and one type-3) is removed. We make this choice randomly. Whenever an annihilation happens from a particle jumping from site $x$ to $x+v$, we consider all labels in $\llbracket x-R,x+R\rrbracket$ before the annihilation. Note that there is at least one type-2 label and at least one type-3 label among this set. We uniformly choose one type-2 label to remove, and uniformly choose one type-3 label to remove. The remaining type-2 (resp. type-3) labels are assigned to the remaining type-2 (resp. type-3) particles in the window $\llbracket x-R,x+R\rrbracket$, in the unique way that ensures the labels increase from left to right. (See the bottom panel of Figure 2.)
Takeovers. For any type-2 label $X$, a ‘takeover’ for $X$ refers to the event where a type-3 label moves from its right to its left for the first time. In addition, in an annihilation where another type-2 label and a type-3 label are removed, if the type-2 label was to the left of $X$ before the annihilation, and the type-3 label was to the right of $X$ before the annihilation, and has never been to the left of $X$ before, then this is also a ‘takeover’ for $X$. This is because we think of the type-3 label first going to the left of $X$ before being removed.
We note that multiple takeovers for a given type-2 label $X$ can happen as a consequence of one move. Moreover, $X$ may have a takeover when an annihilation happens within distance $R$ of $X$, due to the randomized label removal procedure at annihilation.
For any type-3 label, we can define takeovers analogously, to be the event that a type-2 label goes from its left to its right for the first time.
Lemma 8.12.
There are constants $\hat{C}>0$ and $10<\theta<$ 1 depending only on $p$, such that for any type-2 label $X$ and $m\in\mathbb{N}$, the probability that at least $m$ takeovers of $X$ happen (before $X$ is removed in an annihilation) is at most $\hat{C}\theta^{m}$.
The above bound on takeovers is essentially reduced to the following statement on the probability of removal for a given label.
Lemma 8.13.
There is a constant $10<\vartheta<$ 1 depending only on $p$ such that the following is true. Take any $s\geq 0$. Conditional on the configuration $\gamma_{s}$, if there are sites $x,y\in\mathbb{Z}$ with $\gamma_{s}(x)=2,\gamma_{s}(y)=3$ and $R-2r\leq|x-y|<R-r$, with probability $>\vartheta$ the type-2 label at $x$ will be removed before another $R$ takeovers of it.
Proof.
Under the setup stated, without loss of generality, we assume that $x<y$, i.e. $y-x\in\llbracket R-2r,R-r-1\rrbracket$. The main task of this proof is to establish the following combinatorial statement.
Claim: Starting from the configuration $\gamma_{s}$, one can construct a sequence of moves within $\llbracket x,y\rrbracket$ (i.e. all jumps, swaps, and annihilations involve a pair of sites in $\llbracket x,y\rrbracket$) such that the last move is an annihilation.
To prove the claim, we introduce some notations. Denote $R_{1}=R/10=r^{r}$. Denote $g_{-}\in\mathbb{N}$ to be the greatest common divisor of $\{v\in\mathbb{Z}_{-}:p(v)>0\}$, and $g_{+}\in\mathbb{N}$ to be the greatest common divisor of $\{v\in\mathbb{Z}_{+}:p(v)>0\}$. In the case where either of the sets is empty, set $g_{-}=0$ or $g_{+}=0$. From the assumption that $\{v:p(v)+p(-v)>0\}$ additively generates $\mathbb{Z}$, $\gcd(g_{-},g_{+})=1$. Note that at least one of $g_{-}$ and $g_{+}$ is nonzero. Without loss of generality we assume that $g_{-}\neq 0$. We let $G_{-}\subset\mathbb{Z}_{-}$ be the semi-group that is additively generated by $\{v\in\mathbb{Z}_{-}:p(v)>0\}$, and $G_{+}\subset\mathbb{Z}_{+}$ be the semi-group that is additively generated by $\{v\in\mathbb{Z}_{+}:p(v)>0\}$. Then we must have $g_{-}\mathbb{Z}\cap(-\infty,-R_{1}]\subset G_{-}$, and $g_{+}\mathbb{Z}\cap[R_{1},\infty)\subset G_{+}$ if $g_{+}\neq 0$.
We now prove the above claim.
Proof: We argue by contradiction. Namely, below we assume that, starting from $\gamma_{s}$, we cannot find a sequence of moves within $\llbracket x,y\rrbracket$ that contains an annihilation.
For any configuration $\gamma:\mathbb{Z}\to\{0,1,2,3\}$, we define its weight to be the sum $\sum_{z\in\llbracket x+1,y-1\rrbracket:\gamma(z)\in\{1,3\}}z$. We now consider all configurations that can be obtained from $\gamma_{s}$ by taking a sequence of moves within $\llbracket x+1,y-1\rrbracket$, and let $\gamma_{*}$ be the one with the smallest weight. If there are multiple such configurations, choose one arbitrarily. Note that by our assumption above, no move in this sequence (from $\gamma_{s}$ to $\gamma_{*}$) can be an annihilation. We then denote $\Gamma=\{z\in\llbracket x+1,y-1\rrbracket:\gamma_{*}(z)\in\{1,3\}\}$. Then for any $z\in\Gamma$, and $v\in\mathbb{Z}_{-}$ with $z+v\geq x+1$ and $p(v)>0$, necessarily $z+v\in\Gamma$. More generally, $(\Gamma+G_{-})\cap\llbracket x+1,y-1\rrbracket\subset\Gamma$.
Now we consider two scenarios: either (1) $\llbracket y-3R_{1},y-1\rrbracket\cap\Gamma=\emptyset$, or (2) there exists some $z_{*}\in\llbracket y-3R_{1},y-1\rrbracket$ with $z_{*}\in\Gamma$.
In the first scenario, starting from $\gamma_{*}$, we can take a sequence of jumps of the type-3 particle that is at $y$ in $\gamma_{*}$, so that it ends up at some $y_{1}\in(y+G_{-})\cap\llbracket y-3R_{1},y-1\rrbracket$. Then if $g_{+}\neq 0$, we can take another sequence of jumps of this type-3 particle, so that it ends up at some $y_{2}\in(y_{1}+G_{+})\cap\llbracket y_{1},y-1\rrbracket$. If $g_{+}=0$, we just let $y_{2}=y_{1}$. Since $g_{-}\mathbb{Z}\cap(-\infty,-R_{1}]\subset G_{-}$, and $g_{+}\mathbb{Z}\cap[R_{1},\infty)\subset G_{+}$ if $g_{+}\neq 0$, and the greatest common divisor of $g_{-}$ and $g_{+}$ is $1$, $y_{2}$ can be any number in $\llbracket y-R_{1}-r,y-R_{1}\rrbracket$. Take any $v_{-}\in\mathbb{Z}_{-}$ such that $p(v_{-})>0$. (As we assumed that $g_{-}\neq 0$, such a $v_{-}$ exists). We can choose $y_{2}$ such that $y_{2}\equiv x\mod|v_{-}|$. Now that there is a type-2 particle at $y_{2}$, and a type-3 particle at $x$, we can take a sequence of moves within $(x+|v_{-}|\mathbb{Z})\cap\llbracket x,y_{2}\rrbracket$, each move involving two locations $z$ and $z+v_{-}$ for some $z\in(x+|v_{-}|\mathbb{Z})\cap\llbracket x+|v_{-}|,y_{2}\rrbracket$, moving all ordinary particles as far as possible to the left and empty sites as far as possible to the right. Then there will be some $z$ and $z+v_{-}$ in $(x+|v_{-}|\mathbb{Z})\cap\llbracket x,y_{2}\rrbracket$, such that there is a type-2 particle at one of these sites and a type-3 particle at the other. Moving one of these particles onto the other then yields an annihilation, which contradicts the assumption.
In the second scenario, we have that $\llbracket x+1,y-4R_{1}\rrbracket\cap(z_{*}+b_{-}\mathbb{Z})\subset\Gamma$, since $\llbracket x+1,y-4R_{1}\rrbracket\cap(z_{*}+g_{-}\mathbb{Z})\subset z_{*}+G_{-}$. If $g_{+}=0$, necessarily $g_{-}=1$, then we have $\llbracket x+1,y-4R_{1}\rrbracket\subset\Gamma$. If $g_{+}\neq 0$, we take any $v_{+}\in\mathbb{Z}_{+}$ such that $p(v_{+})>0$. Then for any $z^{\prime}\in(z_{*}+G_{-})\cap\llbracket x+1,z_{*}-v_{+}\rrbracket$, we must have $z^{\prime}+v_{+}\in\Gamma$. Indeed, otherwise, we can find a sequence of moves leading to a new configuration $\gamma_{*}^{\prime}$, such that for the set
$\Gamma^{\prime}=\{z\in\llbracket x+1,y-1\rrbracket:\gamma_{*}^{\prime}(z)\in\{1,3\}\},$
$\Gamma^{\prime}\setminus\Gamma=\{z^{\prime}+v\}$ and $\Gamma\setminus\Gamma^{\prime}=\{z_{*}\}$. To see that such a sequence exists, let $z_{0}=z_{*},z_{1},\dots,z_{k}=z^{\prime}$ be a sequence with $z_{i}-z_{i+1}\in\{v\in\mathbb{Z}_{-}:p(v)>0\}$. Since $(\Gamma+G_{-})\cap\llbracket x+1,y-1\rrbracket\subset\Gamma$, $z_{i}\in\Gamma$ for all $i$. Then first moving the particle at site $z_{k}$ to $z^{\prime}+v$, then the particle at $z_{k-1}$ to $z_{k}$, $z_{k-2}$ to $z_{k-1}$ and so on until $z_{0}$ moves to $z_{1}$ yields $\gamma_{*}^{\prime}$. The new configuration $\gamma_{*}^{\prime}$ would have a smaller weight than $\gamma_{*}$, contradicting our choice of $\gamma_{*}$.
Therefore, for any $z^{\prime}\in(z_{*}+G_{-})\cap\llbracket x+1,z_{*}-v_{+}\rrbracket$, we must have $(z^{\prime}+G_{+})\cap\llbracket x+1,z_{*}\rrbracket\subset\Gamma$. Combining this with the fact that $\llbracket x+1,y-4R_{1}\rrbracket\cap(z_{*}+g_{-}\mathbb{Z})\subset z_{*}+G_{-}$, and that $\gcd(g_{-},g_{+})=1$, we necessarily have $\llbracket x+2R_{1},z_{*}\rrbracket\subset\Gamma$. Further using that $(z+G_{-})\cap\llbracket x+1,y-1\rrbracket\subset\Gamma$ for any $z\in\Gamma$, we have that $\llbracket x+1,z_{*}\rrbracket\subset\Gamma$. In summary, whether $g_{+}=0$ or not, we have $\llbracket x+1,y-4R_{1}\rrbracket\subset\Gamma$.
At this point we apply a reflection to the entire picture, by defining a new configuration $\gamma^{*}$ as follows: for any $z\in\mathbb{Z}$, we let $\gamma^{*}(z)=1$ if $\gamma_{*}(x+y-z)=0$, $\gamma^{*}(z)=0$ if $\gamma_{*}(x+y-z)=1$, $\gamma^{*}(z)=2$ if $\gamma_{*}(x+y-z)=3$, and $\gamma^{*}(z)=3$ if $\gamma_{*}(x+y-z)=2$. The configuration $\gamma^{*}$ puts us back in scenario (1), and any sequence of moves starting from $\gamma_{*}$ can be reflected to a sequence of moves starting from $\gamma^{*}$. As we showed above, we can create an annihilation starting from $\gamma^{*}$, so we can also find a sequence of moves that reach an annihilation from $\gamma_{*}$. This contradicts our ‘no annihilation’ assumption in the second scenario as well. $\blacksquare$
Now, the number of moves required by the claim is upper bounded by a constant depending only on $p$, since there are only finitely many configurations in $\llbracket x,y\rrbracket$. Therefore, with positive probability depending only on $p$, this exact sequence of moves occurs, and no other move in $\llbracket x-r,y+r\rrbracket$ happens before the end of this sequence. When any annihilation happens (in this sequence of moves), with probability at least $1/(2R)$ the type-2 label at $x$ in $\gamma_{s}$ is removed. Also note that at most $R$ takeovers of the type-2 label happen before the end of this sequence, and so the conclusion follows.
We can now deduce the bound on the number of takeovers by repeatedly using the removal estimate.
Proof of Lemma 8.12.
For a type-2 label, we say that it is ready-to-remove if there exists one potential annihilation which could remove it. Then there exists some $10<c_{0}<$ 1 depending only on $p$, such that for any $t>0$, given the configuration at time $t$ with a type-2 label that is ready-to-remove, with probability at least $c_{0}$ the type-2 label will be removed before its next takeover.
Below we take a type-2 label $X$ at time 00. Let $t_{0}=0$. For any $i\in\mathbb{N}$, let $t_{i}$ be the time of its $i$-th takeover, and set $t_{i}=\infty$ if less than $i$ takeovers happen to label $X$. For $i\in\mathbb{Z}_{\geq 0}$, we then let $\tau_{i}$ be the first time after $t_{i}$, such that either $X$ is ready-to-remove, or there is a type-3 particle to its right, with distance $\geq R-2r$ and $<R-r$. Then $\tau_{i}$ is a stopping time, and may equal $\infty$. (In particular, if $t_{i}=\infty$ we have $\tau_{i}=\infty$.)
For any $i\in\mathbb{Z}_{\geq 0}$, we note that at most $R$ takeovers can happen to $X$ between $t_{i}$ and $\tau_{i}$. Also given $\tau_{i}$ which is <absent<$\infty$< ∞, by Lemma 8.13 and the statement at the beginning of this proof, with probability $>\vartheta\wedge c_{0}$ the label $X$ will be removed before $R$ takeovers after $\tau_{i}$. Therefore, we have that $\mathbb{P}[t_{i+R}<\infty\mid t_{i}<\infty]<1-\vartheta\wedge c_{0}$. Thus the conclusion follows. ∎
Finally, we derive the approximate monotonicity of AEP, via estimating height functions using the number of takeovers.
Proof of Lemma 8.9.
Let $C_{*}$ be a sufficiently large constant depending on $p$. Consider the following events:
  • $\mathcal{E}_{1}$:
    Every type-2 or type-3 label that enters $\llbracket-w-3R,w+3R\rrbracket$ between time 00 and $t$ is contained in the set $\llbracket-C_{*}(t+w),C_{*}(t+w)\rrbracket$ at time 00.
  • $\mathcal{E}_{2}$:
    For every type-2 or type-3 label that is contained in $\llbracket-C_{*}(t+w),C_{*}(t+w)\rrbracket$ at time 00, at most $m$ takeovers happen (before it is removed).
Claim: Under $\mathcal{E}_{1}\cap\mathcal{E}_{2}$, we must have $Y_{t,w}\leq m+2R$.
Proof: For any $x\in\llbracket-w,w\rrbracket$, we consider the last time (before $t$) that a type-2 or type-3 label goes from $\leq x-1$ to $\geq x$, or vice versa. This also includes the event where an annihilation happens between a label $\leq x-1$ and a label $\geq x$. If no such time exists, then $\eta^{-}_{t}(x)-\eta^{+}_{t}(x)=\eta^{-}_{0}(x)-\eta^{+}_{0}(x)\leq 0$. Otherwise, let $s\in(0,t]$ be the last such time. Without loss of generality, we assume that a type-2 label goes from $\leq x-1$ to $\geq x$ at time $s$, or it is at $\leq x-1$ right before time $s$ and annihilates with a type-3 label $\geq x$ at time $s$. Denote the trajectory of that type-2 label by $(w(s^{\prime}))_{0\leq s^{\prime}<s}$. Then we have $\eta^{-}_{t}(x)-\eta^{+}_{t}(x)\leq\eta^{-}_{s-}(w(s-))-\eta^{+}_{s-}(w(s-))+2R$. On the other hand, since $\eta^{-}_{0}(w(0))\leq\eta^{+}_{0}(w(0))$, we have that for any $0\leq s^{\prime}<s$, $\eta^{-}_{s^{\prime}}(w(s^{\prime}))-\eta^{+}_{s^{\prime}}(w(s^{\prime}))$ is at most the number of takeovers that happened to this type-2 label between time 00 and $s^{\prime}$. By $\mathcal{E}_{1}$ we have that $w(0)\in\llbracket-C_{*}(t+w),C_{*}(t+w)\rrbracket$. By $\mathcal{E}_{2}$ we have that at most $m$ takeovers happen to this type-2 label. Thus $\eta^{-}_{s-}(w(s-))-\eta^{+}_{s-}(w(s-))\leq m$, further implying that $\eta^{-}_{t}(x)-\eta^{+}_{t}(x)\leq m+2R$. $\blacksquare$
It now remains to lower bound the probability of $\mathcal{E}_{1}\cap\mathcal{E}_{2}$.
By Lemma 8.12, we have $\mathbb{P}[\mathcal{E}_{2}]>1-C(t+w)\theta^{m}$. As for $\mathcal{E}_{1}$, we consider any type-2 or type-3 label that is at location $x<-C_{*}(t+w)$ at time 00. We can couple it with a random walk with a different jump rate $p^{\prime}$ (depending on $p$), such that it is always to the left of that random walk. Therefore, by taking $C_{*}$ large enough (depending on $p$), with have that with probability $>1-C\exp(c(w+x))$, it is to the left of $-w-3R$ up to time $t$. By taking a union bound over $x<-C_{*}(t+w)$, we have that with probability $>1-C\exp(-c(t+w))$, any type-2 or type-3 label that is to the left of $-C_{*}(t+w)$ at time 00 is to the left of $-w-3R$ up to time $t$.
Similarly, we have that with probability $>1-C\exp(-c(t+w))$, any type-2 or type-3 label that is to the right of $C_{*}(t+w)$ at time 00 is to the right of $w+3R$ up to time $t$. Thus we get that $\mathbb{P}[\mathcal{E}_{1}]>1-C\exp(-ct)$, and so $\mathbb{P}[\mathcal{E}_{1}\cap\mathcal{E}_{2}]>1-C(t+w)\theta^{m}-C\exp(-ct)$. The conclusion follows. ∎